### v1

# Regression-Based Single-Point Zeroth-Order Optimization

Abstract
Zeroth-order optimization (ZO) is widely used for solving black-box optimization and control problems. In particular, single-point ZO (SZO) is well-suited to online or dynamic problem settings due to its requirement of only a single function evaluation per iteration. However, SZO suffers from high gradient estimation variance and slow convergence, which severely limit its practical applicability. Despite recent advances, the convergence of existing SZO methods remains inferior to that of two-point ZO methods. To overcome this limitation, we propose a novel yet simple SZO framework, termed regression-based SZO (RESZO), which substantially enhances the convergence rate. Unlike conventional ZO methods that rely solely on function evaluations at the current point for gradient estimation, RESZO improves gradient estimation by effectively leveraging historical function evaluations from previous iterations. Specifically, RESZO constructs a surrogate function via regression using recent historical evaluations and employs the gradient of this surrogate function for iterative updates. Two variants of RESZO, which fit linear and quadratic surrogate functions respectively, are introduced. Theoretically, we provide a non-asymptotic convergence analysis for the linear variant of RESZO, showing that its convergence rates are comparable to those of two-point ZO methods for both smooth nonconvex and strongly convex functions. Moreover, extensive numerical experiments demonstrate that RESZO not only matches but empirically outperforms two-point ZO in terms of function query complexity.
1 Introduction
In this paper, we consider solving a generic black-box unconstrained optimization problem (1):
$\min_{\bm{x}\in\mathbb{R}^{d}}\,f(\bm{x}),$ (1)
where $\bm{x}\in\mathbb{R}^{d}$ is the decision variable of dimension $d$. The objective function $f:\mathbb{R}^{d}\to\mathbb{R}$ is unknown, and only its zeroth-order information, namely function evaluations, is accessible. As a result, the first-order gradients of function $f$ are unavailable, and problem (1) can be solved solely based on function evaluations. This type of black-box optimization problems has attracted considerable recent attention and covers a broad range of applications, including power system control [1, 2], reinforcement learning [3, 4], neural network training [5, 6], sensor management [7], etc.
Zeroth-order optimization (ZO) methods [7, 8] have been extensively developed and applied to solve black-box optimization problems, offering theoretical convergence guarantees. Essentially, ZO methods estimate unknown gradients using perturbed function evaluations. According to the number of queried function evaluations per iteration, ZO methods can be categorized into single-point, two-point, and multi-point ZO schemes [7]. Single-point ZO (SZO) methods [9, 10, 11] require only a single function evaluation per iteration for gradient estimation, making them well-suited for online optimization or dynamic control applications. Because in such settings, the system environment is non-stationary, or each function evaluation involves implementing a decision that directly changes the system environment, rendering multiple function queries at the same instant infeasible. However, single-point ZO suffers from large estimation variance and slow convergence, severely limiting its practical applicability. Despite recent advancements [10, 11, 12, 13, 14], the convergence of single-point ZO remains slower than that of two-point ZO or multi-point ZO, which use multiple function evaluations for gradient estimation per iteration and achieve lower variance.
Motivation. At each iteration $t$, classic ZO methods estimate the gradient $\nabla f(\bm{x}_{t})$ solely based on the function evaluations at the current point $\bm{x}_{t}$, while all historical function evaluations obtained from previous iterations are discarded. However, these past function evaluations contain valuable information about the unknown function and can help with the gradient estimations for subsequent iterations, especially those obtained in the most recent iterations. Moreover, in many practical applications, such as clinical trials, real-world experiments, and long-term complex simulations, obtaining a single function evaluation can be costly and laborious, making it crucial to fully exploit all available evaluations to enhance efficiency. Then, a natural question is
Can available historical function evaluations from previous iterations be leveraged to enhance single-point zeroth-order gradient estimation?"
Contribution. Motivated by this question, we propose a novel yet simple single-point ZO framework, termed Regression-based Single-point Zeroth-order Optimization (RESZO). In contrast to classic ZO methods that rely solely on function evaluations at the current point for gradient estimation, the proposed RESZO method significantly enhances gradient estimation by leveraging available historical function evaluations from previous iterations. Specifically, the RESZO method queries the function value only once at each iteration, and fits a surrogate model based on recent historical function evaluations via regression. Then, the surrogate model’s gradient is used as the estimated gradient for iterative updates. We introduce two variants of the RESZO method, termed L-RESZO and Q-RESZO, which fit linear and quadratic surrogate functions, respectively. A key ingredient of the proposed RESZO method is adding random perturbations to function evaluations to ensure sufficient exploration. Theoretically (see Table 1), we show that for both smooth nonconvex and strongly convex functions, the linear variant of our algorithm (L-RESZO) achieves convergence rates that are only worse than those of two-point ZO methods by a factor $C_{d}>0$, where $C_{d}$ is a parameter defined in (21) in Assumption 2. This represents a substantial improvement over the rates achieved by existing SZO methods (see Table 1). While theoretically bounding the term $C_{d}$ is challenging, strong empirical evidence shows that $C_{d}$ scales as $\mathcal{O}(\sqrt{d})$ under appropriate algorithmic settings; see Section 4 for more details. To the best of our knowledge, this is the first work to present single-point ZO methods whose convergence rates match those of two-point ZO methods. Moreover, extensive numerical experiments demonstrate that RESZO not only matches but empirically outperforms two-point ZO in terms of function query complexity.
Table 1: Iteration complexity of ZO methods for solving Lipschitz and smooth objective functions.
Method Literature Complexity
$\mu$-Strongly convex Non-convex
Classic SZO [14] $\mathcal{O}\!\bigl{(}d^{2}/\varepsilon^{3}\bigr{)}$
Residual-feedback SZO [15] $\mathcal{O}\!\bigl{(}d^{3}/\varepsilon^{\frac{3}{2}}\bigr{)}$ $\mathcal{O}\!\bigl{(}d^{3}/\varepsilon^{\frac{3}{2}}\bigr{)}$
HLF-SZO [11] $\mathcal{O}\!\bigl{(}d^{\frac{3}{2}}/\varepsilon^{\frac{3}{2}}\bigr{)}$ $\mathcal{O}\!\bigl{(}d^{\frac{3}{2}}/\varepsilon^{\frac{3}{2}}\bigr{)}$
L-RESZO This work $\mathcal{O}\!\Bigl{(}C_{d}\frac{d}{\mu}\log\!\frac{1}{\varepsilon}\Bigr{)}$ $\mathcal{O}\!\bigl{(}C_{d}d/\varepsilon\bigr{)}$
Two-point ZO [8] $\mathcal{O}\!\Bigl{(}\frac{d}{\mu}\log\!\frac{1}{\varepsilon}\Bigr{)}$ $\mathcal{O}\!\bigl{(}d/\varepsilon\bigr{)}$
*In the strongly convex case, the accuracy is measured by $\,f(x_{T})-f^{*}\leq\varepsilon$, where $f^{*}:=\min_{x\in\mathbb{R}^{d}}f(x)$. In the non-convex case, the accuracy is measured by $\tfrac{1}{T}\sum_{k=1}^{T}\lVert\nabla f(x_{k})\rVert^{2}\leq\varepsilon$.
Related Work. The classic SZO method, formulated in (2), is first proposed by [9], which estimates gradients by querying only a single function value per iteration. In [13], a self-concordant barrier function is used to enhance the convergence of SZO for solving bandit smooth convex optimization. By leveraging high-order smoothness, reference [16] derives an improved complexity bound for SZO in the context of stochastic convex optimization. To further enhance the convergence of SZO, reference [15] proposes a residual-feedback SZO method, formulated in (2), which reuses the function evaluation from the last iteration for gradient estimation, resulting in significant performance improvement. Additionally, reference [11] leverages the idea of high-pass and low-pass filters from extremum seeking control to accelerate SZO convergence. Notably, the integration of a high-pass filter coincides with the residual-feedback approach in [15]. This paper follows a similar direction to [11, 15]; however, rather than reusing only the most recent function evaluation, we leverage many historical function evaluations from previous iterations to fit surrogate models. In addition, our RESZO method is related to the linear regression and interpolation approaches [17, 18, 19] used in derive-free optimization (DFO) [20]. The fundamental difference is that these methods [17, 18, 19] perform multiple function evaluations near the current point for linear interpolation or regression. In contrast, our approach queries the function value only once at each iteration and fits surrogate models by fully recycling historical function evaluations from past iterations.
2 Preliminaries on Zeroth-Order Optimization
The classic single-point ZO (SZO) method [9] that solves (1) is given by (2):
[…]
where $\delta\!>\!0$ is a scalar parameter called smoothing radius, $\eta\!>\!0$ is the step size, and $\bm{u}_{t}\in\mathbb{R}^{d}$ is an independent identically distributed random direction vector sampled from $\mathrm{Unif}(\mathbb{S}_{d-1})$, namely the uniform distribution on the unit sphere $\mathbb{S}_{d-1}$, for all $k=0,1,2,\cdots$. The random direction vector $\bm{u}_{t}$ can also be sampled from the multivariate Gaussian distribution $\mathcal{N}(\mathbf{0},\bm{I})$ or other zero-mean distributions. In (2), $\mathsf{G}^{1}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)$ denotes the SZO gradient estimator of function $f$.
Recent works [15, 11] have introduced a new variant of SZO, referred to as residual-feedback SZO (RSZO) [15] or high-pass filter-integrated SZO [11], which is given by (2):
[…]
$\qquad\bm{x}_{t+1}=\bm{x}_{t}-\eta\cdot\underbrace{\frac{d}{\delta}\Big{(}f(\bm{x}_{t}+\delta\bm{u}_{t})-f(\bm{x}_{t-1}+\delta\bm{u}_{t-1})\Big{)}\bm{u}_{t}}_{\coloneq\,\mathsf{G}^{\mathrm{r}1}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)},\qquad t=1,2,\cdots.$ (3)
At each iteration $t$, the gradient estimator $\mathsf{G}^{\mathrm{r}1}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)$ of the RSZO method (2) reuses the function evaluation $f(\bm{x}_{t\!-\!1}\!+\!r\bm{u}_{t\!-\!1})$ queried at the last iteration $t\!-\!1$. This approach substantially reduces gradient estimation variance and improves convergence compared with SZO (2).
The classic two-point ZO (TZO) method [8] that solves problem (1) is given by (4):
[…]
$\qquad\qquad\bm{x}_{t+1}=\bm{x}_{t}-\eta\cdot\underbrace{\frac{d}{2\delta}\Big{(}f(\bm{x}_{t}+\delta\bm{u}_{t})-f(\bm{x}_{t}-\delta\bm{u}_{t})\Big{)}\bm{u}_{t}}_{\coloneq\,\mathsf{G}^{2}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)},\quad t=0,1,2,\cdots,$ (4)
which employs two function evaluations to estimate the gradient of function $f$ with the two-point gradient estimator $\mathsf{G}^{2}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)$.
The classic ZO gradient estimators above share the same expectation [8]:
$\mathbb{E}_{\bm{u}\sim\mathrm{Unif}(\mathbb{S}_{d-1})}\big{[}\mathsf{G}^{1}_{f}(\bm{x};\bm{u},\delta)\big{]}=\,$ $\mathbb{E}_{\bm{u}\sim\mathrm{Unif}(\mathbb{S}_{d-1})}\big{[}\mathsf{G}^{\mathrm{r}1}_{f}(\bm{x};\bm{u},\delta)\big{]}$
=\,= $\mathbb{E}_{\bm{u}\sim\mathrm{Unif}(\mathbb{S}_{d-1})}\big{[}\mathsf{G}^{2}_{f}(\bm{x};\bm{u},\delta)\big{]}=\nabla\hat{f}(\bm{x};\delta),$ (5)
where $\hat{f}(\bm{x};\delta):=\mathbb{E}_{\bm{v}\sim\mathrm{Unif}(\mathbb{B}_{d})}\big{[}f(\bm{x}+\delta\bm{v})\big{]}$ is a smoothed version of the original function $f$ [9]. Despite having the same expectation, the two-point ZO gradient estimator $\mathsf{G}^{2}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)$ exhibits lower variance and achieves faster convergence, compared with the single-point ZO gradient estimators $\mathsf{G}^{1}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)$ and $\mathsf{G}^{\mathrm{r}1}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)$. See [8, 11] for more details.
3 Regression-Based Single-Point Zeroth-Order Optimization
In this paper, we focus on the class of SZO methods that require only one function evaluation per iteration. Below, we introduce the proposed RESZO method and two specific variants, L-RESZO and Q-RESZO, which fit linear and quadratic surrogate models using historical function evaluations.
3.1 Linear Regression-Based Single-Point Zeroth-Order Optimization (L-RESZO)
At each iteration $t$, the available function evaluations from all previous iterations are given by $\{f(\hat{\bm{x}}_{k})\}_{k=0}^{t}$ with $\hat{\bm{x}}_{k}=\bm{x}_{k}+\delta\bm{u}_{k}$. Here, $\bm{u}_{k}\in\mathbb{R}^{d}$ is a random direction vector sampled from the uniform distribution $\mathrm{Unif}(\mathbb{S}_{d-1})$ over the unit sphere $\mathbb{S}_{d-1}$, or from the multivariate Gaussian distribution $\mathcal{N}(\bm{0},\bm{I})$; and $\delta>0$ is a scalar parameter called smoothing radius. These are consistent with the classic ZO methods (2)-(4). Note that function evaluations are performed at the perturbed points $\hat{\bm{x}}_{k}$ rather than the iterate points ${\bm{x}}_{k}$ to promote sufficient exploration of the optimization space. The necessity of introducing such perturbations is further illustrated by Remark 1 and the numerical experiments in Section 5.2.
Based on the first-order Taylor expansion at the current perturbed point $\hat{\bm{x}}_{t}$, the available function evaluations $\{f(\hat{\bm{x}}_{k})\}_{k=0}^{t}$ can be expressed as (6):
$\qquad\qquad f(\hat{\bm{x}}_{k})=f(\hat{\bm{x}}_{t})+\nabla f(\hat{\bm{x}}_{t})^{\top}(\hat{\bm{x}}_{k}-\hat{\bm{x}}_{t})+\epsilon_{t,k},\qquad k=0,1,\cdots,t,$ (6)
where $\epsilon_{t,k}$ denotes the corresponding high-order residual term. Accordingly, we propose to fit a linear surrogate function $f^{\mathrm{s1}}_{t}(\bm{x})$ in (7) to locally approximate the original function $f(\bm{x})$ in (1):
$f^{\mathrm{s1}}_{t}(\bm{x})\coloneq f(\hat{\bm{x}}_{t})+\bm{g}_{t}^{\top}({\bm{x}}-\hat{\bm{x}}_{t})+c_{t},$ (7)
where the linear coefficient vector $\bm{g}_{t}\in\mathbb{R}^{d}$ and the intercept $c_{t}\in\mathbb{R}$ can be obtained through linear regression on the available sample points $\{\hat{\bm{x}}_{k},f(\hat{\bm{x}}_{k})\}_{k=0}^{t}$. Then, the gradient of the surrogate function $\nabla f^{\mathrm{s1}}_{t}(\bm{x})=\bm{g}_{t}$ can be used as the estimator for the true gradient $\nabla f(\bm{x})$.
To ensure that the fitted surrogate model $f^{\mathrm{s1}}_{t}(\bm{x})$ remains up-to-date with the most recent function evaluations, we employ the recursive linear regression in a sliding window, which uses only the latest $m$ function evaluations $\{\hat{\bm{x}}_{t-i+1},f(\hat{\bm{x}}_{t-i+1})\}_{i=1}^{m}$ to fit the linear model at each iteration $t$. Specifically, at each iteration $t$, we solve the following least squares problem (8) to obtain $\bm{g}_{t}$ and $c_{t}$:
$\min_{\bm{g}_{t},c_{t}}\,\sum_{i=1}^{m}\Big{[}f^{\mathrm{s1}}_{t}(\hat{\bm{x}}_{t-i+1})-f(\hat{\bm{x}}_{t-i+1})\Big{]}^{2}\ \Longrightarrow\ \begin{bmatrix}\bm{g}_{t}\\ c_{t}\end{bmatrix}=(X_{t}^{\top}X_{t})^{\dagger}X_{t}^{\top}\bm{y}_{t},$ (8)
where $\dagger$ denotes the Moore–Penrose pseudoinverse, $\bm{\Delta}^{x}_{t,i}\!\coloneq\!\hat{\bm{x}}_{t-i+1}\!-\!\hat{\bm{x}}_{t},\Delta^{f}_{t,i}\!\coloneq\!f(\hat{\bm{x}}_{t-i+1})\!-\!f(\hat{\bm{x}}_{t})$,
$X_{t}\coloneq\begin{bmatrix}{\bm{\Delta}^{x}_{t,1}}^{\top}&1\\ {\bm{\Delta}^{x}_{t,2}}^{\top}&1\\ \vdots&\vdots\\ {\bm{\Delta}^{x}_{t,m}}^{\top}&1\end{bmatrix}\in\mathbb{R}^{m\times(d+1)},\quad\bm{y}_{t}\coloneq\begin{bmatrix}\Delta^{f}_{t,1}\\ \Delta^{f}_{t,2}\\ \vdots\\ \Delta^{f}_{t,m}\end{bmatrix}\in\mathbb{R}^{m}.$ (9)
Consequently, the proposed Linear Regression-based Single-point Zeroth-order Optimization (L-RESZO) method for solving problem (1) is presented as Algorithm 1. Since the linear regression requires $m$ function evaluations, we initially employ a classical ZO algorithm, such as the RSZO method (2), to perform iterative updates and accumulate sufficient sample points, as shown in Step 3. After $m$ iterations, the algorithm switches to using the linear regression-based gradient estimator $\bm{g}_{t}$ for subsequent iterative updates, as shown in Steps 6-8.
1:  Inputs: Initial point $\bm{x}_{0}$, stepsizes $\tilde{\eta},\eta$, smoothing radius values $\tilde{\delta},\delta$, sample size $m$, and total iteration number $T$.
2:  for $t=0:m-1$ do
3:     Perform the RSZO method (2) with $\tilde{\eta},\tilde{\delta}$ for iterative updates.
4:  end for
5:  for $t=m:T$ do
6:     Randomly sample $\bm{u}_{t}\sim\mathrm{Unif}(\mathbb{S}_{d-1})$; perform a function evaluation on the perturbed value $\hat{\bm{x}}_{t}=\bm{x}_{t}+\delta\bm{u}_{t}$ to obtain $f(\hat{\bm{x}}_{t})$; collect $m$ latest sample points $\{\hat{\bm{x}}_{t-i+1},f(\hat{\bm{x}}_{t-i+1})\}_{i=1}^{m}$.
7:     Solve for the linear coefficient $\bm{g}_{t}$ via (8).
8:     Update with $\bm{x}_{t+1}=\bm{x}_{t}-\eta\cdot\bm{g}_{t}$.
9:  end for
Algorithm 1 Linear Regression-based Single-point Zeroth-order Optimization (L-RESZO)
In addition, the sample size $m$ is typically chosen to be comparable to the dimension $d$ of $\bm{x}$, or slightly larger, to ensure that $X_{t}$ is of full column rank and the matrix $X_{t}^{\top}X_{t}$ in (8) is invertible. In this case, one can replace the pseudoinverse $(\cdot)^{\dagger}$ in (8) by the regular matrix inverse $(\cdot)^{-1}$.
Fast Computation of Matrix Inverse. When the matrix $X_{t}^{\top}X_{t}$ in (8) is invertible, the rank-1 update approach [21] can be employed to enable fast computation of the matrix inverse. Note that performing linear regression based on the sample point set $\{\bm{\Delta}^{x}_{t,i},\Delta^{f}_{t,i}\}_{i=1}^{m}$ yields the same linear coefficient $\bm{g}_{t}$ as linear regression on the sample point set $\{\hat{\bm{x}}_{t-i+1},f(\hat{\bm{x}}_{t-i+1})\}_{i=1}^{m}$. This equivalence holds because the former set is simply a translational shift of the latter by $(\hat{\bm{x}}_{t},f(\hat{\bm{x}}_{t}))$ in the coordinate space. Hence, the least squares problem (8) based on $\{\bm{\Delta}^{x}_{t,i},\Delta^{f}_{t,i}\}_{i=1}^{m}$ can be replaced by the least squares problem (10) based on $\{\hat{\bm{x}}_{t-i+1},f(\hat{\bm{x}}_{t-i+1})\}_{i=1}^{m}$ to obtain $\bm{g}_{t}$:
$\min_{\bm{g}_{t},c_{t}}\,\sum_{i=1}^{m}\Big{(}\bm{g}_{t}^{\top}\hat{\bm{x}}_{t-i+1}+c_{t}-f(\hat{\bm{x}}_{t-i+1})\Big{)}^{2}\ \Longrightarrow\ \begin{bmatrix}\bm{g}_{t}\\ c_{t}\end{bmatrix}=(\hat{X}_{t}^{\top}\hat{X}_{t})^{-1}\hat{X}_{t}^{\top}\hat{\bm{y}}_{t},$ (10)
where
$\hat{X}_{t}\coloneq\begin{bmatrix}{\hat{\bm{x}}_{t}}^{\top}&1\\ {\hat{\bm{x}}_{t-1}}^{\top}&1\\ \vdots&\vdots\\ {\hat{\bm{x}}_{t-m+1}}^{\top}&1\end{bmatrix}\in\mathbb{R}^{m\times(d+1)},\quad\hat{\bm{y}}_{t}\coloneq\begin{bmatrix}f(\hat{\bm{x}}_{t})\\ f(\hat{\bm{x}}_{t-1})\\ \vdots\\ f(\hat{\bm{x}}_{t-m+1})\end{bmatrix}\in\mathbb{R}^{m}.$ (11)
Consequently, when moving from iteration $t$ to $t+1$, the matrix $\hat{X}_{t+1}$ is actually formed by dropping the oldest sample point $\hat{\bm{x}}_{t-m+1}$ from $\hat{X}_{t}$ and adding the latest sample point $\hat{\bm{x}}_{t+1}$ to it. This leads to the update rule (12):
$\hat{X}_{t+1}^{\top}\hat{X}_{t+1}=\hat{X}_{t}^{\top}\hat{X}_{t}-\hat{\bm{x}}_{t-m+1}\hat{\bm{x}}_{t-m+1}^{\top}+\hat{\bm{x}}_{t+1}\hat{\bm{x}}_{t+1}^{\top}.$ (12)
Hence, given the available matrix inverse $(\hat{X}_{t}^{\top}\hat{X}_{t})^{-1}$ obtained in the last iteration, the new matrix inverse $(\hat{X}_{t+1}^{\top}\hat{X}_{t+1})^{-1}$ can be computed efficiently as (13) by using the Sherman–Morrison–Woodbury formula [21] twice in a row: one for dropping $\hat{\bm{x}}_{t-m+1}$ and the other for adding $\hat{\bm{x}}_{t+1}$;
$B=A+\frac{A\hat{\bm{x}}_{t-m+1}\hat{\bm{x}}_{t-m+1}^{\top}A}{1-\hat{\bm{x}}_{t-m+1}^{\top}A\hat{\bm{x}}_{t-m+1}},\qquad(\hat{X}_{t+1}^{\top}\hat{X}_{t+1})^{-1}=B-\frac{B\hat{\bm{x}}_{t+1}\hat{\bm{x}}_{t+1}^{\top}B}{1+\hat{\bm{x}}_{t+1}^{\top}B\hat{\bm{x}}_{t+1}},$ (13)
where $A\coloneq(\hat{X}_{t}^{\top}\hat{X}_{t})^{-1}$ and $B\coloneq(\hat{X}_{t}^{\top}\hat{X}_{t}-\hat{\bm{x}}_{t-m+1}\hat{\bm{x}}_{t-m+1}^{\top})^{-1}$.
This approach avoids recomputing the new matrix inverse $(\hat{X}_{t+1}^{\top}\hat{X}_{t+1})^{-1}$ from scratch and reduces the computational cost of matrix inversion from $\mathcal{O}(d^{3})$ to $\mathcal{O}(d^{2})$.
Remark 1.
(Role of Perturbation). From (11), it is seen that the rows of $\hat{X}_{t}$ are constructed by ${\hat{\bm{x}}_{k}}=\bm{x}_{k}\!+\!\delta\bm{u}_{k}$ for $k=t,\cdots,t\!-\!m\!+\!1$. Intuitively, adding the random perturbations $\delta\bm{u}_{k}$ promotes the full row rank property of $\hat{X}_{t}$, thus making $\hat{X}_{t+1}^{\top}\hat{X}_{t+1}$ tend to be invertible. This enhances the algorithmic stability of the RESZO method. The magnitude of perturbation is controlled by the value of the smoothing radius parameter $\delta$. Since the RESZO method essentially estimates the gradient at the perturbed point, i.e., $\nabla f(\hat{\bm{x}})=\nabla f(\bm{x}+\delta\bm{u}_{t})$, a smaller $\delta$ generally results in lower gradient estimation bias and promotes convergence precision. Accordingly, rather than selecting a fixed $\delta$, we propose an adaptive $\delta$ scheme for the RESZO algorithm in Section 4. The impacts and necessity of incorporating perturbations are further illustrated through numerical experiments in Section 5.2.
3.2 Quadratic Regression-Based Single-Point Zeroth-Order Optimization (Q-RESZO)
In addition to fitting a linear model, other more sophisticated surrogate models, such as a quadratic function, can be constructed based on available historical function evaluations to further enhance gradient estimation. At each iteration $t$, based on the second-order Taylor expansion at the current perturbed point $\hat{\bm{x}}_{t}$, the available function evaluations $\{f(\hat{\bm{x}}_{k})\}_{k=0}^{t}$ can be expressed as:
$f(\hat{\bm{x}}_{k})\!=\!f(\hat{\bm{x}}_{t})\!+\!\nabla f(\hat{\bm{x}}_{t})^{\top}(\hat{\bm{x}}_{k}\!-\!\hat{\bm{x}}_{t})\!+\!\frac{1}{2}(\hat{\bm{x}}_{k}\!-\!\hat{\bm{x}}_{t})^{\top}H_{f}(\hat{\bm{x}}_{t})(\hat{\bm{x}}_{k}\!-\!\hat{\bm{x}}_{t})\!+\!\epsilon_{t,k},\ \ k\!=\!0,1,\cdots,t,$ (14)
where $H_{f}(\cdot)$ denotes the Hessian matrix of function $f$, and $\epsilon_{t,k}$ denotes the corresponding higher-order residual term. Accordingly, we can fit a quadratic surrogate function $f^{\mathrm{s2}}_{t}(\bm{x})$ (15) to locally approximate the original function $f(\bm{x})$ in (1):
$f^{\mathrm{s2}}_{t}(\bm{x})\coloneq f(\hat{\bm{x}}_{t})+\bm{g}_{t}^{\top}({\bm{x}}-\hat{\bm{x}}_{t})+\frac{1}{2}({\bm{x}}-\hat{\bm{x}}_{t})^{\top}\text{diag}(\bm{h}_{t})({\bm{x}}-\hat{\bm{x}}_{t})+c_{t}.$ (15)
Here, we use the diagonal matrix $\text{diag}(\bm{h}_{t})$ with vector $\bm{h}_{t}\in\mathbb{R}^{d}$ to approximate the Hessian matrix, which reduces the number of variables for fitting. Similarly, the recursive linear regression in a sliding window is employed, which uses only the latest $m$ function evaluations $\{\hat{\bm{x}}_{t-i+1},f(\hat{\bm{x}}_{t-i+1})\}_{i=1}^{m}$ to fit the quadratic surrogate model at each iteration $t$.
Then, we solve the following least squares problem (16) to obtain the coefficients $(\bm{g}_{t},\bm{h}_{t},c_{t})$:
$\min_{\bm{g}_{t},\bm{h}_{t},c_{t}}\,\sum_{i=1}^{m}\Big{[}f^{\mathrm{s2}}_{t}(\hat{\bm{x}}_{t-i+1})-f(\hat{\bm{x}}_{t-i+1})\Big{]}^{2}\Longrightarrow\begin{bmatrix}\bm{g}_{t}\\ \bm{h}_{t}\\ c_{t}\end{bmatrix}=(\tilde{X}_{t}^{\top}\tilde{X}_{t})^{\dagger}\tilde{X}_{t}^{\top}{\bm{y}}_{t},$ (16)
where $\odot$ denotes the element-wise multiplication of two column vectors, and
$\tilde{X}_{t}\coloneq\begin{bmatrix}{\bm{\Delta}^{x}_{t,1}}^{\top}&\frac{1}{2}\Big{(}{\bm{\Delta}^{x}_{t,1}}\odot\bm{\Delta}^{x}_{t,1}\Big{)}^{\top}&1\\ {\bm{\Delta}^{x}_{t,2}}^{\top}&\frac{1}{2}\Big{(}{\bm{\Delta}^{x}_{t,2}}\odot\bm{\Delta}^{x}_{t,2}\Big{)}^{\top}&1\\ \vdots&\vdots\\ {\bm{\Delta}^{x}_{t,m}}^{\top}&\frac{1}{2}\Big{(}{\bm{\Delta}^{x}_{t,m}}\odot\bm{\Delta}^{x}_{t,m}\Big{)}^{\top}&1\end{bmatrix}\in\mathbb{R}^{m\times(2d+1)},\quad{\bm{y}}_{t}\coloneq\begin{bmatrix}\Delta^{f}_{t,1}\\ \Delta^{f}_{t,2}\\ \vdots\\ \Delta^{f}_{t,m}\end{bmatrix}\in\mathbb{R}^{m}.$ (17)
Accordingly, our proposed Quadratic Regression-based Single-point Zeroth-order Optimization (Q-RESZO) method for solving problem (1) follows the same framework as the L-RESZO method (i.e., Algorithm 1), except that it solves for $(\bm{g}_{t},\bm{h}_{t},c_{t})$ via (16) in Step 7 and performs the iterative update by (18) with the gradient estimator $\nabla f^{\mathrm{s2}}_{t}(\bm{x})$ in Step 8.
$\bm{x}_{t+1}=\bm{x}_{t}-\eta\,\nabla f^{\mathrm{s2}}_{t}(\bm{x}_{t})=\bm{x}_{t}-\eta\,\big{(}\bm{g}_{t}-\delta\bm{h}_{t}\odot\bm{u}_{t}\big{)}.$ (18)
Additionally, the RESZO algorithms presented above are designed for solving unconstrained optimization problems as formulated in (1); nevertheless, they can be readily extended to constrained optimization problems using projection techniques [22] and primal-dual algorithms [23].
4 Theoretical Convergence Analysis
This section provides non-asymptotic convergence analysis, and we focus on the L-RESZO method. Throughout the analysis, we utilize the notation $[N]:=\{1,2,\dots,N\}$. We also denote $\xi_{t}:=\bm{g}_{t}-\nabla f(\bm{x}_{t})$ to represent the gradient estimation error at an iteration $t$.
Throughout our analysis, we use the following standard assumption on the smoothness of the function $f$.
Assumption 1.
The function $f$ is $L$-smooth, i.e., $\|\nabla f(\bm{x})\!-\!\nabla f(\bm{y})\|\leq L\|\bm{x}-\bm{y}\|,\,\forall\bm{x},\bm{y}\in\mathbb{R}^{d}.$
For convenience, throughout the analysis, we will study the case where we do not include an intercept $c_{t}$. In this case, the approximate gradient $\bm{g}_{t}$ is the solution to the equation (19):
$\ \bm{g}_{t}=(S_{t}S_{t}^{\top})^{\dagger}S_{t}\Delta_{t}^{f},\quad S_{t}^{\top}:=\begin{bmatrix}(\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t})^{\top}\\ (\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t})^{\top}\\ \vdots\\ (\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t})^{\top}\\ \end{bmatrix},\quad\Delta_{t}^{f}:=\begin{bmatrix}f(\hat{\bm{x}}_{t-1})-f(\hat{\bm{x}}_{t})\\ f(\hat{\bm{x}}_{t-2})-f(\hat{\bm{x}}_{t})\\ \vdots\\ f(\hat{\bm{x}}_{t-m+1})-f(\hat{\bm{x}}_{t})\\ \end{bmatrix},$ (19)
where $S_{t}^{\top}\in\mathbb{R}^{(m-1)\times d},\Delta_{t}^{f}\in\mathbb{R}^{m-1}$.
Inspired by our theoretical analysis, we propose an adaptive $\delta$ scheme, given by (20), to enhance convergence precision:
[…]
where the smoothing radius parameter $\delta_{t}$ at each iteration $t$ is adaptively adjusted based on the norm of the last gradient estimate $\bm{g}_{t-1}$.
In our analysis, we will study the L-RESZO algorithm that employs this adaptive $\delta$ scheme, and numerical experiments on this scheme are performed in Section 5.2. We also make the following assumption on the regression procedure.
Assumption 2.
Suppose $\bm{g}_{t}$ is as defined in (19). We assume that for any $t\geq m$,
$\|\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})\|\leq C_{d}\frac{L}{2}\max_{i=1:m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|,$ (21)
for some constant $C_{d}>0$ depending on the dimension $d$ and smoothing radius sequence $\{\delta_{t}\}_{t}$.
Assumption 2 bounds the error of the approximate gradient by the distance between points used in the regression $\hat{\bm{x}}_{t-i}$ and the current point $\hat{\bm{x}}$; for clarity, we note that the error $\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})$ that appears in (21) is distinct from $\xi_{t}:=\bm{g}_{t}-\nabla f(\bm{x}_{t})$, since the former is the bias compared to the gradient at the perturbed point while the latter is the bias compared to the gradient at the actual iterate of the algorithm. While this assumption may appear unmotivated, we show in Appendix A.1 that it holds for $L$-smooth functions when $d=1$ with $C_{d}=1$. For general cases of $d$, Appendix C presents empirical evidence suggesting that the factor $C_{d}$ in Assumption 2 scales in the order of $\mathcal{O}(\sqrt{d})$ under the adaptive $\delta_{t}$ scheme in (20). While establishing a theoretical bound for $C_{d}$ is desired, it remains very challenging due to the technical difficulties discussed in Appendix A.1.1.
Before deriving our theoretical convergence results, we first specify the following condition on the initial $\tilde{\eta}$, the step-size during the initial update phase (before regression), which states that it should be sufficiently small in a certain sense. Our results later will all require $\tilde{\eta}$ to satisfy this condition.
Condition 1.
Recall that $\tilde{\eta}$ is the step-size of RESZO during the initial update phase (before regression). We pick $\tilde{\eta}$ such that for any $t\in\{0,1,\dots,m-1\}$, it satisfies [^1] We note that this is a relatively mild condition, and that we expect this to hold for any $\tilde{\eta}$ sufficiently small.
$\|\bm{x}_{t+1}-\bm{x}_{t}\|=\|\tilde{\eta}\bm{g}_{t}\|\leq 2\eta\|\nabla f(\bm{x}_{t})\|,$ (22)
where $\bm{g}_{t}$ denotes the estimator of $\nabla f(\bm{x}_{t})$ computed using the RSZO method in (2).
We proceed with introducing our theoretical results. Our analysis is underpinned by the following key technical result, which bounds the gradient approximation error in terms of the norm of the true gradient. To be precise, we show that for any $10<c<$ 1, we can inductively guarantee that the gradient estimation bias $\|\xi_{t}\|$ always satisfies $\|\xi_{t}\|\leq c\|\nabla f(\bm{x}_{t})\|$ for all $t\geq m$.
Lemma 1.
Suppose that Assumptions 1 and 2 hold. Suppose $2\leq m=\mathcal{O}(d)$. Suppose we pick $\tilde{\eta}$ such that it satisfies Condition 1. Then, for any $10<c<$ 1, by picking $\eta=C\frac{1}{dC_{d}L}$ for some sufficiently small absolute constant $C>0$, we can ensure that
$\|\xi_{t}\|=\|\bm{g}_{t}-\nabla f(\bm{x}_{t})\|\leq c\|\nabla f(\bm{x}_{t})\|$
for any $t\geq m$.
We provide a brief proof sketch here, deferring proof details to Appendix A.2. The proof consists of three key steps:
  1. 1.
    First, utilizing Assumption 2 and the adaptive smoothing radius selection, we can bound the gradient estimation error $\xi_{t}:=\bm{g}_{t}-\nabla f(x_{t})$ as
    $\|\xi_{t}\|\leq\mathcal{O}(C_{d}L)\sum_{j=1}^{m}\|\bm{x}_{t-j+1}-\bm{x}_{t-j}\|,$ (23)
    i.e. relate the gradient estimation error to a sum of the increments in the past $m$ updates.
  2. 2.
    Next, via an inductive argument (utilizing our condition on the initial step-size $\tilde{\eta}$), for any $t\geq m$ and $j\in[m]$, we show that we can always bound the one-step increment $\|\bm{x}_{t-j+1}-\bm{x}_{t-j}\|$ as
    $\|\bm{x}_{t-j+1}-\bm{x}_{t-j}\|\leq 2\eta\|\nabla f(\bm{x}_{t-j})\|,$ (24)
    i.e. the size of each update can be bounded in terms of the step-size times the norm of the gradient at that step.
  3. 3.
    Utilizing a simple technical result (Lemma 5 in the Appendix A.2), we can show that whenever (24) holds, with an appropriate choice of $\eta$, the gradient norm changes only by an $O(1)$ multiplicative constant across $m$ iterations. Then, combining this fact with (23) and (24), we have that
    $\|\xi_{t}\|\leq$ $\ \mathcal{O}(C_{d}L)\sum_{j=1}^{m}\|\bm{x}_{t-j+1}-\bm{x}_{t-j}\|$
    $\leq$ $\ \mathcal{O}(C_{d}L)\eta\sum_{j=1}^{m}\|\nabla f(\bm{x}_{t-j})\|$
    $\leq$ $\ \mathcal{O}(C_{d}L)\eta\sum_{j=1}^{m}\|\nabla f(\bm{x}_{t})\|.$
    Using the condition that $m=\mathcal{O}(d)$, it follows then that for any $10<c<$ 1, by selecting $\eta=\frac{C}{dC_{d}L}$ for an appropriately small constant $C>0$, we can ensure that
    $\|\xi_{t}\|\leq c\|\nabla f(\bm{x}_{t})\|,$
    as desired.
With Lemma 1 in hand, we can guarantee the following function value improvement per step.
Lemma 2.
Suppose that Assumptions 1 and 2 hold. Suppose $2\leq m=\mathcal{O}(d)$. Suppose we pick $\tilde{\eta}$ such that it satisfies Condition 1. Then, by picking $\eta=C\frac{1}{dC_{d}L}$ for some sufficiently small absolute constant $C>0$, we can ensure that
$f(\bm{x}_{t+1})-f(\bm{x}_{t})\leq-\frac{\eta}{8}\|\nabla f(\bm{x}_{t})\|^{2}$
for any $t\geq m$.
Proof.
Consider any positive integer $t\geq m$. Fix any $10<c<$ 1. By our assumptions, applying Lemma 1, we find that there exists a sufficiently small absolute constant $C>0$ such that when $\eta=\frac{C}{dC_{d}L}$, we have
$\|\xi_{t}\|\leq c\|\nabla f(\bm{x}_{t})\|$
for any $t\geq m$.
Next, observe that
$\ f(\bm{x}_{t+1})-f(\bm{x}_{t})$
$\stackrel{{\scriptstyle\textnormal{(i)}}}{{\mathstrut{\leq}}}$ $\ -\eta\langle\nabla f(\bm{x}_{t}),\bm{g}_{t}\rangle+\frac{L}{2}\|\eta\bm{g}_{t}\|^{2}$
$=$ $\ -\eta\langle\nabla f(\bm{x}_{t}),\nabla f(\bm{x}_{t})\rangle-\eta\langle\nabla f(\bm{x}_{t}),\bm{g}_{t}-\nabla f(\bm{x}_{t})\rangle+\frac{L}{2}\|\eta(\bm{g}_{t}-\nabla f(\bm{x}_{t})+\nabla f(\bm{x}_{t}))\|^{2}$
$\stackrel{{\scriptstyle\textnormal{(ii)}}}{{\mathstrut{\leq}}}$ $\ -\eta\|\nabla f(\bm{x}_{t})\|^{2}+\frac{\eta}{2}\left(\|\nabla f(\bm{x}_{t})\|^{2}+\|\xi_{t}\|^{2}\right)+\eta^{2}L\left(\|\nabla f(\bm{x}_{t})\|^{2}+\|\xi_{t}\|^{2}\right)$
$\leq$ $\ -\frac{\eta}{4}\|\nabla f(\bm{x}_{t})\|^{2}+\eta\|\xi_{t}\|^{2},$ (25)
where (4) follows by $L$-smoothness of $f$, (4) follows by application of Young’s equality as well as the notation $\xi_{t}:=\bm{g}_{t}-\nabla f(\bm{x}_{t})$, and the final inequality follows by choosing $\eta\leq\frac{1}{4L}$.
Thus, continuing from (25), by picking $c$ to be smaller than $\frac{1}{\sqrt{8}}$, such that
$\|\xi_{t}\|\leq\frac{1}{\sqrt{8}}\|\nabla f(\bm{x}_{t})\|,$
we find that
$f(\bm{x}_{t+1})-f(\bm{x}_{t})\leq$ $\ -\frac{\eta}{4}\|\nabla f(\bm{x}_{t})\|^{2}+\eta\|\xi_{t}\|^{2}.$
$\leq$ $\ -\frac{\eta}{4}\|\nabla f(\bm{x}_{t})\|^{2}+\frac{\eta}{8}\|\nabla f(\bm{x}_{t})\|^{2}$
$\leq$ $\ -\frac{\eta}{8}\|\nabla f(\bm{x}_{t})\|^{2},$
as desired.
Equipped with Lemma 2, we can easily show by standard arguments that the algorithm converges at a sublinear rate for smooth nonconvex functions, and at a linear rate for smooth strongly convex functions. We defer proof details to Appendix A.2.
Theorem 1.
Suppose Assumptions 1 and 2 hold. Suppose $2\leq m=\mathcal{O}(d)$. Suppose we pick $\tilde{\eta}$ such that it satisfies Condition 1. Then, by picking $\eta=C\frac{1}{dC_{d}L}$ for some sufficiently small absolute constant $C>0$, for any $T>m$, the L-RESZO algorithm achieves
$\frac{1}{T}\sum_{t=m}^{T-1}\|\nabla f(\bm{x}_{t})|^{2}\leq\mathcal{O}\left(\frac{dC_{d}L(f(\bm{x}_{m})-f^{*}))}{T}\right).$
This convergence rate scales as $\mathcal{O}(\frac{dC_{d}L}{T})$, decaying at the rate $1/T$. When the term $C_{d}$ is on the order of $\sqrt{d}$ (as observed empirically under the adaptive choice of $\delta_{t}=\eta\|\bm{g}_{t-1}\|$ (20); see Appendix C for more details), this matches the standard convergence rate for two-point ZO of smooth nonconvex functions, up to a factor of $\sqrt{d}$.
Next, we present a linear convergence rate for strongly convex functions.
Theorem 2.
Suppose Assumptions 1 and 2 hold. Suppose $2\leq m=\mathcal{O}(d)$. Suppose we pick $\tilde{\eta}$ such that it satisfies Condition 1. Then, by picking $\eta=C\frac{1}{dC_{d}L}$ for some sufficiently small absolute constant $C>0$, for any $T>m$, the L-RESZO algorithm achieves
$f(\bm{x}_{T})-f^{*}\leq\left(1-\mathcal{O}\left(\frac{\mu}{C_{d}Ld}\right)\right)^{T-m}(f(\bm{x}_{m})-f^{*}).$
Theorem 2 shows that with the adaptive $\delta_{t}$ sequence, the L-RESZO achieves a linear convergence rate that is worse than that of two-point ZO methods for strongly convex functions only by a factor of $C_{d}$. We provide strong empirical evidence in Appendix C that $C_{d}$ scales as $\mathcal{O}(\sqrt{d})$ for both ridge and logistic regression problems. This implies that, for these problems, the theoretical linear convergence rate of L-RESZO is worse than that of two-point ZO only by a factor of $\sqrt{d}$.
5 Numerical Experiments
This section demonstrates the performance of the proposed RESZO methods in comparison with classic ZO methods through numerical experiments [^2] All numerical experiments are coded in Matlab and implemented in a computing environment with Intel(R) Core(TM) i7-1185G7 CPUs running at 3.00 GHz and with 16 GB RAM., and studies the impact of adding perturbations.
5.1 Comparison with Classic ZO Methods
We compare the performance of the proposed L-RESZO (i.e., Algorithm 1), Q-RESZO (18) methods with the RSZO (2) and TZO (4) methods across four optimization tasks, including ridge regression, logistic regression, Rosenbrock function minimization, and neural networks training.
Case (a) Ridge Regression. Consider solving the ridge regression problem (26) [24]:
$\min_{\bm{x}\in\mathbb{R}^{d}}\,f(\bm{x})=\frac{1}{2}||\bm{y}-H\bm{x}||_{2}^{2}+\frac{\lambda}{2}||\bm{x}||_{2}^{2}.$ (26)
In our experiments, each entry of the matrix $H\in\mathbb{R}^{N\times d}$ is independently drawn from the standard Gaussian distribution $\mathcal{N}(0,1)$. The column vector $\bm{y}\in\mathbb{R}^{N}$ is constructed by letting $\bm{y}=\frac{1}{2}H\bm{1}_{d}+\bm{\epsilon}$, where $\bm{1}_{d}$ denotes the $d$-dimensional all-ones vector and $\bm{\epsilon}\sim\mathcal{N}(\bm{0},0.1\bm{I})$ represents the additive Gaussian noise. The regularization parameter is fixed as $\lambda=0.1$, and the initial point is set to $\bm{x}_{0}=\bm{0}_{d}$. We set $N=1000$ and the problem dimension $d=100$.
Case (b) Logistic Regression. Consider solving the logistic regression problem (27) [24]:
$\min_{\bm{x}\in\mathbb{R}^{d}}f(\bm{x})=\frac{1}{2}\sum_{i=1}^{N}\log\big{(}1+\exp(-y_{i}\cdot\bm{s}_{i}^{\top}\bm{x})\big{)}+\frac{\lambda}{2}||\bm{x}||_{2}^{2},$ (27)
where $\bm{s}_{i}\in\mathbb{R}^{d}$ is one of the data samples, $y_{i}\in\{-1,1\}$ is the corresponding class label, and $N=1000$ is the sample size. In our experiments, each element of a data sample $\bm{s}_{i}$ is independently drawn from the uniform distribution $\mathrm{Unif}([-1,1])$, and the corresponding label is computed as $y_{i}=\mathrm{sign}(\frac{1}{2}\bm{s}_{i}^{\top}\mathbf{1}_{d})$. The regularization parameter is fixed as $\lambda=0.1$, and we set $d=100$.
Case (c) Rosenbrock Function. Consider minimizing the nonconvex Rosenbrock function (28):
$\min_{\bm{x}\in\mathbb{R}^{d}}f(\bm{x})=\sum_{i=1}^{d}\Big{[}100\big{(}(x_{i}+1)^{2}-x_{i+1}-1\big{)}^{2}+x_{i}^{2}\Big{]},$ (28)
which is a well-known benchmark problem for numerical optimization [25]. The global optimal solution and objective value are $\bm{x}^{*}\!=\!\bm{0}$ and $f^{*}\!=\!0$. We set $d\!=\!200$ and the initial point $\bm{x}_{0}\!=\!0.5\bm{1}_{d}$.
Case (d) Neural Network Training. Consider training a three-layer fully connected neural network by minimizing the loss function (29):
$\min_{\bm{x}\in\mathbb{R}^{d}}f(\bm{x})=\sum_{i=1}^{N}\Big{[}\bm{w}_{o}^{\top}\sigma(W_{3}\sigma(W_{2}\sigma(W_{1}\bm{s}_{i}+\bm{b}_{1})+\bm{b}_{2})+\bm{b}_{3})-y_{i}\Big{]}^{2},$ (29)
where the decision variable is $\bm{x}\coloneq(W_{1},W_{2},W_{3},\bm{b}_{1},\bm{b}_{2},\bm{b}_{3},\bm{w}_{o})$, and $\sigma(\cdot)$ denotes the Sigmoid activation function. Let $\bm{s}_{i}\in\mathbb{R}^{n}$ and $y_{i}\in\mathbb{R}$ be the input and output of each sample $i$ with $n=6$. Thus, the problem dimension is $d=3n^{2}+4n=132$. We set the sample size to $N=500$. Each sample input $\bm{s}_{i}$ is randomly drawn from the standard Gaussian distribution $\mathcal{N}(0,1)$, and the corresponding output is computed by the neural network with ground truth $\bm{x}^{*}$, where each entry of $\bm{x}^{*}$ is randomly generated from $\mathcal{N}(0,1)$. Let the initial point be $\bm{x}_{0}\!=\!\bm{x}^{*}+\bm{\varepsilon}$, and each entry of $\bm{\varepsilon}$ is randomly drawn from the uniform distribution $\mathrm{Unif}([-1,1])$.
For each ZO method, we optimize the stepsize $\eta$ and the smoothing radius $\delta$ via grid search to achieve the fastest convergence, and the selected parameters are summarized in Table 2. The same sample size $m$ is used for L-RESZO and Q-RESZO, which is set to 110, 110, 210, and 6 in Cases (a)-(d) [^3] In Cases (a)-(c), the sample size is chosen as $m=d+10$ to be slightly larger than the problem dimension $d$. However, in Case (d), a smaller sample size $m$ yields better convergence, and thus a very small sample size $m=6$ is used. It implies that, for certain problems, a small sample size might be preferable and only few recent function evaluations are needed to construct an effective gradient estimate with the proposed RESZO method., respectively. For each of the four optimization tasks described above, we run each ZO method 100 times and calculate the mean and $80\%$-confidence interval (CI) of the optimality gap $f(\bm{x})\!-\!f^{*}$. Figure 1 illustrates the convergence results using the TZO, RSZO, L-RESZO, and Q-RESZO methods, evaluated based on the number of function queries.
Table 2: Selected parameters of smoothing radius $\delta$ and stepsize $\eta$ for the ZO methods.
Methods Ridge Reg. Logistic Reg. Rosenbrock Func. Neural Network
$\boldsymbol{\delta}$ $\boldsymbol{\eta}$ $\boldsymbol{\delta}$ $\boldsymbol{\eta}$ $\boldsymbol{\delta}$ $\boldsymbol{\eta}$ $\boldsymbol{\delta}$ $\boldsymbol{\eta}$
TZO 0.002 $1.1\times 10^{-5}$ 0.01 $1.6\times 10^{-3}$ 0.01 $4.5\times 10^{-6}$ 0.01 $3.8\times 10^{-4}$
RSZO 0.2 $2.5\times 10^{-6}$ 2 $5\times 10^{-4}$ 0.5 $2\times 10^{-6}$ 0.05 $1.1\times 10^{-4}$
L-RESZO 0.002 $8\times 10^{-6}$ 0.1 $2\times 10^{-3}$ 0.02 $4.2\times 10^{-6}$ 0.001 $1.7\times 10^{-3}$
Q-RESZO 0.002 $1.6\times 10^{-5}$ 0.01 $5\times 10^{-3}$ 0.02 $1\times 10^{-5}$ 0.001 $1.7\times 10^{-3}$
As shown in Figure 1, the proposed L-RESZO and Q-RESZO methods outperform the TZO method in terms of function query complexity across all test cases and converge significantly faster than the RSZO method. Regarding iteration complexity, L-RESZO and Q-RESZO achieve a convergence rate comparable to that of TZO, which is consistent with our theoretical analysis. However, since L-RESZO and Q-RESZO require only one function query per iteration, they are approximately twice as fast as TZO with respect to function query complexity. Q-RESZO exhibits slightly faster convergence than L-RESZO, attributable to more accurate gradient estimation enabled by quadratic surrogate functions. Nevertheless, TZO ultimately attains higher convergence precision than L-RESZO and Q-RESZO. As illustrated in the next subsection, the convergence precision of L-RESZO and Q-RESZO is controlled by the smoothing radius $\delta$.
Figure 1: The convergence results of applying the TZO (4), RSZO (2), L-RESZO (Algorithm 1) and Q-RESZO (18) methods to solve ridge regression (26), logistic regression (27), Rosenbrock function minimization (28), and neural network training (29).
5.2 Impact of Random Perturbations on RESZO
Figure 2: Convergence of L-RESZO and Q-RESZO for solving the ridge regression problem (26) under different levels of perturbations and using the adaptive $\delta$ scheme (20).
To analyze the impact of perturbations added for function evaluations, we test the proposed L-RESZO and Q-RESZO methods under varying levels of the smoothing radius $\delta$ when solving the ridge regression problem (26). The experimental settings remain the same as those described in the previous subsection. Specifically, the smoothing radius $\delta$ is progressively reduced from 0.01, 0.005, 0.002, 0.001, to zero (no perturbation), and we also perform the adaptive smoothing radius scheme given by (20). The test results are illustrated in Figure 2. It is observed that a smaller smoothing radius $\delta$ leads to improved convergence precision. That is because the RESZO method essentially estimates the gradient at the perturbed point, i.e., $\nabla f(\hat{\bm{x}})=\nabla f(\bm{x}+\delta\bm{u}_{t})$, and thus a smaller $\delta$ results in lower estimation errors. However, the introduction of random perturbations is necessary for the stability of RESZO. Figure 2 shows that, when no perturbation is added ($\delta=0$), the Q-RESZO method exhibits large oscillations and poor convergence precision, while the L-RESZO method even diverges. Moreover, corroborating our theoretical results, our adaptive $\delta$ scheme given by (20) can effectively enhance convergence precision by adjusting $\delta$ based on the updated gradient estimate.
6 Conclusion
In this paper, we propose a novel yet simple Single-point Zeroth-order Optimization (SZO) framework, termed Regression-Based SZO (RESZO), which significantly improves gradient estimation by leveraging historical function evaluations from previous iterations. We theoretically establish that the RESZO method achieves convergence rates comparable to those of the two-point ZO method for both smooth nonconvex and strongly convex functions. Moreover, extensive numerical experiments demonstrate that RESZO not only matches but achieves approximately twice as fast convergence as the two-point ZO method in terms of function query complexity in all test cases. To the best of our knowledge, this is the first work to present single-point ZO that achieves a comparable convergence rate to two-point ZO and empirically outperforms it.
Limitations. Despite showing comparable convergence rates to two-point ZO, the theoretical analysis in this paper can be further enhanced by (i) rigorously bounding the term $C_{d}$ in Assumption 2 and (ii) improving the convergence rate bounds to exactly match those of two-point ZO, as observed in practice. Additionally, this work focuses exclusively on the noiseless setting; extending the analysis and algorithms to handle noisy function evaluations would better reflect practical scenarios. Future studies can be conducted to explore various strategies to further improve the RESZO framework, such as fitting more advanced surrogate functions, incorporating weighting schemes for historical function evaluations, adopting adaptive stepsize rules, and integrating trust region methods, etc.
Appendix A Proofs for Theoretical Convergence Analysis
A.1 More on Assumption 2
To facilitate our analysis, we first show the following result which provides an alternative characterization of the approximation error term $\|\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})\|$ that appears in Assumption 2.
Lemma 3.
Suppose $\bm{g}_{t}$ is defined as in (19), i.e.
$\ \bm{g}_{t}=(S_{t}S_{t}^{\top})^{\dagger}S_{t}\Delta_{t}^{f},\quad S_{t}^{\top}:=\begin{bmatrix}(\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t})^{\top}\\ (\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t})^{\top}\\ \vdots\\ (\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t})^{\top}\\ \end{bmatrix},\quad\Delta_{t}^{f}:=\begin{bmatrix}f(\hat{\bm{x}}_{t-1})-f(\hat{\bm{x}}_{t})\\ f(\hat{\bm{x}}_{t-2})-f(\hat{\bm{x}}_{t})\\ \vdots\\ f(\hat{\bm{x}}_{t-m+1})-f(\hat{\bm{x}}_{t})\\ \end{bmatrix}.$
Define the errors of the first-order Taylor expansion as:
$\epsilon_{t}:=\begin{bmatrix}f(\hat{\bm{x}}_{t-1})-(f(\hat{\bm{x}}_{t})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t}\rangle)\\ \vdots\\ f(\hat{\bm{x}}_{t-m+1})-(f(\hat{\bm{x}}_{t})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t}\rangle)\end{bmatrix}.$ (30)
Suppose $S_{t}S_{t}^{\top}\in\mathbb{R}^{d\times d}$ is invertible. Then,
$\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})=(S_{t}S_{t}^{\top})^{-1}S_{t}\epsilon_{t}.$
Proof.
Using the assumption that $S_{t}S_{t}^{\top}\in\mathbb{R}^{d\times d}$ is invertible, we have
$\bm{g}_{t}=$ $\ (S_{t}S_{t}^{\top})^{-1}S_{t}\Delta_{t}^{f}=\ (S_{t}S_{t}^{\top})^{-1}S_{t}\left(S_{t}^{\top}\nabla f(\hat{\bm{x}}_{t})+\epsilon_{t}\right)$
$=$ $\ \nabla f(\hat{\bm{x}}_{t})+(S_{t}S_{t}^{\top})^{-1}S_{t}\epsilon_{t},$
as desired. ∎
We first prove the following result (under a very mild condition stated in (31)) showing that Assumption 2 holds with $C_{d}=1$ in the case $d=1$. Define $[N]\coloneq\{1,2,\cdots,N\}$.
Lemma 4.
Suppose Assumption 1 holds. Suppose $d=1$. Suppose also that there exists some $i\in[m-1]$ such that
$|\hat{\bm{x}}_{t,i}-\hat{\bm{x}}|>0.$ (31)
Then, with $\bm{g}_{t}$ being defined as in (19), for any $t\geq m$, Assumption 2 holds with $C_{d}=1$, i.e.
$|\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})|\leq\frac{L}{2}\max_{i=1:m-1}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|.$
Proof.
Note that $d=1$. Since we assume there exists some $i\in[m-1]$ such that $|\hat{\bm{x}}_{t,i}-\hat{\bm{x}}|>0$, we know that $S_{t}S_{t}^{\top}\in\mathbb{R}$ is invertible, satisfying the condition in Lemma 3. Thus, using Lemma 3, we note that the result is equivalent to showing that
$|(S_{t}S_{t}^{\top})^{-1}S_{t}\epsilon_{t}|\leq\frac{L}{2}\max_{i=1:m-1}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|.$
Let $\epsilon_{t,i}$ denote the $i$-th entry of $\epsilon_{t}\in\mathbb{R}^{m-1}$. By $L$-smoothness of $f$, we observe that for any $i\in[m-1]$, we have
$|\epsilon_{t,i}|=\ |f(\hat{\bm{x}}_{t-i})-(f(\hat{\bm{x}}_{t})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\rangle)|\leq\ \frac{L}{2}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|^{2}.$
Using the above result, in the case $d=1$, we observe that
$|(S_{t}S_{t}^{\top})^{-1}S_{t}\epsilon_{t}|=$ $\ \Big{|}\frac{1}{\sum_{i=1}^{m-1}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|^{2}}\sum_{i=1}^{m-1}(\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t})\epsilon_{t,i}\Big{|}$
$\leq$ $\ \frac{1}{\sum_{i=1}^{m-1}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|^{2}}\max_{i=1:m-1}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|\sum_{i=1}^{m-1}|\epsilon_{t,i}|$
$\leq$ $\ \frac{1}{\sum_{i=1}^{m-1}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|^{2}}\max_{i=1:m-1}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|\sum_{i=1}^{m-1}\frac{L}{2}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|^{2}$
$\leq$ $\ \frac{L}{2}\max_{i=1:m-1}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|,$
as desired. ∎
A.1.1 Technical Challenges of Proving Assumption 2 for General Cases with $d>1$
From Lemma 3, suppose that $S_{t}S_{t}^{\top}$ is invertible, proving Assumption 2 is equivalent to showing that
$\|(S_{t}S_{t}^{\top})^{-1}S_{t}\epsilon_{t}\|\leq\frac{C_{d}L}{2}\max_{i=1:m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|,$
for some $C_{d}>0$. A natural approach to proving the above inequality is to use the standard bound
$\|(S_{t}S_{t}^{\top})^{-1}S_{t}\epsilon_{t}\|\leq\|(S_{t}S_{t}^{\top})^{-1}\|_{2}\|S_{t}\epsilon_{t}\|.$
However, it is (i) unclear how we might theoretically bound $\|(S_{t}S_{t}^{\top})^{-1}\|_{2}$, and (ii), empirically, when $d<m=\mathcal{O}(d)$ regression samples are used, on the ridge regression, we observe that $\|(S_{t}S_{t}^{\top})^{-1}\|_{2}$ scales as [^4] We recall that the notation $\Omega(\cdot)$ is defined as follows: for any non-negative variable $n$, saying that a positive function $f(n)$ is $\Omega(g(n))$, where $g(n)$ is some positive funcion of $n$, means that there exists some absolute constant $C_{1}>0$ and a number $n_{0}$ such that for all $n\geq n_{0}$, we have $f(n)\geq C_{1}g(n)>0$.
$\|(S_{t}S_{t}^{\top})^{-1}\|_{2}=\Omega\left(\frac{d^{3}}{\sum_{i=1}^{m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|^{2}}\right).$
Combining this with the argument in Lemma 4 bounding $\|\epsilon_{t,i}\|$ as
$|\epsilon_{t,i}|=\ |f(\hat{\bm{x}}_{t-i})-(f(\hat{\bm{x}}_{t})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\rangle)|\leq\ \frac{L}{2}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|^{2},$
it yields
$\|(S_{t}S_{t}^{\top})^{-1}S_{t}\epsilon_{t}\|\leq$ $\ \|(S_{t}S_{t}^{\top})^{-1}\|_{2}\|S_{t}\epsilon_{t}\|$
$\leq$ $\ \Omega\left(\frac{d^{3}}{\sum_{i=1}^{m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|^{2}}\right)\left\|\sum_{i=1}^{m-1}(\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t})\epsilon_{t,i}\right\|$
$\leq$ $\ \Omega\left(\frac{d^{3}}{\sum_{i=1}^{m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|^{2}}\right)\max_{i=1:m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|\sum_{i=1}^{m-1}\frac{L}{2}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|^{2}$
$\leq$ $\ \Omega\left(d^{3}\right)\frac{L}{2}\max_{i=1:m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|,$
yielding $C_{d}=\Omega(d^{3})$, which as we later demonstrate empirically in Appendix C, vastly overestimates the true size of $C_{d}$ that tends to scale as $\mathcal{O}(\sqrt{d})$. This suggests that the natural approach of estimating $C_{d}$ by first bounding $\|(S_{t}S_{t}^{\top})^{-1}\|_{2}$ does not yield tight bounds; additionally, it is in the first place also non-trivial to theoretically prove an upper bound for $\|(S_{t}S_{t}^{\top})^{-1}\|_{2}$.
Thus, we regard the task of accurately bounding $\|(S_{t}S_{t}^{\top})^{-1}S_{t}\epsilon_{t}\|$ as a very challenging theoretical problem which we leave to future work.
A.2 Proofs for Lemma 1, Theorem 1, and Theorem 2
Throughout the analysis, without loss of generality, we assume that $C_{d}\geq 1$. We also recall what we stated earlier, which is that we will study a variant of the L-RESZO algorithm with an adaptive choice of smoothing radius $\delta_{t}$ as (20),
$\delta_{t}:=\eta\|\bm{g}_{t-1}\|.$
We first prove the following useful helper result, which bounds the change in gradient norm across $m$ iterations whenever the gradient estimation bias can be bounded in terms of the true gradient norm.
Lemma 5.
Suppose Assumption 1 holds. Suppose $2\leq m=\mathcal{O}(d)$. Consider any $t\geq m$. Suppose that
$\|\bm{x}_{j+1}-\bm{x}_{j}\|\leq 2\eta\|\nabla f(\bm{x}_{j})\|$ (32)
for all $j<t$.
Then, by picking $\eta=\mathcal{O}\left(\frac{1}{dL}\right)$, we can ensure that
$\|\nabla f(\bm{x}_{t-j})\|\leq 4\|\nabla f(\bm{x}_{t})\|$
for any $j\in[m]$.
Proof.
Consider any $t\geq m$, and consider any $0\leq j<t$. Since by assumption
$\|\bm{x}_{j+1}-\bm{x}_{j}\|\leq 2\eta\|\nabla f(\bm{x}_{j})\|,$
we have that
[…]
where the first line utilizes $L$-smoothness and the second line utilizes our assumption in (32). Continuing, we observe that
$\|\nabla f(\bm{x}_{j})\|\leq$ $\ \|\nabla f(\bm{x}_{j+1})\|+\|\nabla f(\bm{x}_{j})-\nabla f(\bm{x}_{j+1})\|$
$\leq$ $\|\nabla f(\bm{x}_{j+1})\|+2\eta L\|\nabla f(\bm{x}_{j})\|,$
where the first inequality uses triangle inequality and the second inequality follows from (33). Rearranging, we find that
$\ \left(1-2\eta L\right)\|\nabla f(\bm{x}_{j})\|\leq\|\nabla f(\bm{x}_{j+1})\|$
$\implies$ $\ \|\nabla f(\bm{x}_{j})\|\leq\frac{1}{1-2\eta L}\|\nabla f(\bm{x}_{j+1})\|.$
Suppose we pick $\eta=\frac{1}{2mL}$. Then, continuing, we find that
$\|\nabla f(\bm{x}_{j})\|\leq$ $\ \frac{1}{1-2\eta L}\|\nabla f(\bm{x}_{j+1})\|$
$\leq$ $\ \frac{1}{1-\frac{1}{m}}\|\nabla f(\bm{x}_{j+1})\|$
$=$ $\ \frac{m}{m-1}\|\nabla f(\bm{x}_{j+1})\|.$
It follows then that for any $t\geq m$, for any $j\in[m]$, we have
$\|\nabla f(\bm{x}_{t-j})\|\leq$ $\ \left(\frac{m}{m-1}\right)^{j}\|\nabla f(\bm{x}_{t})\|$
$\leq$ $\ \left(\frac{m}{m-1}\right)^{m}\|\nabla f(\bm{x}_{t})\|$
$\leq$ $\ 4\|\nabla f(\bm{x}_{t})\|,$
where the final inequality follows from the fact that the sequence $(\frac{n}{n-1})^{n}\leq 4$ for all $n\geq 2$. ∎
Next, we restate and prove Lemma 1, which shows that for any $10<c<$ 1, we can inductively guarantee that the gradient estimation bias $\|\xi_{t}\|$ always satisfies $\|\xi_{t}\|\leq c\|\nabla f(\bm{x}_{t})\|$ for all $t\geq m$.
See 1
Proof.
We consider an inductive proof. Our inductive claim is as follows. Pick any $t\geq m$. suppose that the inductive assumption
$\|\bm{x}_{j+1}-\bm{x}_{j}\|\leq 2\eta\|\nabla f(\bm{x}_{j})\|$
holds for all $j<t$. Then, for any $10<c<$ 1, by picking $\eta=C\frac{1}{dC_{d}L}$ for some sufficiently small absolute constant $C>0$, we will show that the inductive claim implies that
$\ \|\xi_{t}\|\leq c\|\nabla f(\bm{x}_{t})\|$
$\implies$ $\ \|\bm{x}_{t+1}-\bm{x}_{t}\|=\|\eta\bm{g}_{t}\|\leq\eta(\|\nabla f(\bm{x}_{t})\|+\|\xi_{t}\|)\leq 2\eta\|\nabla f(\bm{x}_{t})\|,$
which inductively guarantees that $\|\xi_{t}\|\leq c\|\nabla f(\bm{x}_{t})\|$ for all $t^{\prime}\geq t$.
We first observe that our inductive assumption holds for the base case $t=m$, since by our assumption,
$\|\bm{x}_{j+1}-\bm{x}_{j}\|\leq 2\eta\|\nabla f(\bm{x}_{j})\|$
holds for all $j<m$.
Let $\eta_{t}$ denote the step-size at time $t$, which is $\tilde{\eta}$ for any $t<m$ and $\eta$ after that. Observe that for any $t\geq m$, [^5] For notational simplicity, our analysis below assumes that the first $m$ iterations also utilizes the adaptive smoothing scheme. However, this is by no means necessary and we note that with some more tedious derivations, we can easily generalize the derivation to the case with no adaptive smoothing in the first $m$ iterations and achieve the same bound.
$\|\xi_{t}\|\leq$ $\ \|\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})\|+\|\nabla f(\hat{\bm{x}}_{t})-\nabla f(\bm{x}_{t})\|$
$\stackrel{{\scriptstyle\textnormal{(i)}}}{{\mathstrut{\leq}}}$ $\ \|\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})\|+L\|\hat{\bm{x}}_{t}-\bm{x}_{t}\|$
$\stackrel{{\scriptstyle\textnormal{(ii)}}}{{\mathstrut{\leq}}}$ $\ \|\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})\|+\eta_{t-1}L\|\bm{g}_{t-1}\|$
$\stackrel{{\scriptstyle\textnormal{(iii)}}}{{\mathstrut{\leq}}}$ $\ C_{d}\frac{L}{2}\max_{i=1:m-1}\left\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\right\|+\eta_{t-1}L\|\bm{g}_{t-1}\|$
$=$ $\ \frac{C_{d}L}{2}\max_{i=1:m-1}\left\|\sum_{j=1}^{i}(\hat{\bm{x}}_{t-j}-\hat{\bm{x}}_{t-j+1})\right\|+\eta_{t-1}L\|\bm{g}_{t-1}\|$
$=$ $\ \frac{C_{d}L}{2}\max_{i=1:m-1}\left\|\sum_{j=1}^{i}(\bm{x}_{t-j}-\bm{x}_{t-j+1}+\delta_{t-j}\bm{u}_{t-j}-\delta_{t-j+1}\bm{u}_{t-j+1})\right\|+\eta_{t-1}L\|\bm{g}_{t-1}\|$
$=$ $\ \frac{C_{d}L}{2}\max_{i=1:m-1}\left\|\sum_{j=1}^{i}(\bm{x}_{t-j}-\bm{x}_{t-j+1}+\eta_{t-j-1}\|\bm{g}_{t-j-1}\|\bm{u}_{t-j}-\eta_{t-j}\|\bm{g}_{t-j}\|\bm{u}_{t-j+1})\right\|$
$\ \quad\quad+\eta_{t-1}L\|\bm{g}_{t-1}\|$
$\stackrel{{\scriptstyle\textnormal{(iv)}}}{{\mathstrut{\leq}}}$ $\ \frac{C_{d}L}{2}\max_{i=1:m-1}\left(\sum_{j=1}^{i}\left(\eta_{t-j}\left\|\bm{g}_{t-j}\right\|+\eta_{t-j-1}\|\bm{g}_{t-j-1}\|+\eta_{t-j}\|\bm{g}_{t-j}\|\right)\right)+\eta_{t-1}L\|\bm{g}_{t-1}\|$
$\leq$ $\ \frac{C_{d}L}{2}\max_{i=1:m-1}\left(\sum_{j=1}^{i}\left(2\eta_{t-j}\|\bm{g}_{t-j}\|+\eta_{t-j-1}\|\bm{g}_{t-j-1}\|\right)\right)+\eta_{t-1}L\|\bm{g}_{t-1}\|$
$\stackrel{{\scriptstyle\textnormal{(v)}}}{{\mathstrut{\leq}}}$ $\ \frac{C_{d}L}{2}\left(\sum_{j=1}^{m-1}\left(4\eta_{t-j}\|\bm{g}_{t-j}\|+\eta_{t-j-1}\|\bm{g}_{t-j-1}\|\right)\right)$
$\leq$ $\ \frac{C_{d}L}{2}\sum_{j=1}^{m}4\eta_{t-j}\|\bm{g}_{t-j}\|$
=$\stackrel{{\scriptstyle\textnormal{(vi)}}}{{\mathstrut{=}}}$ $\ \frac{C_{d}L}{2}\sum_{j=1}^{m}4\|\bm{x}_{t-j+1}-\bm{x}_{t-j}\|$
$\stackrel{{\scriptstyle\textnormal{(vii)}}}{{\mathstrut{\leq}}}$ $\ 2C_{d}L\sum_{j=1}^{m}2\eta\|\nabla f(\bm{x}_{t-j})\|,$
where (A.2) holds by $L$-smoothness of $f$, (A.2) holds by definition of $\hat{\bm{x}}_{t}:=\hat{\bm{x}}_{t}+\eta_{t-1}\|\bm{g}_{t-1}\|\bm{u}_{t}$ due to the adaptive selection of the smoothing radius $\delta_{t}:=\eta\|\bm{g}_{t-1}\|$ and the fact that each $\bm{u}_{t}$ has unit norm, (A.2) utilizes Assumption 2, (A.2) again uses the fact that each $\bm{u}_{t}$ has a unit norm, (A.2) holds by our assumption at the start of this subsection that $C_{d}\geq 1$, (A.2) holds since $\bm{x}_{t-j+1}=\bm{x}_{t-j}+\eta_{t-j}\bm{g}_{t-j}$, and (A.2) follows by our inductive assumption.
Now, utilizing our inductive assumption, we may also use Lemma 5 to find that for any $j\in[m]$, $\|\nabla f(\bm{x}_{t-j})\|\leq 4\|\nabla f(\bm{x}_{t})\|$. Thus, continuing from above, we find that
$\|\xi_{t}\|\leq$ $\ 2C_{d}L\sum_{j=1}^{m}2\eta\|\nabla f(\bm{x}_{t-j})\|$
$\leq$ $\ 4C_{d}L\sum_{j=1}^{m}4\eta\|\nabla f(\bm{x}_{t})\|$
$\leq$ $\ 16C_{d}L\eta\sum_{j=1}^{m}\|\nabla f(\bm{x}_{t})\|.$
Now, (recalling that $m=\mathcal{O}(d)$), by picking $\eta=\frac{C}{dC_{d}L}$ for a sufficiently small $C>0$, we may conclude that
$\ \|\xi_{t}\|\leq c\|\nabla f(\bm{x}_{t})\|,$
$\implies$ $\ \|\bm{x}_{t+1}-\bm{x}_{t}\|=\|\eta\bm{g}_{t}\|\leq\eta(\|\nabla f(\bm{x}_{t})\|+\|\xi_{t}\|)\leq 2\eta\|\nabla f(\bm{x}_{t})\|.$
Thus, by induction, since the base case of $t=m$ holds, we can conclude that for all $t\geq m$, we have
$\ \|\xi_{t}\|\leq c\|\nabla f(\bm{x}_{t})\|,$
as desired. ∎
We next restate and prove Theorem 1, which provides a convergence result for smooth nonconvex functions.
See 1
Proof.
By Lemma 2, we have that
$f(\bm{x}_{t+1})-f(\bm{x}_{t})\leq-\frac{\eta}{8}\|\nabla f(\bm{x}_{t})\|^{2}.$
Summing this from $t=m$ to $t=T$ and plugging in our choice of $\eta=\mathcal{O}\left(\frac{1}{dC_{d}L}\right)$, we observe that
$\ f(\bm{x}_{T})-f(\bm{x}_{m})\leq-\mathcal{O}\left(\frac{1}{dC_{d}L}\sum_{t=m}^{T-1}\|\nabla f(\bm{x}_{t})\|^{2}\right)$
$\implies$ $\ \mathcal{O}\left(\frac{1}{dC_{d}L}\sum_{t=m}^{T-1}\|\nabla f(\bm{x}_{t})\|^{2}\right)\leq f(\bm{x}_{m})-f(\bm{x}_{T})$
$\implies$ $\ \frac{1}{T}\sum_{t=m}^{T-1}\|\nabla f(\bm{x}_{t})\|^{2}\leq\mathcal{O}\left(\frac{dC_{d}L\left(f(\bm{x}_{m})-f^{*}\right)}{T}\right),$
where the final inequality uses the simple fact that $f(\bm{x}_{t})\geq f^{*}$. ∎
We next restate and prove Theorem 2, which shows a linear convergence rate for smooth and strongly convex functions.
See 2
Proof.
Consider any positive integer $t\geq m$. By Lemma 2, we have that
$f(\bm{x}_{t+1})-f(\bm{x}_{t})\leq$ $\ -\frac{\eta}{8}\|\nabla f(\bm{x}_{t})\|^{2}$
$\leq$ $\ -\frac{\eta\mu}{4}(f(\bm{x}_{t})-f^{*}),$
$\Longrightarrow\ f(\bm{x}_{t+1})-f^{*}$ $\leq\left(1-\frac{\eta\mu}{4}\right)(f(\bm{x}_{t})-f^{*}),$
where the fourth inequality follows by $\mu$-strong convexity which ensures that $\|\nabla f(\bm{x}_{t})\|^{2}\geq 2\mu(f(\bm{x}_{t})-f^{*})$. Our desired result then follows. ∎
Appendix B Empirical Study of $C_{d}$ in Assumption 2
We recall from Assumption 2 that the $C_{d}$ term is defined as a positive ratio term such that
$\|\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})\|\leq C_{d}\frac{L}{2}\max_{i=1:m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|.$
In our empirical calculations below, for simplicity, we calculate this ratio $C_{d}$ at the $t$-th iteration of the algorithm as
$C_{d}(t):=\frac{\|\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})\|}{\frac{L}{2}\|\hat{\bm{x}}_{t-(m-1)}-\hat{\bm{x}}_{t}\|}\geq\frac{\|\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})\|}{\frac{L}{2}\max_{i=1:m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|},$
and use our analytical knowledge of the ridge and logistic regression problems in (26) and (27) to compute the smoothness constant $L$.
B.1 Ridge Regression
We experimented on the ridge regression problem in (26) for three values for the dimension $d$: $d=100,400,900$. We computed the $C_{d}$ ratio statistics for a trial of the algorithm for these three dimensions, and display this information in Table 3, which is based on the data from Figure 3; the corresponding convergence behavior is shown in Figure 4. We observe that the maximum $C_{d}$ scales as around $3\sqrt{d}$, i.e., $\mathcal{O}(\sqrt{d})$.
Table 3: Statistics of the $C_{d}$ ratio across different dimensions $d$ for the ridge regression problem (26).
Dimension $d$ Max 99th Percentile Mean
$100$ $25.1522$ $14.9690$ $3.9955$
$400$ $77.3924$ $53.6369$ $10.4881$
$900$ $104.1276$ $85.5617$ $18.3887$
(a) $d=100$
(b) $d=400$
(c) $d=900$
Figure 3: Empirical distribution of the ratio $C_{d}$ for dimensions $d=100,400,900$ for ridge regression problem in (26). These plots are the sources from which the max/99-percentile/mean in Table 3 are computed.
(a) $d=100$
(b) $d=400$
(c) $d=900$
Figure 4: Convergence of L-RESZO on the ridge-regression problem in (26) for dimensions $d=100,400,900$. The $C_{d}$ ratios shown in Figure 3 are based on the algorithmic trajectories here.
B.2 Logistic Regression
We experimented on the logistic regression problem (27) for three values for the dimension $d$: $d=100,400,900.$ We computed the $C_{d}$ ratio statistics for a trial of the algorithm for these three dimensions, and display this information in Table 4, which is based on the data from Figure 5; the corresponding convergence behavior is shown in Figure 6. We observe that the maximum $C_{d}$ scales as around $2.5\sqrt{d}$, i.e., $\mathcal{O}(\sqrt{d})$.
Table 4: Statistics of the $C_{d}$ ratio with different dimensions $d$ for the logistic regression problem (27).
Dimension $d$ Max 99th Percentile Mean
$100$ $6.9935$ $2.8267$ $0.5665$
$400$ $47.2857$ $24.1032$ $3.7831$
$900$ $75.4320$ $64.0339$ $16.4693$
(a) $d=100$
(b) $d=400$
(c) $d=900$
Figure 5: Empirical distribution of the ratio $C_{d}$ for dimensions $d=100,400,900$ for logistic regression problem in (27). These plots are the sources from which the max/99-percentile/mean in Table 4 were computed.
(a) $d=100$
(b) $d=400$
(c) $d=900$
Figure 6: Convergence of L-RESZO on the logistic regression problem in (27) for dimensions $d=100,400,900$. The $C_{d}$ ratios shown in Figure 5 are based on the algorithmic trajectories here.

### v2

@@
-Regression-Based Single-Point Zeroth-Order Optimization
+Accelerating Single-Point Zeroth-Order Optimization with Regression-Based Gradient Surrogates
 Abstract
@@
 However, SZO suffers from high gradient estimation variance and slow convergence, which severely limit its practical applicability.
-Despite recent advances, the convergence of existing SZO methods remains inferior to that of two-point ZO methods.
-To overcome this limitation, we propose a novel yet simple SZO framework, termed regression-based SZO (RESZO), which substantially enhances the convergence rate.
-Unlike conventional ZO methods that rely solely on function evaluations at the current point for gradient estimation, RESZO improves gradient estimation by effectively leveraging historical function evaluations from previous iterations.
-Specifically, RESZO constructs a surrogate function via regression using recent historical evaluations and employs the gradient of this surrogate function for iterative updates.
-Two variants of RESZO, which fit linear and quadratic surrogate functions respectively, are introduced.
-Theoretically, we provide a non-asymptotic convergence analysis for the linear variant of RESZO, showing that its convergence rates are comparable to those of two-point ZO methods for both smooth nonconvex and strongly convex functions.
-Moreover, extensive numerical experiments demonstrate that RESZO not only matches but empirically outperforms two-point ZO in terms of function query complexity.
+To overcome this limitation, we propose a novel yet simple SZO framework termed regression-based SZO (ReSZO), which substantially enhances the convergence rate.
+Specifically, ReSZO constructs a surrogate function via regression using historical function evaluations and employs the gradient of this surrogate function for iterative updates.
+Two instantiations of ReSZO, which fit linear and quadratic surrogate functions respectively, are introduced.
+Moreover, we provide a non-asymptotic convergence analysis for the linear instantiation of ReSZO, showing that its convergence rates are comparable to those of two-point ZO methods.
+Extensive numerical experiments demonstrate that ReSZO empirically converges two to three times faster than two-point ZO in terms of function query complexity.
 1 Introduction
@@
 As a result, the first-order gradients of function $f$ are unavailable, and problem (1) can be solved solely based on function evaluations.
-This type of black-box optimization problems has attracted considerable recent attention and covers a broad range of applications, including power system control [1, 2], reinforcement learning [3, 4], neural network training [5, 6], sensor management [7], etc.
+This black-box optimization has attracted considerable recent attention and covers a broad range of applications, including power system control [4, 8], reinforcement learning [18, 16], neural network training [2, 3], sensor management [17], etc.
 Zeroth-order optimization (ZO) methods [17, 19] have been extensively developed and applied to solve black-box optimization problems, offering theoretical convergence guarantees.
@@
 Motivation.
-At each iteration $t$, classic ZO methods estimate the gradient $\nabla f(\bm{x}_{t})$ solely based on the function evaluations at the current point $\bm{x}_{t}$, while all historical function evaluations obtained from previous iterations are discarded.
+At each iteration $t$, classic ZO methods estimate the gradient $\nabla f(\bm{x}_{t})$ solely based on the function evaluations at the current iterate $\bm{x}_{t}$, while all historical function evaluations obtained from previous iterations are discarded.
 However, these past function evaluations contain valuable information about the unknown function and can help with the gradient estimations for subsequent iterations, especially those obtained in the most recent iterations.
-Moreover, in many practical applications, such as clinical trials, real-world experiments, and long-term complex simulations, obtaining a single function evaluation can be costly and laborious, making it crucial to fully exploit all available evaluations to enhance efficiency.
+Moreover, in many practical applications, e.g., clinical trials, real-world experiments, and long-horizon simulations, obtaining even a single function evaluation can be costly and laborious, making it crucial to fully exploit all available evaluations to improve efficiency.
 Then, a natural question is
@@
 Contribution.
-Motivated by this question, we propose a novel yet simple single-point ZO framework, termed Regression-based Single-point Zeroth-order Optimization (RESZO).
-In contrast to classic ZO methods that rely solely on function evaluations at the current point for gradient estimation, the proposed RESZO method significantly enhances gradient estimation by leveraging available historical function evaluations from previous iterations.
-Specifically, the RESZO method queries the function value only once at each iteration, and fits a surrogate model based on recent historical function evaluations via regression.
-Then, the surrogate model’s gradient is used as the estimated gradient for iterative updates.
-We introduce two variants of the RESZO method, termed L-RESZO and Q-RESZO, which fit linear and quadratic surrogate functions, respectively.
-A key ingredient of the proposed RESZO method is adding random perturbations to function evaluations to ensure sufficient exploration.
-Theoretically (see Table 1), we show that for both smooth nonconvex and strongly convex functions, the linear variant of our algorithm (L-RESZO) achieves convergence rates that are only worse than those of two-point ZO methods by a factor $C_{d}>0$, where $C_{d}$ is a parameter defined in (21) in Assumption 2.
-This represents a substantial improvement over the rates achieved by existing SZO methods (see Table 1).
-While theoretically bounding the term $C_{d}$ is challenging, strong empirical evidence shows that $C_{d}$ scales as $\mathcal{O}(\sqrt{d})$ under appropriate algorithmic settings; see Section 4 for more details.
-To the best of our knowledge, this is the first work to present single-point ZO methods whose convergence rates match those of two-point ZO methods.
-Moreover, extensive numerical experiments demonstrate that RESZO not only matches but empirically outperforms two-point ZO in terms of function query complexity.
+Motivated by this question, we propose a novel yet simple single-point ZO framework termed Regression-based Single-point Zeroth-order Optimization (ReSZO).
+Unlike classical ZO methods that estimate gradients using only function evaluations at the current iterate, ReSZO exploits recent historical function evaluations to significantly improve gradient estimation while requiring only one function query per iteration.
+Our main contributions are summarized as follows:
+  • 1)
+    (New Framework and Algorithms).
+We introduce the ReSZO framework and develop two concrete instantiations, namely L-ReSZO and Q-ReSZO, which construct linear and quadratic surrogate models, respectively.
+By leveraging historical function evaluations together with an adaptive perturbation scheme, the proposed methods substantially improve convergence efficiency over existing SZO methods and demonstrate strong empirical performance.
+  • 2)
+    (Theoretical Analysis).
+We provide theoretical analysis for the convergence of ReSZO.
+Specifically, for both smooth nonconvex and strongly convex objectives, we show that under certain conditions, L-ReSZO achieves convergence rates that are within a factor $C_{d}\!>\!0$ of those attained by two-point ZO methods (see Table 1).
+While tightly bounding $C_{d}$ is challenging, we offer theoretical justification that $C_{d}$ scales as $\mathcal{O}(d)$.
+We also provide extensive empirical evidence indicating that, under suitable algorithmic settings, the practical scaling follows $C_{d}=\mathcal{O}(\sqrt{d})$ (see Section 4).
+  • 3)
+    (Empirical Evaluation).
+We conduct extensive numerical experiments on a variety of benchmark problems, demonstrating that ReSZO not only matches but empirically converges two to three times faster than two-point ZO in terms of function query complexity.
 Table 1: Iteration complexity of ZO methods for solving Lipschitz and smooth objective functions.
@@
 In addition, our ReSZO method is related to the linear regression and interpolation approaches [9, 1, 25] used in derive-free optimization (DFO) [15].
-The fundamental difference is that these methods [17, 18, 19] perform multiple function evaluations near the current point for linear interpolation or regression.
-In contrast, our approach queries the function value only once at each iteration and fits surrogate models by fully recycling historical function evaluations from past iterations.
-2 Preliminaries on Zeroth-Order Optimization
+The fundamental difference is that these methods [9, 1, 25] perform multiple function evaluations near the current iterate for linear interpolation or regression.
+In contrast, our method queries the function value only once at each iteration and fits surrogate models by recycling historical function evaluations.
+We also note a recent SZO algorithm [14] that achieves strong convergence rates.
+However, unlike our approach, it assumes access to an oracle capable of handling complex-valued queries, which limits its applicability in standard zeroth-order settings where the oracle takes only real-valued queries.
+2 Preliminaries
 The classic single-point ZO (SZO) method [12] that solves (1) is given by (2):
-[…]
-where $\delta\!>\!0$ is a scalar parameter called smoothing radius, $\eta\!>\!0$ is the step size, and $\bm{u}_{t}\in\mathbb{R}^{d}$ is an independent identically distributed random direction vector sampled from $\mathrm{Unif}(\mathbb{S}_{d-1})$, namely the uniform distribution on the unit sphere $\mathbb{S}_{d-1}$, for all $k=0,1,2,\cdots$.
+$\ \textbf{Single-point ZO (SZO)}:\ \bm{x}_{t+1}=\bm{x}_{t}-\eta\cdot\underbrace{\frac{d}{\delta}f(\bm{x}_{t}+\delta\bm{u}_{t})\bm{u}_{t}}_{:=\,\mathsf{G}^{1}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)},\quad t=0,1,2,\cdots$ (2)
+where $\delta\!>\!0$ is a scalar parameter called smoothing radius, $\eta\!>\!0$ is the step size, and $\bm{u}_{t}\in\mathbb{R}^{d}$ is an independent identically distributed random direction vector sampled from $\mathrm{Unif}(\mathbb{S}_{d-1})$, namely the uniform distribution on the unit sphere $\mathbb{S}_{d-1}$, for all $t=0,1,2,\cdots$.
 The random direction vector $\bm{u}_{t}$ can also be sampled from the multivariate Gaussian distribution $\mathcal{N}(\mathbf{0},\bm{I})$ or other zero-mean distributions.
 In (2), $\mathsf{G}^{1}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)$ denotes the SZO gradient estimator of function $f$.
-Recent works [15, 11] have introduced a new variant of SZO, referred to as residual-feedback SZO (RSZO) [15] or high-pass filter-integrated SZO [11], which is given by (2):
-[…]
-$\qquad\bm{x}_{t+1}=\bm{x}_{t}-\eta\cdot\underbrace{\frac{d}{\delta}\Big{(}f(\bm{x}_{t}+\delta\bm{u}_{t})-f(\bm{x}_{t-1}+\delta\bm{u}_{t-1})\Big{)}\bm{u}_{t}}_{\coloneq\,\mathsf{G}^{\mathrm{r}1}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)},\qquad t=1,2,\cdots.$ (3)
-At each iteration $t$, the gradient estimator $\mathsf{G}^{\mathrm{r}1}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)$ of the RSZO method (2) reuses the function evaluation $f(\bm{x}_{t\!-\!1}\!+\!r\bm{u}_{t\!-\!1})$ queried at the last iteration $t\!-\!1$.
+Recent works have introduced a new variant of SZO, referred to as residual-feedback SZO (RSZO) [27] or high-pass filter-integrated SZO [7], which is given by (2):
+$\ \textbf{Residual-feedback SZO (RSZO)}:\text{for }t=1,2,\cdots,$
+$\ \bm{x}_{t+1}=\bm{x}_{t}\!-\!\eta\cdot\underbrace{\frac{d}{\delta}\Big(f(\bm{x}_{t}\!+\!\delta\bm{u}_{t})\!-\!f(\bm{x}_{t-1}\!+\!\delta\bm{u}_{t-1})\Big)\bm{u}_{t}}_{\coloneq\,\mathsf{G}^{\mathrm{r}1}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)}.$ (3)
+At each iteration $t$, the gradient estimator $\mathsf{G}^{\mathrm{r}1}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)$ of the RSZO method (2) reuses the function evaluation $f(\bm{x}_{t\!-\!1}\!+\!r\bm{u}_{t\!-\!1})$ queried at the last iteration t1t\!-\!1.
 This approach substantially reduces gradient estimation variance and improves convergence compared with SZO (2).
 The classic two-point ZO (TZO) method [19] that solves problem (1) is given by (4):
-[…]
-$\qquad\qquad\bm{x}_{t+1}=\bm{x}_{t}-\eta\cdot\underbrace{\frac{d}{2\delta}\Big{(}f(\bm{x}_{t}+\delta\bm{u}_{t})-f(\bm{x}_{t}-\delta\bm{u}_{t})\Big{)}\bm{u}_{t}}_{\coloneq\,\mathsf{G}^{2}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)},\quad t=0,1,2,\cdots,$ (4)
+$\ \textbf{Two-point ZO (TZO)}:\text{ for }t=0,1,2,\cdots$
+$\ \bm{x}_{t+1}=\bm{x}_{t}-\eta\cdot\underbrace{\frac{d}{2\delta}\Big(f(\bm{x}_{t}+\delta\bm{u}_{t})-f(\bm{x}_{t}-\delta\bm{u}_{t})\Big)\bm{u}_{t}}_{\coloneq\,\mathsf{G}^{2}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)},$ (4)
 which employs two function evaluations to estimate the gradient of function $f$ with the two-point gradient estimator $\mathsf{G}^{2}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)$.
 The classic ZO gradient estimators above share the same expectation [19]:
-$\mathbb{E}_{\bm{u}\sim\mathrm{Unif}(\mathbb{S}_{d-1})}\big{[}\mathsf{G}^{1}_{f}(\bm{x};\bm{u},\delta)\big{]}=\,$ $\mathbb{E}_{\bm{u}\sim\mathrm{Unif}(\mathbb{S}_{d-1})}\big{[}\mathsf{G}^{\mathrm{r}1}_{f}(\bm{x};\bm{u},\delta)\big{]}$
-=\,= $\mathbb{E}_{\bm{u}\sim\mathrm{Unif}(\mathbb{S}_{d-1})}\big{[}\mathsf{G}^{2}_{f}(\bm{x};\bm{u},\delta)\big{]}=\nabla\hat{f}(\bm{x};\delta),$ (5)
-where $\hat{f}(\bm{x};\delta):=\mathbb{E}_{\bm{v}\sim\mathrm{Unif}(\mathbb{B}_{d})}\big{[}f(\bm{x}+\delta\bm{v})\big{]}$ is a smoothed version of the original function $f$ [9].
+$\mathbb{E}_{\bm{u}\sim\mathrm{Unif}(\mathbb{S}_{d-1})}\big[\mathsf{G}^{1}_{f}(\bm{x};\bm{u},\delta)\big]$
+$=\,$ $\mathbb{E}_{\bm{u}\sim\mathrm{Unif}(\mathbb{S}_{d-1})}\big[\mathsf{G}^{\mathrm{r}1}_{f}(\bm{x};\bm{u},\delta)\big]$
+$=\,$ $\mathbb{E}_{\bm{u}\sim\mathrm{Unif}(\mathbb{S}_{d-1})}\big[\mathsf{G}^{2}_{f}(\bm{x};\bm{u},\delta)\big]=\nabla\hat{f}(\bm{x};\delta),$ (5)
+where $\hat{f}(\bm{x};\delta):=\mathbb{E}_{\bm{v}\sim\mathrm{Unif}(\mathbb{B}_{d})}\big[f(\bm{x}+\delta\bm{v})\big]$ is a smoothed version of function $f$ [12].
 Despite having the same expectation, the two-point ZO gradient estimator $\mathsf{G}^{2}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)$ exhibits lower variance and achieves faster convergence, compared with the single-point ZO gradient estimators $\mathsf{G}^{1}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)$ and $\mathsf{G}^{\mathrm{r}1}_{f}(\bm{x}_{t};\bm{u}_{t},\delta)$.
@@
 In this paper, we focus on the class of SZO methods that require only one function evaluation per iteration.
-Below, we introduce the proposed RESZO method and two specific variants, L-RESZO and Q-RESZO, which fit linear and quadratic surrogate models using historical function evaluations.
+Below, we introduce the ReSZO method and its two instantiations, L-ReSZO and Q-ReSZO, which fit linear and quadratic surrogate models using historical function evaluations.
+At each iteration $t$, the available function evaluations from all previous iterations are given by $\{f(\hat{\bm{x}}_{k})\}_{k=0}^{t}$ with $\hat{\bm{x}}_{k}=\bm{x}_{k}+\delta\bm{u}_{k}$.
+Here, $\bm{u}_{k}\in\mathbb{R}^{d}$ is a random direction vector sampled from the uniform distribution $\mathrm{Unif}(\mathbb{S}_{d-1})$ over the unit sphere $\mathbb{S}_{d-1}$, and $\delta>0$ is a scalar parameter called smoothing radius.
+These are the same as the classic ZO methods (2)-(4).
+Note that function evaluations are performed at the perturbed points $\hat{\bm{x}}_{k}$ rather than the iterate points ${\bm{x}}_{k}$ to promote exploration of the optimization space.
+The necessity of introducing such perturbations is further illustrated by Remark 1 and the numerical experiments in Section 5.2.
 3.1 Linear Regression-Based Single-Point Zeroth-Order Optimization (L-ReSZO)
-At each iteration $t$, the available function evaluations from all previous iterations are given by $\{f(\hat{\bm{x}}_{k})\}_{k=0}^{t}$ with $\hat{\bm{x}}_{k}=\bm{x}_{k}+\delta\bm{u}_{k}$.
-Here, $\bm{u}_{k}\in\mathbb{R}^{d}$ is a random direction vector sampled from the uniform distribution $\mathrm{Unif}(\mathbb{S}_{d-1})$ over the unit sphere $\mathbb{S}_{d-1}$, or from the multivariate Gaussian distribution $\mathcal{N}(\bm{0},\bm{I})$; and $\delta>0$ is a scalar parameter called smoothing radius.
-These are consistent with the classic ZO methods (2)-(4).
-Note that function evaluations are performed at the perturbed points $\hat{\bm{x}}_{k}$ rather than the iterate points ${\bm{x}}_{k}$ to promote sufficient exploration of the optimization space.
-The necessity of introducing such perturbations is further illustrated by Remark 1 and the numerical experiments in Section 5.2.
-Based on the first-order Taylor expansion at the current perturbed point $\hat{\bm{x}}_{t}$, the available function evaluations $\{f(\hat{\bm{x}}_{k})\}_{k=0}^{t}$ can be expressed as (6):
-$\qquad\qquad f(\hat{\bm{x}}_{k})=f(\hat{\bm{x}}_{t})+\nabla f(\hat{\bm{x}}_{t})^{\top}(\hat{\bm{x}}_{k}-\hat{\bm{x}}_{t})+\epsilon_{t,k},\qquad k=0,1,\cdots,t,$ (6)
-where $\epsilon_{t,k}$ denotes the corresponding high-order residual term.
+Based on the first-order Taylor expansion at the current perturbed point $\hat{\bm{x}}_{t}$, the available function evaluations $\{f(\hat{\bm{x}}_{k})\}_{k=0}^{t}$ can be expressed as (6): for $k=0,1\dots,t$,
+$f(\hat{\bm{x}}_{k})\!=\!f(\hat{\bm{x}}_{t})\!+\!\nabla f(\hat{\bm{x}}_{t})^{\top}\big(\hat{\bm{x}}_{k}\!-\!\hat{\bm{x}}_{t}\big)\!+\!\epsilon_{t,k}(\hat{\bm{x}}_{k}\!-\!\hat{\bm{x}}_{t}),$ (6)
+where the function $\epsilon_{t,k}(\cdot)$ represents the corresponding high-order residual term.
 Accordingly, we propose to fit a linear surrogate function $f^{\mathrm{s1}}_{t}(\bm{x})$ in (7) to locally approximate the original function $f(\bm{x})$ in (1):
-$f^{\mathrm{s1}}_{t}(\bm{x})\coloneq f(\hat{\bm{x}}_{t})+\bm{g}_{t}^{\top}({\bm{x}}-\hat{\bm{x}}_{t})+c_{t},$ (7)
-where the linear coefficient vector $\bm{g}_{t}\in\mathbb{R}^{d}$ and the intercept $c_{t}\in\mathbb{R}$ can be obtained through linear regression on the available sample points $\{\hat{\bm{x}}_{k},f(\hat{\bm{x}}_{k})\}_{k=0}^{t}$.
+$f^{\mathrm{s1}}_{t}(\bm{x})\coloneq f(\hat{\bm{x}}_{t})+\bm{g}_{t}^{\top}\big({\bm{x}}-\hat{\bm{x}}_{t}\big),$ (7)
+where the linear coefficient vector $\bm{g}_{t}\in\mathbb{R}^{d}$ can be obtained through linear regression on the available sample points $\{\hat{\bm{x}}_{k},f(\hat{\bm{x}}_{k})\}_{k=0}^{t}$.
 Then, the gradient of the surrogate function $\nabla f^{\mathrm{s1}}_{t}(\bm{x})=\bm{g}_{t}$ can be used as the estimator for the true gradient $\nabla f(\bm{x})$.
-To ensure that the fitted surrogate model $f^{\mathrm{s1}}_{t}(\bm{x})$ remains up-to-date with the most recent function evaluations, we employ the recursive linear regression in a sliding window, which uses only the latest $m$ function evaluations $\{\hat{\bm{x}}_{t-i+1},f(\hat{\bm{x}}_{t-i+1})\}_{i=1}^{m}$ to fit the linear model at each iteration $t$.
-Specifically, at each iteration $t$, we solve the following least squares problem (8) to obtain $\bm{g}_{t}$ and $c_{t}$:
-$\min_{\bm{g}_{t},c_{t}}\,\sum_{i=1}^{m}\Big{[}f^{\mathrm{s1}}_{t}(\hat{\bm{x}}_{t-i+1})-f(\hat{\bm{x}}_{t-i+1})\Big{]}^{2}\ \Longrightarrow\ \begin{bmatrix}\bm{g}_{t}\\ c_{t}\end{bmatrix}=(X_{t}^{\top}X_{t})^{\dagger}X_{t}^{\top}\bm{y}_{t},$ (8)
-where $\dagger$ denotes the Moore–Penrose pseudoinverse, $\bm{\Delta}^{x}_{t,i}\!\coloneq\!\hat{\bm{x}}_{t-i+1}\!-\!\hat{\bm{x}}_{t},\Delta^{f}_{t,i}\!\coloneq\!f(\hat{\bm{x}}_{t-i+1})\!-\!f(\hat{\bm{x}}_{t})$,
-$X_{t}\coloneq\begin{bmatrix}{\bm{\Delta}^{x}_{t,1}}^{\top}&1\\ {\bm{\Delta}^{x}_{t,2}}^{\top}&1\\ \vdots&\vdots\\ {\bm{\Delta}^{x}_{t,m}}^{\top}&1\end{bmatrix}\in\mathbb{R}^{m\times(d+1)},\quad\bm{y}_{t}\coloneq\begin{bmatrix}\Delta^{f}_{t,1}\\ \Delta^{f}_{t,2}\\ \vdots\\ \Delta^{f}_{t,m}\end{bmatrix}\in\mathbb{R}^{m}.$ (9)
+A key feature of the linear surrogate function (7) is the omission of a constant intercept term, which enforces $f^{\mathrm{s1}}_{t}(\hat{\bm{x}}_{t})=f(\hat{\bm{x}}_{t})$ and reduces model bias.
+To ensure that the fitted surrogate model $f^{\mathrm{s1}}_{t}(\bm{x})$ remains up-to-date with the most recent function evaluations, we employ recursive linear regression in a sliding window, which uses only the latest $m$ function evaluations $\{\hat{\bm{x}}_{t-i+1},f(\hat{\bm{x}}_{t-i+1})\}_{i=1}^{m}$ to fit the linear model at each iteration $t$.
+Specifically, at each iteration $t$, we solve the following least squares problem (8) to obtain $\bm{g}_{t}$:
+$\ \min_{\bm{g}_{t}}\,\sum_{i=1}^{m}\Big[f^{\mathrm{s1}}_{t}(\hat{\bm{x}}_{t-i+1})-f(\hat{\bm{x}}_{t-i+1})\Big]^{2}\Longrightarrow\ \bm{g}_{t}=(X_{t}^{\top}X_{t})^{\dagger}X_{t}^{\top}\bm{y}_{t},$ (8)
+where $\dagger$ denotes the Moore–Penrose pseudoinverse, $\bm{\Delta}^{x}_{t,i}\!\coloneq\!\hat{\bm{x}}_{t-i}\!-\!\hat{\bm{x}}_{t},\Delta^{f}_{t,i}\!\coloneq\!f(\hat{\bm{x}}_{t-i})\!-\!f(\hat{\bm{x}}_{t})$,
+$X_{t}\coloneq\!\!\begin{bmatrix}{\bm{\Delta}^{x}_{t,1}}^{\top}\\ {\bm{\Delta}^{x}_{t,2}}^{\top}\\ \vdots\\ {\bm{\Delta}^{x}_{t,m-1}}^{\top}\end{bmatrix}\!\!\in\mathbb{R}^{(m\!-\!1)\times d},\ \bm{y}_{t}\coloneq\!\!\begin{bmatrix}\Delta^{f}_{t,1}\\ \Delta^{f}_{t,2}\\ \vdots\\ \Delta^{f}_{t,m-1}\end{bmatrix}\!\!\in\mathbb{R}^{m\!-\!1}.$ (9)
 Consequently, the proposed Linear Regression-based Single-point Zeroth-order Optimization (L-ReSZO) method for solving problem (1) is presented as Algorithm 1.
-Since the linear regression requires $m$ function evaluations, we initially employ a classical ZO algorithm, such as the RSZO method (2), to perform iterative updates and accumulate sufficient sample points, as shown in Step 3.
-After $m$ iterations, the algorithm switches to using the linear regression-based gradient estimator $\bm{g}_{t}$ for subsequent iterative updates, as shown in Steps 6-8.
-1:  Inputs: Initial point $\bm{x}_{0}$, stepsizes $\tilde{\eta},\eta$, smoothing radius values $\tilde{\delta},\delta$, sample size $m$, and total iteration number $T$.
-2:  for $t=0:m-1$ do
-3:     Perform the RSZO method (2) with $\tilde{\eta},\tilde{\delta}$ for iterative updates.
+Since the method requires $m$ function evaluations for regression, we initially employ a classical ZO algorithm, e.g., the RSZO method (2), to perform iterative updates and accumulate sufficient sample points, as shown in Steps 2-4.
+After $m$ iterations, the algorithm switches to using the linear regression-based gradient estimator $\bm{g}_{t}$ for subsequent iterative updates, as shown in Steps 5-10.
+The sample size $m$ is typically chosen to be comparable to the dimension $d$ of $\bm{x}$, or slightly larger, to ensure that $X_{t}$ is of full column rank and the matrix $X_{t}^{\top}X_{t}$ in (8) is invertible.
+In this case, one can replace the pseudoinverse $(\cdot)^{\dagger}$ in (8) by the regular matrix inverse $(\cdot)^{-1}$.
+1:Input: Initial point $\bm{x}_{0}$; step sizes $\tilde{\eta},\eta$; smoothing radius $\tilde{\delta}$; sample size $m$; total iterations $T$.
+2:for $t=0$ to m1m-1 do
+3:  Perform RSZO update (i.e., Eq. (2)) with $\tilde{\eta},\tilde{\delta}$.
 4:end for
-5:  for $t=m:T$ do
-6:     Randomly sample $\bm{u}_{t}\sim\mathrm{Unif}(\mathbb{S}_{d-1})$; perform a function evaluation on the perturbed value $\hat{\bm{x}}_{t}=\bm{x}_{t}+\delta\bm{u}_{t}$ to obtain $f(\hat{\bm{x}}_{t})$; collect $m$ latest sample points $\{\hat{\bm{x}}_{t-i+1},f(\hat{\bm{x}}_{t-i+1})\}_{i=1}^{m}$.
-7:     Solve for the linear coefficient $\bm{g}_{t}$ via (8).
-8:     Update with $\bm{x}_{t+1}=\bm{x}_{t}-\eta\cdot\bm{g}_{t}$.
-9:  end for
+5:for $t=m$ to T1T-1 do
+6:  Use an adaptive smoothing radius by setting:
+$\delta_{t}\!=\!\|\bm{x}_{t}\!-\!\bm{x}_{t-1}\|.$ (10)
+7:  Sample $\bm{u}_{t}\sim\mathrm{Unif}(\mathbb{S}_{d-1})$; set $\hat{\bm{x}}_{t}=\bm{x}_{t}+\delta_{t}\bm{u}_{t}$; evaluate $f(\hat{\bm{x}}_{t})$; and collect the latest $m$ samples $\{(\hat{\bm{x}}_{t-i+1},f(\hat{\bm{x}}_{t-i+1}))\}_{i=1}^{m}$.
+8:  Solve for the linear coefficient $\bm{g}_{t}$ via Eq. (8).
+9:  Update $\bm{x}_{t+1}=\bm{x}_{t}-\eta\,\bm{g}_{t}$.
+10:end for
 Algorithm 1 Linear Regression-based Single-point Zeroth-order Optimization (L-ReSZO)
-In addition, the sample size $m$ is typically chosen to be comparable to the dimension $d$ of $\bm{x}$, or slightly larger, to ensure that $X_{t}$ is of full column rank and the matrix $X_{t}^{\top}X_{t}$ in (8) is invertible.
-In this case, one can replace the pseudoinverse $(\cdot)^{\dagger}$ in (8) by the regular matrix inverse $(\cdot)^{-1}$.
+In particular, the following two key approaches are proposed to further improve the performance of Algorithm 1.
+Adaptive Smoothing Radius.
+In Step 6, we propose an adaptive smoothing radius scheme (10) to enhance convergence instead of using a fixed $\delta$.
+The key idea is to adjust the smoothing radius $\delta_{t}$ in proportion to the magnitude of $\|\bm{g}_{t}\|$, since the update in Step 9 follows $\|\bm{x}_{t}\!-\!\bm{x}_{t-1}\|=\eta\|\bm{g}_{t}\|$.
+Using a fixed $\delta$ would cause convergence to stall at an error level on the order of $\mathcal{O}(\delta)$.
+In contrast, the adaptive scheme gradually reduces the perturbation size as the gradient norm decreases, enabling high convergence precision.
+See the proof of Lemma 1 in the Appendix for more details, where this adaptive scheme is used to control the error magnitude of the gradient estimator.
 Fast Computation of Matrix Inverse.
@@
 Hence, the least squares problem (8) based on $\{\bm{\Delta}^{x}_{t,i},\Delta^{f}_{t,i}\}_{i=1}^{m}$ can be replaced by the least squares problem (11) based on $\{\hat{\bm{x}}_{t-i+1},f(\hat{\bm{x}}_{t-i+1})\}_{i=1}^{m}$ to obtain $\bm{g}_{t}$:
-$\min_{\bm{g}_{t},c_{t}}\,\sum_{i=1}^{m}\Big{(}\bm{g}_{t}^{\top}\hat{\bm{x}}_{t-i+1}+c_{t}-f(\hat{\bm{x}}_{t-i+1})\Big{)}^{2}\ \Longrightarrow\ \begin{bmatrix}\bm{g}_{t}\\ c_{t}\end{bmatrix}=(\hat{X}_{t}^{\top}\hat{X}_{t})^{-1}\hat{X}_{t}^{\top}\hat{\bm{y}}_{t},$ (10)
+$\ \min_{\bm{g}_{t}}\,\sum_{i=1}^{m}\Big(\bm{g}_{t}^{\top}\hat{\bm{x}}_{t-i+1}-f(\hat{\bm{x}}_{t-i+1})\Big)^{2}\ \Longrightarrow\ \bm{g}_{t}=(\hat{X}_{t}^{\top}\hat{X}_{t})^{-1}\hat{X}_{t}^{\top}\hat{\bm{y}}_{t},$ (11)
 where
-$\hat{X}_{t}\coloneq\begin{bmatrix}{\hat{\bm{x}}_{t}}^{\top}&1\\ {\hat{\bm{x}}_{t-1}}^{\top}&1\\ \vdots&\vdots\\ {\hat{\bm{x}}_{t-m+1}}^{\top}&1\end{bmatrix}\in\mathbb{R}^{m\times(d+1)},\quad\hat{\bm{y}}_{t}\coloneq\begin{bmatrix}f(\hat{\bm{x}}_{t})\\ f(\hat{\bm{x}}_{t-1})\\ \vdots\\ f(\hat{\bm{x}}_{t-m+1})\end{bmatrix}\in\mathbb{R}^{m}.$ (11)
-Consequently, when moving from iteration $t$ to $t+1$, the matrix $\hat{X}_{t+1}$ is actually formed by dropping the oldest sample point $\hat{\bm{x}}_{t-m+1}$ from $\hat{X}_{t}$ and adding the latest sample point $\hat{\bm{x}}_{t+1}$ to it.
+$\hat{X}_{t}\!\coloneq\!\!\!\begin{bmatrix}{\hat{\bm{x}}_{t-1}}^{\top}\\ {\hat{\bm{x}}_{t-2}}^{\top}\\ \vdots\\ {\hat{\bm{x}}_{t-m+1}}^{\top}\end{bmatrix}\!\!\in\!\mathbb{R}^{(m-\!1)\!\times\!d},\ \hat{\bm{y}}_{t}\!\coloneq\!\!\!\begin{bmatrix}f(\hat{\bm{x}}_{t-1})\\ f(\hat{\bm{x}}_{t-2})\\ \vdots\\ f(\hat{\bm{x}}_{t-m+1})\end{bmatrix}\!\!\in\!\mathbb{R}^{m-1}.$ (12)
+As a result, when moving from iteration $t$ to $t+1$, the matrix $\hat{X}_{t+1}$ is actually formed by dropping the oldest sample point $\hat{\bm{x}}_{t-m+1}$ from $\hat{X}_{t}$ and adding the latest sample point $\hat{\bm{x}}_{t}$ to it.
 This leads to the update rule (13):
-$\hat{X}_{t+1}^{\top}\hat{X}_{t+1}=\hat{X}_{t}^{\top}\hat{X}_{t}-\hat{\bm{x}}_{t-m+1}\hat{\bm{x}}_{t-m+1}^{\top}+\hat{\bm{x}}_{t+1}\hat{\bm{x}}_{t+1}^{\top}.$ (12)
-Hence, given the available matrix inverse $(\hat{X}_{t}^{\top}\hat{X}_{t})^{-1}$ obtained in the last iteration, the new matrix inverse $(\hat{X}_{t+1}^{\top}\hat{X}_{t+1})^{-1}$ can be computed efficiently as (13) by using the Sherman–Morrison–Woodbury formula [21] twice in a row: one for dropping $\hat{\bm{x}}_{t-m+1}$ and the other for adding $\hat{\bm{x}}_{t+1}$;
-$B=A+\frac{A\hat{\bm{x}}_{t-m+1}\hat{\bm{x}}_{t-m+1}^{\top}A}{1-\hat{\bm{x}}_{t-m+1}^{\top}A\hat{\bm{x}}_{t-m+1}},\qquad(\hat{X}_{t+1}^{\top}\hat{X}_{t+1})^{-1}=B-\frac{B\hat{\bm{x}}_{t+1}\hat{\bm{x}}_{t+1}^{\top}B}{1+\hat{\bm{x}}_{t+1}^{\top}B\hat{\bm{x}}_{t+1}},$ (13)
-where $A\coloneq(\hat{X}_{t}^{\top}\hat{X}_{t})^{-1}$ and $B\coloneq(\hat{X}_{t}^{\top}\hat{X}_{t}-\hat{\bm{x}}_{t-m+1}\hat{\bm{x}}_{t-m+1}^{\top})^{-1}$.
+$\hat{X}_{t+1}^{\top}\!\hat{X}_{t+1}\!=\!\hat{X}_{t}^{\top}\!\hat{X}_{t}\!-\!\hat{\bm{x}}_{t-m+1}\hat{\bm{x}}_{t-m+1}^{\top}\!+\!\hat{\bm{x}}_{t}\hat{\bm{x}}_{t}^{\top}.$ (13)
+Given the available matrix inverse $(\hat{X}_{t}^{\top}\hat{X}_{t})^{-1}$ obtained in the last iteration, the new matrix inverse $(\hat{X}_{t+1}^{\top}\hat{X}_{t+1})^{-1}$ can be computed efficiently as (14) by using the Sherman–Morrison–Woodbury formula [11] twice: one for dropping $\hat{\bm{x}}_{t-m+1}$ and the other for adding $\hat{\bm{x}}_{t}$;
+$B=A+\frac{A\hat{\bm{x}}_{t-m+1}\hat{\bm{x}}_{t-m+1}^{\top}A}{1-\hat{\bm{x}}_{t-m+1}^{\top}A\hat{\bm{x}}_{t-m+1}},$ (14a)
+$(\hat{X}_{t+1}^{\top}\hat{X}_{t+1})^{-1}=B-\frac{B\hat{\bm{x}}_{t}\hat{\bm{x}}_{t}^{\top}B}{1+\hat{\bm{x}}_{t}^{\top}B\hat{\bm{x}}_{t}},$ (14b)
+$A\coloneq(\hat{X}_{t}^{\top}\hat{X}_{t})^{-1},\ B\coloneq(\hat{X}_{t}^{\top}\hat{X}_{t}\!-\!\hat{\bm{x}}_{t-m+1}\hat{\bm{x}}_{t-m+1}^{\top})^{-1}.$
 This approach avoids recomputing the new matrix inverse $(\hat{X}_{t+1}^{\top}\hat{X}_{t+1})^{-1}$ from scratch and reduces the computational cost of matrix inversion from $\mathcal{O}(d^{3})$ to $\mathcal{O}(d^{2})$.
@@
 (Role of Perturbation).
-From (11), it is seen that the rows of $\hat{X}_{t}$ are constructed by ${\hat{\bm{x}}_{k}}=\bm{x}_{k}\!+\!\delta\bm{u}_{k}$ for $k=t,\cdots,t\!-\!m\!+\!1$.
-Intuitively, adding the random perturbations $\delta\bm{u}_{k}$ promotes the full row rank property of $\hat{X}_{t}$, thus making $\hat{X}_{t+1}^{\top}\hat{X}_{t+1}$ tend to be invertible.
-This enhances the algorithmic stability of the RESZO method.
-The magnitude of perturbation is controlled by the value of the smoothing radius parameter $\delta$.
-Since the RESZO method essentially estimates the gradient at the perturbed point, i.e., $\nabla f(\hat{\bm{x}})=\nabla f(\bm{x}+\delta\bm{u}_{t})$, a smaller $\delta$ generally results in lower gradient estimation bias and promotes convergence precision.
-Accordingly, rather than selecting a fixed $\delta$, we propose an adaptive $\delta$ scheme for the RESZO algorithm in Section 4.
-The impacts and necessity of incorporating perturbations are further illustrated through numerical experiments in Section 5.2.
+In (12), the rows of $\hat{X}_{t}$ are constructed by ${\hat{\bm{x}}_{k}}=\bm{x}_{k}\!+\!\delta_{k}\bm{u}_{k}$ for $k=t,\cdots,t\!-\!m\!+\!1$.
+Intuitively, adding random perturbations $\delta_{k}\bm{u}_{k}$ promotes the full row rank property of $\hat{X}_{t}$, thus making $\hat{X}_{t+1}^{\top}\hat{X}_{t+1}$ tend to be invertible.
+This is critical for ensuring the algorithmic stability of the ReSZO method.
+The magnitude of perturbations is controlled by the smoothing radius parameter $\delta_{k}$.
+In general, a smaller $\delta$ reduces the bias in gradient estimation and improves convergence precision, whereas a larger $\delta$ promotes greater exploration during gradient estimation.
+The proposed adaptive smoothing radius scheme (10) effectively balances these two aspects based on the current optimization state.
+The necessity of introducing perturbations and the effectiveness of this adaptive scheme are further demonstrated through numerical experiments in Section 5.2.
 3.2 Quadratic Regression-Based Single-Point Zeroth-Order Optimization (Q-ReSZO)
 In addition to fitting a linear model, other more sophisticated surrogate models, such as a quadratic function, can be constructed based on available historical function evaluations to further enhance gradient estimation.
-At each iteration $t$, based on the second-order Taylor expansion at the current perturbed point $\hat{\bm{x}}_{t}$, the available function evaluations $\{f(\hat{\bm{x}}_{k})\}_{k=0}^{t}$ can be expressed as:
-$f(\hat{\bm{x}}_{k})\!=\!f(\hat{\bm{x}}_{t})\!+\!\nabla f(\hat{\bm{x}}_{t})^{\top}(\hat{\bm{x}}_{k}\!-\!\hat{\bm{x}}_{t})\!+\!\frac{1}{2}(\hat{\bm{x}}_{k}\!-\!\hat{\bm{x}}_{t})^{\top}H_{f}(\hat{\bm{x}}_{t})(\hat{\bm{x}}_{k}\!-\!\hat{\bm{x}}_{t})\!+\!\epsilon_{t,k},\ \ k\!=\!0,1,\cdots,t,$ (14)
-where $H_{f}(\cdot)$ denotes the Hessian matrix of function $f$, and $\epsilon_{t,k}$ denotes the corresponding higher-order residual term.
-Accordingly, we can fit a quadratic surrogate function $f^{\mathrm{s2}}_{t}(\bm{x})$ (15) to locally approximate the original function $f(\bm{x})$ in (1):
-$f^{\mathrm{s2}}_{t}(\bm{x})\coloneq f(\hat{\bm{x}}_{t})+\bm{g}_{t}^{\top}({\bm{x}}-\hat{\bm{x}}_{t})+\frac{1}{2}({\bm{x}}-\hat{\bm{x}}_{t})^{\top}\text{diag}(\bm{h}_{t})({\bm{x}}-\hat{\bm{x}}_{t})+c_{t}.$ (15)
+At each iteration $t$, based on the second-order Taylor expansion at the current perturbed point $\hat{\bm{x}}_{t}$, the available function evaluations $\{f(\hat{\bm{x}}_{k})\}_{k=0}^{t}$ can be expressed as: for all $k\!=\!0,1,\cdots,t$,
+$f(\hat{\bm{x}}_{k})=f(\hat{\bm{x}}_{t})\!+\!\nabla f(\hat{\bm{x}}_{t})^{\top}\big(\hat{\bm{x}}_{k}\!-\!\hat{\bm{x}}_{t}\big)+\frac{1}{2}\big(\hat{\bm{x}}_{k}\!-\!\hat{\bm{x}}_{t}\big)^{\top}H_{f}(\hat{\bm{x}}_{t})\big(\hat{\bm{x}}_{k}\!-\!\hat{\bm{x}}_{t}\big)\!+\!\epsilon_{t,k}(\hat{\bm{x}}_{k}\!-\!\hat{\bm{x}}_{t}),$ (15)
+where $H_{f}(\cdot)$ is the Hessian matrix of function $f$, and $\epsilon_{t,k}(\cdot)$ denotes the corresponding higher-order residual term.
+Accordingly, we can fit a quadratic surrogate function $f^{\mathrm{s2}}_{t}(\bm{x})$ (16) to locally approximate the original function $f$ in (1):
+$\ f^{\mathrm{s2}}_{t}(\bm{x})\coloneq f(\hat{\bm{x}}_{t})+\bm{g}_{t}^{\top}\big({\bm{x}}-\hat{\bm{x}}_{t}\big)+\frac{1}{2}\big({\bm{x}}-\hat{\bm{x}}_{t}\big)^{\top}\text{diag}(\bm{h}_{t})\big({\bm{x}}-\hat{\bm{x}}_{t}\big).$ (16)
 Here, we use the diagonal matrix $\text{diag}(\bm{h}_{t})$ with vector $\bm{h}_{t}\in\mathbb{R}^{d}$ to approximate the Hessian matrix, which reduces the number of variables for fitting.
-Similarly, the recursive linear regression in a sliding window is employed, which uses only the latest $m$ function evaluations $\{\hat{\bm{x}}_{t-i+1},f(\hat{\bm{x}}_{t-i+1})\}_{i=1}^{m}$ to fit the quadratic surrogate model at each iteration $t$.
-Then, we solve the following least squares problem (16) to obtain the coefficients $(\bm{g}_{t},\bm{h}_{t},c_{t})$:
-$\min_{\bm{g}_{t},\bm{h}_{t},c_{t}}\,\sum_{i=1}^{m}\Big{[}f^{\mathrm{s2}}_{t}(\hat{\bm{x}}_{t-i+1})-f(\hat{\bm{x}}_{t-i+1})\Big{]}^{2}\Longrightarrow\begin{bmatrix}\bm{g}_{t}\\ \bm{h}_{t}\\ c_{t}\end{bmatrix}=(\tilde{X}_{t}^{\top}\tilde{X}_{t})^{\dagger}\tilde{X}_{t}^{\top}{\bm{y}}_{t},$ (16)
-where $\odot$ denotes the element-wise multiplication of two column vectors, and
-$\tilde{X}_{t}\coloneq\begin{bmatrix}{\bm{\Delta}^{x}_{t,1}}^{\top}&\frac{1}{2}\Big{(}{\bm{\Delta}^{x}_{t,1}}\odot\bm{\Delta}^{x}_{t,1}\Big{)}^{\top}&1\\ {\bm{\Delta}^{x}_{t,2}}^{\top}&\frac{1}{2}\Big{(}{\bm{\Delta}^{x}_{t,2}}\odot\bm{\Delta}^{x}_{t,2}\Big{)}^{\top}&1\\ \vdots&\vdots\\ {\bm{\Delta}^{x}_{t,m}}^{\top}&\frac{1}{2}\Big{(}{\bm{\Delta}^{x}_{t,m}}\odot\bm{\Delta}^{x}_{t,m}\Big{)}^{\top}&1\end{bmatrix}\in\mathbb{R}^{m\times(2d+1)},\quad{\bm{y}}_{t}\coloneq\begin{bmatrix}\Delta^{f}_{t,1}\\ \Delta^{f}_{t,2}\\ \vdots\\ \Delta^{f}_{t,m}\end{bmatrix}\in\mathbb{R}^{m}.$ (17)
-Accordingly, our proposed Quadratic Regression-based Single-point Zeroth-order Optimization (Q-RESZO) method for solving problem (1) follows the same framework as the L-RESZO method (i.e., Algorithm 1), except that it solves for $(\bm{g}_{t},\bm{h}_{t},c_{t})$ via (16) in Step 7 and performs the iterative update by (18) with the gradient estimator $\nabla f^{\mathrm{s2}}_{t}(\bm{x})$ in Step 8.
-$\bm{x}_{t+1}=\bm{x}_{t}-\eta\,\nabla f^{\mathrm{s2}}_{t}(\bm{x}_{t})=\bm{x}_{t}-\eta\,\big{(}\bm{g}_{t}-\delta\bm{h}_{t}\odot\bm{u}_{t}\big{)}.$ (18)
-Additionally, the RESZO algorithms presented above are designed for solving unconstrained optimization problems as formulated in (1); nevertheless, they can be readily extended to constrained optimization problems using projection techniques [22] and primal-dual algorithms [23].
+Similarly, a constant intercept term is omitted from the quadratic surrogate function (16) to enforce $f^{\mathrm{s2}}_{t}(\hat{\bm{x}}_{t})=f(\hat{\bm{x}}_{t})$ and to mitigate model bias.
+In addition, the recursive linear regression in a sliding window is employed, which uses only the latest $m$ function evaluations $\{\hat{\bm{x}}_{t-i+1},f(\hat{\bm{x}}_{t-i+1})\}_{i=1}^{m}$ to fit the quadratic surrogate model (16) at each iteration $t$.
+Then, we solve the following least squares problem (17) to obtain the coefficients $\{\bm{g}_{t},\bm{h}_{t}\}$:
+$\min_{\bm{g}_{t},\bm{h}_{t}}\,\sum_{i=1}^{m}\Big[f^{\mathrm{s2}}_{t}(\hat{\bm{x}}_{t-i+1})-f(\hat{\bm{x}}_{t-i+1})\Big]^{2}\ \Longrightarrow\ \begin{bmatrix}\bm{g}_{t}\\ \bm{h}_{t}\end{bmatrix}=(\tilde{X}_{t}^{\top}\tilde{X}_{t})^{\dagger}\tilde{X}_{t}^{\top}{\bm{y}}_{t},$ (17)
+where $\odot$ denotes the element-wise multiplication of two column vectors, $\tilde{X}_{t}\in\mathbb{R}^{(m-1)\times 2d}$ and ${\bm{y}}_{t}\in\mathbb{R}^{m-1}$ with
+$\tilde{X}_{t}\!\coloneq\!\!\begin{bmatrix}{\bm{\Delta}^{x}_{t,1}}^{\top}&\frac{1}{2}\big({\bm{\Delta}^{x}_{t,1}}\!\odot\!\bm{\Delta}^{x}_{t,1}\big)^{\top}\\ {\bm{\Delta}^{x}_{t,2}}^{\top}&\frac{1}{2}\big({\bm{\Delta}^{x}_{t,2}}\!\odot\!\bm{\Delta}^{x}_{t,2}\big)^{\top}\\ \vdots&\vdots\\ {\bm{\Delta}^{x}_{t,m\!-\!1}}^{\top}&\frac{1}{2}\big({\bm{\Delta}^{x}_{t,m\!-\!1}}\!\odot\!\bm{\Delta}^{x}_{t,m\!-\!1}\big)^{\top}\end{bmatrix}\!,{\bm{y}}_{t}\!\coloneq\!\!\begin{bmatrix}\Delta^{f}_{t,1}\\ \Delta^{f}_{t,2}\\ \vdots\\ \Delta^{f}_{t,m\!-\!1}\end{bmatrix}\!$.
+(18)
+Accordingly, our proposed Quadratic Regression-based Single-point Zeroth-order Optimization (Q-ReSZO) method for solving problem (1) follows the same framework as the L-ReSZO method (i.e., Algorithm 1), except that it solves for $\{\bm{g}_{t},\bm{h}_{t}\}$ via (17) in Step 8 and performs the iterative update by (19) with the gradient estimator $\nabla f^{\mathrm{s2}}_{t}(\bm{x})$ in Step 9.
+$\bm{x}_{t+1}\!=\!\bm{x}_{t}\!-\!\eta\nabla f^{\mathrm{s2}}_{t}(\bm{x}_{t})\!=\!\bm{x}_{t}\!-\!\eta\big(\bm{g}_{t}-\delta_{t}\bm{h}_{t}\odot\bm{u}_{t}\big).$ (19)
+The ReSZO algorithms presented above aim for solving unconstrained optimization problems as formulated in (1); nevertheless, they can be readily extended to constrained optimization problems using projection techniques [6] and primal-dual algorithms [5].
 4 Theoretical Convergence Analysis
-This section provides non-asymptotic convergence analysis, and we focus on the L-RESZO method.
-Throughout the analysis, we utilize the notation $[N]:=\{1,2,\dots,N\}$.
-We also denote $\xi_{t}:=\bm{g}_{t}-\nabla f(\bm{x}_{t})$ to represent the gradient estimation error at an iteration $t$.
-Throughout our analysis, we use the following standard assumption on the smoothness of the function $f$.
+This section presents the non-asymptotic convergence analysis for the L-ReSZO method.
+Throughout the analysis, we use the notation $N]\!\coloneqq\!\{1,2,\dots,N\}$, and $\xi_{t}\!\coloneqq\!\bm{g}_{t}\!-\!\nabla f(\bm{x}_{t})$ denotes the gradient estimation error at iteration $t$.
+Throughout our analysis, we make the following standard assumption on the smoothness of the function $f$.
 Assumption 1.
 The function $f$ is $L$-smooth, i.e., $\|\nabla f(\bm{x})\!-\!\nabla f(\bm{y})\|\leq L\|\bm{x}-\bm{y}\|,\,\forall\bm{x},\bm{y}\in\mathbb{R}^{d}.$
-For convenience, throughout the analysis, we will study the case where we do not include an intercept $c_{t}$. In this case, the approximate gradient $\bm{g}_{t}$ is the solution to the equation (19):
-$\ \bm{g}_{t}=(S_{t}S_{t}^{\top})^{\dagger}S_{t}\Delta_{t}^{f},\quad S_{t}^{\top}:=\begin{bmatrix}(\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t})^{\top}\\ (\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t})^{\top}\\ \vdots\\ (\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t})^{\top}\\ \end{bmatrix},\quad\Delta_{t}^{f}:=\begin{bmatrix}f(\hat{\bm{x}}_{t-1})-f(\hat{\bm{x}}_{t})\\ f(\hat{\bm{x}}_{t-2})-f(\hat{\bm{x}}_{t})\\ \vdots\\ f(\hat{\bm{x}}_{t-m+1})-f(\hat{\bm{x}}_{t})\\ \end{bmatrix},$ (19)
-where $S_{t}^{\top}\in\mathbb{R}^{(m-1)\times d},\Delta_{t}^{f}\in\mathbb{R}^{m-1}$.
-Inspired by our theoretical analysis, we propose an adaptive $\delta$ scheme, given by (20), to enhance convergence precision:
-[…]
-where the smoothing radius parameter $\delta_{t}$ at each iteration $t$ is adaptively adjusted based on the norm of the last gradient estimate $\bm{g}_{t-1}$.
-In our analysis, we will study the L-RESZO algorithm that employs this adaptive $\delta$ scheme, and numerical experiments on this scheme are performed in Section 5.2.
+We recall the approximate gradient $\bm{g}_{t}$ can be written as equation (20):
+$\qquad\qquad\qquad\qquad\bm{g}_{t}=(S_{t}S_{t}^{\top})^{\dagger}S_{t}\Delta_{t}^{f},$ (20)
+$S_{t}^{\top}\!:=\!\!\begin{bmatrix}(\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t})^{\top}\\ (\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t})^{\top}\\ \vdots\\ (\hat{\bm{x}}_{t-m+1}\!-\!\hat{\bm{x}}_{t})^{\top}\\ \end{bmatrix}\!,\Delta_{t}^{f}\!:=\!\!\begin{bmatrix}f(\hat{\bm{x}}_{t-1})\!-\!f(\hat{\bm{x}}_{t})\\ f(\hat{\bm{x}}_{t-2})\!-\!f(\hat{\bm{x}}_{t})\\ \vdots\\ f(\hat{\bm{x}}_{t-m+1})\!-\!f(\hat{\bm{x}}_{t})\\ \end{bmatrix}$,
+where $S_{t}^{\top}=X_{t}\in\mathbb{R}^{(m-1)\times d},\Delta_{t}^{f}=\bm{y}_{t}\in\mathbb{R}^{m-1}$.
 We also make the following assumption on the regression procedure.
 Assumption 2.
-Suppose $\bm{g}_{t}$ is as defined in (19). We assume that for any $t\geq m$,
-$\|\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})\|\leq C_{d}\frac{L}{2}\max_{i=1:m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|,$ (21)
-for some constant $C_{d}>0$ depending on the dimension $d$ and smoothing radius sequence $\{\delta_{t}\}_{t}$.
-Assumption 2 bounds the error of the approximate gradient by the distance between points used in the regression $\hat{\bm{x}}_{t-i}$ and the current point $\hat{\bm{x}}$; for clarity, we note that the error $\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})$ that appears in (21) is distinct from $\xi_{t}:=\bm{g}_{t}-\nabla f(\bm{x}_{t})$, since the former is the bias compared to the gradient at the perturbed point while the latter is the bias compared to the gradient at the actual iterate of the algorithm.
-While this assumption may appear unmotivated, we show in Appendix A.1 that it holds for $L$-smooth functions when $d=1$ with $C_{d}=1$.
-For general cases of $d$, Appendix C presents empirical evidence suggesting that the factor $C_{d}$ in Assumption 2 scales in the order of $\mathcal{O}(\sqrt{d})$ under the adaptive $\delta_{t}$ scheme in (20).
-While establishing a theoretical bound for $C_{d}$ is desired, it remains very challenging due to the technical difficulties discussed in Appendix A.1.1.
-Before deriving our theoretical convergence results, we first specify the following condition on the initial $\tilde{\eta}$, the step-size during the initial update phase (before regression), which states that it should be sufficiently small in a certain sense.
-Our results later will all require $\tilde{\eta}$ to satisfy this condition.
+Suppose that $\bm{g}_{t}$ is as defined in (20), and that $m>d$. We assume that for any $t\geq m$,
+$\|\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})\|\leq C_{d,m}\frac{L}{2}\max_{i=1:m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|,$ (21)
+for some constant $C_{d,m}>0$ depending on the dimension $d$ and $m$. In particular, when the regression sample size is picked as $m=\mathcal{O}(d)$, we denote $C_{d}:=C_{d,m}$.
+Assumption 2 bounds the error of the approximate gradient by the distance between points used in the regression $\hat{\bm{x}}_{t-i}$ and the current point $\hat{\bm{x}}_{t}$.
+For clarity, we note that the error $\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})$ that appears in (21) is distinct from $\xi_{t}:=\bm{g}_{t}-\nabla f(\bm{x}_{t})$, since the former is the bias compared to the gradient at the perturbed point while the latter is the bias compared to the gradient at the actual iterate of the algorithm.
+We note that this assumption is difficult to rigorously prove, since it relies on the intricate behavior of the pseudoinverse of the random matrix $S_{t}S_{t}^{\top}$.
+We provide heuristic reasoning for its validity in Appendix B.1, showing that $C_{d,m}$ scales at worst as $\mathcal{O}(\sqrt{d}{\sqrt{m}})=\mathcal{O}(d)$ (when we set $m=\mathcal{O}(d)$).
+In practice, we set $m=\mathcal{O}(d)$ in our algorithm and present empirical evidence in Appendix C suggesting that in this setting, the factor $C_{d}:=C_{d,m}$ in Assumption 2 actually scales on the order of $\mathcal{O}(\sqrt{d})$.
+Before deriving our theoretical convergence results, we first specify the following condition on the initial $\tilde{\eta}$, namely the stepsize during the initial phase (before regression), indicating that it should be sufficiently small in a certain sense.
+Our subsequent results require that $\tilde{\eta}$ satisfies this condition.
 Condition 1.
 Recall that $\tilde{\eta}$ is the step-size of ReSZO during the initial update phase (before regression).
-We pick $\tilde{\eta}$ such that for any $t\in\{0,1,\dots,m-1\}$, it satisfies [^1] We note that this is a relatively mild condition, and that we expect this to hold for any $\tilde{\eta}$ sufficiently small.
+We pick $\tilde{\eta}$ such that for any $t\in\{0,1,\dots,m-1\}$, it satisfies [^1] We note that this is a relatively mild condition and expect it to hold for any sufficiently small $\tilde{\eta}$.
 $\|\bm{x}_{t+1}-\bm{x}_{t}\|=\|\tilde{\eta}\bm{g}_{t}\|\leq 2\eta\|\nabla f(\bm{x}_{t})\|,$ (22)
 where $\bm{g}_{t}$ denotes the estimator of $\nabla f(\bm{x}_{t})$ computed using the RSZO method in (2).
-We proceed with introducing our theoretical results.
 Our analysis is underpinned by the following key technical result, which bounds the gradient approximation error in terms of the norm of the true gradient.
-To be precise, we show that for any $10<c<$ 1, we can inductively guarantee that the gradient estimation bias $\|\xi_{t}\|$ always satisfies $\|\xi_{t}\|\leq c\|\nabla f(\bm{x}_{t})\|$ for all $t\geq m$.
+To be precise, we show that for any $0<c<1$, we can inductively guarantee that the gradient estimation bias $\|\xi_{t}\|$ always satisfies $\|\xi_{t}\|\leq c\|\nabla f(\bm{x}_{t})\|$ for all $t\geq m$.
 Lemma 1.
-Suppose that Assumptions 1 and 2 hold.
-Suppose $2\leq m=\mathcal{O}(d)$.
-Suppose we pick $\tilde{\eta}$ such that it satisfies Condition 1.
-Then, for any $10<c<$ 1, by picking $\eta=C\frac{1}{dC_{d}L}$ for some sufficiently small absolute constant $C>0$, we can ensure that
-$\|\xi_{t}\|=\|\bm{g}_{t}-\nabla f(\bm{x}_{t})\|\leq c\|\nabla f(\bm{x}_{t})\|$
+Suppose that Assumptions 1 and 2 hold, and that $m>d$.
+Suppose that $\tilde{\eta}$ is picked such that Condition 1 is satisfied.
+Then, for any $0<c<1$, by choosing $\eta=C\frac{1}{mC_{d}L}$ for some sufficiently small absolute constant $C>0$, we have that
+$\|\xi_{t}\|=\|\bm{g}_{t}-\nabla f(\bm{x}_{t})\|\leq c\|\nabla f(\bm{x}_{t})\|,$ (23)
 for any $t\geq m$.
-We provide a brief proof sketch here, deferring proof details to Appendix A.2. The proof consists of three key steps:
-  1. 1.
-    First, utilizing Assumption 2 and the adaptive smoothing radius selection, we can bound the gradient estimation error $\xi_{t}:=\bm{g}_{t}-\nabla f(x_{t})$ as
-    $\|\xi_{t}\|\leq\mathcal{O}(C_{d}L)\sum_{j=1}^{m}\|\bm{x}_{t-j+1}-\bm{x}_{t-j}\|,$ (23)
-    i.e. relate the gradient estimation error to a sum of the increments in the past $m$ updates.
-  2. 2.
-    Next, via an inductive argument (utilizing our condition on the initial step-size $\tilde{\eta}$), for any $t\geq m$ and $j\in[m]$, we show that we can always bound the one-step increment $\|\bm{x}_{t-j+1}-\bm{x}_{t-j}\|$ as
-    $\|\bm{x}_{t-j+1}-\bm{x}_{t-j}\|\leq 2\eta\|\nabla f(\bm{x}_{t-j})\|,$ (24)
-    i.e. the size of each update can be bounded in terms of the step-size times the norm of the gradient at that step.
-  3. 3.
-    Utilizing a simple technical result (Lemma 5 in the Appendix A.2), we can show that whenever (24) holds, with an appropriate choice of $\eta$, the gradient norm changes only by an $O(1)$ multiplicative constant across $m$ iterations.
-Then, combining this fact with (23) and (24), we have that
-    $\|\xi_{t}\|\leq$ $\ \mathcal{O}(C_{d}L)\sum_{j=1}^{m}\|\bm{x}_{t-j+1}-\bm{x}_{t-j}\|$
-    $\leq$ $\ \mathcal{O}(C_{d}L)\eta\sum_{j=1}^{m}\|\nabla f(\bm{x}_{t-j})\|$
-    $\leq$ $\ \mathcal{O}(C_{d}L)\eta\sum_{j=1}^{m}\|\nabla f(\bm{x}_{t})\|.$
-    Using the condition that $m=\mathcal{O}(d)$, it follows then that for any $10<c<$ 1, by selecting $\eta=\frac{C}{dC_{d}L}$ for an appropriately small constant $C>0$, we can ensure that
-    $\|\xi_{t}\|\leq c\|\nabla f(\bm{x}_{t})\|,$
-    as desired.
-With Lemma 1 in hand, we can guarantee the following function value improvement per step.
+We defer the proof to Appendix B.
+With Lemma 1 in hand, we can guarantee the following function value improvement per step; the proof of this result is provided in Appendix B.
 Lemma 2.
-Suppose that Assumptions 1 and 2 hold.
-Suppose $2\leq m=\mathcal{O}(d)$.
-Suppose we pick $\tilde{\eta}$ such that it satisfies Condition 1.
-Then, by picking $\eta=C\frac{1}{dC_{d}L}$ for some sufficiently small absolute constant $C>0$, we can ensure that
-$f(\bm{x}_{t+1})-f(\bm{x}_{t})\leq-\frac{\eta}{8}\|\nabla f(\bm{x}_{t})\|^{2}$
+Suppose that Assumptions 1 and 2 hold, and that $m>d$.
+Suppose that $\tilde{\eta}$ is picked such that Condition 1 is satisfied.
+Then, by choosing $\eta=C\frac{1}{mC_{d}L}$ for some sufficiently small absolute constant $C>0$, we have that
+$f(\bm{x}_{t+1})-f(\bm{x}_{t})\leq-\frac{\eta}{8}\|\nabla f(\bm{x}_{t})\|^{2},$ (24)
 for any $t\geq m$.
-Proof.
-Consider any positive integer $t\geq m$.
-Fix any $10<c<$ 1.
-By our assumptions, applying Lemma 1, we find that there exists a sufficiently small absolute constant $C>0$ such that when $\eta=\frac{C}{dC_{d}L}$, we have
-$\|\xi_{t}\|\leq c\|\nabla f(\bm{x}_{t})\|$
-for any $t\geq m$.
-Next, observe that
-$\ f(\bm{x}_{t+1})-f(\bm{x}_{t})$
-$\stackrel{{\scriptstyle\textnormal{(i)}}}{{\mathstrut{\leq}}}$ $\ -\eta\langle\nabla f(\bm{x}_{t}),\bm{g}_{t}\rangle+\frac{L}{2}\|\eta\bm{g}_{t}\|^{2}$
-$=$ $\ -\eta\langle\nabla f(\bm{x}_{t}),\nabla f(\bm{x}_{t})\rangle-\eta\langle\nabla f(\bm{x}_{t}),\bm{g}_{t}-\nabla f(\bm{x}_{t})\rangle+\frac{L}{2}\|\eta(\bm{g}_{t}-\nabla f(\bm{x}_{t})+\nabla f(\bm{x}_{t}))\|^{2}$
-$\stackrel{{\scriptstyle\textnormal{(ii)}}}{{\mathstrut{\leq}}}$ $\ -\eta\|\nabla f(\bm{x}_{t})\|^{2}+\frac{\eta}{2}\left(\|\nabla f(\bm{x}_{t})\|^{2}+\|\xi_{t}\|^{2}\right)+\eta^{2}L\left(\|\nabla f(\bm{x}_{t})\|^{2}+\|\xi_{t}\|^{2}\right)$
-$\leq$ $\ -\frac{\eta}{4}\|\nabla f(\bm{x}_{t})\|^{2}+\eta\|\xi_{t}\|^{2},$ (25)
-where (4) follows by $L$-smoothness of $f$, (4) follows by application of Young’s equality as well as the notation $\xi_{t}:=\bm{g}_{t}-\nabla f(\bm{x}_{t})$, and the final inequality follows by choosing $\eta\leq\frac{1}{4L}$.
-Thus, continuing from (25), by picking $c$ to be smaller than $\frac{1}{\sqrt{8}}$, such that
-$\|\xi_{t}\|\leq\frac{1}{\sqrt{8}}\|\nabla f(\bm{x}_{t})\|,$
-we find that
-$f(\bm{x}_{t+1})-f(\bm{x}_{t})\leq$ $\ -\frac{\eta}{4}\|\nabla f(\bm{x}_{t})\|^{2}+\eta\|\xi_{t}\|^{2}.$
-$\leq$ $\ -\frac{\eta}{4}\|\nabla f(\bm{x}_{t})\|^{2}+\frac{\eta}{8}\|\nabla f(\bm{x}_{t})\|^{2}$
-$\leq$ $\ -\frac{\eta}{8}\|\nabla f(\bm{x}_{t})\|^{2},$
-as desired.
-Equipped with Lemma 2, we can easily show by standard arguments that the algorithm converges at a sublinear rate for smooth nonconvex functions, and at a linear rate for smooth strongly convex functions.
-We defer proof details to Appendix A.2.
+With Lemma 2 in place, we can show by standard arguments that the L-ReSZO algorithm converges at a sublinear rate for smooth nonconvex functions and at a linear rate for smooth strongly convex functions, as stated in Theorems 1 and 2.
+We defer the detailed proof to Appendix B.
 Theorem 1.
-Suppose Assumptions 1 and 2 hold.
-Suppose $2\leq m=\mathcal{O}(d)$.
-Suppose we pick $\tilde{\eta}$ such that it satisfies Condition 1.
-Then, by picking $\eta=C\frac{1}{dC_{d}L}$ for some sufficiently small absolute constant $C>0$, for any $T>m$, the L-RESZO algorithm achieves
-$\frac{1}{T}\sum_{t=m}^{T-1}\|\nabla f(\bm{x}_{t})|^{2}\leq\mathcal{O}\left(\frac{dC_{d}L(f(\bm{x}_{m})-f^{*}))}{T}\right).$
-This convergence rate scales as $\mathcal{O}(\frac{dC_{d}L}{T})$, decaying at the rate $1/T$.
-When the term $C_{d}$ is on the order of $\sqrt{d}$ (as observed empirically under the adaptive choice of $\delta_{t}=\eta\|\bm{g}_{t-1}\|$ (20); see Appendix C for more details), this matches the standard convergence rate for two-point ZO of smooth nonconvex functions, up to a factor of $\sqrt{d}$.
-Next, we present a linear convergence rate for strongly convex functions.
+Suppose that Assumptions 1 and 2 hold, and that $d<m=\mathcal{O}(d)$.
+Suppose that $\tilde{\eta}$ is picked such that Condition 1 is satisfied.
+Then, by choosing $\eta=C\frac{1}{mC_{d}L}$ for some sufficiently small absolute constant $C>0$, for any $T>m$, the L-ReSZO algorithm achieves
+$\frac{1}{T}\sum_{t=m}^{T-1}\|\nabla f(\bm{x}_{t})||^{2}\leq\mathcal{O}\left(\frac{dC_{d}L(f(\bm{x}_{m})-f^{*})}{T}\right).$ (25)
+This convergence rate scales as $\mathcal{O}(\frac{dC_{d}L}{T})$ when we pick $m=\Theta(d)$, decaying at the rate $1/T$.
+We briefly comment on picking $m$ here.
+Our theory suggests that $m$ should be larger than $d$ to ensure that the matrix $S_{t}$ is full-rank, a requirement for our analysis of Assumption 2); while it should not exceed $\Theta(d)$, since $\eta$ scales as $\mathcal{O}(1/m)$ and increasing $m$ slows down the convergence rate.
+Next, we present the linear convergence result for strongly convex functions.
 Theorem 2.
-Suppose Assumptions 1 and 2 hold.
-Suppose $2\leq m=\mathcal{O}(d)$.
-Suppose we pick $\tilde{\eta}$ such that it satisfies Condition 1.
-Then, by picking $\eta=C\frac{1}{dC_{d}L}$ for some sufficiently small absolute constant $C>0$, for any $T>m$, the L-RESZO algorithm achieves
-$f(\bm{x}_{T})-f^{*}\leq\left(1-\mathcal{O}\left(\frac{\mu}{C_{d}Ld}\right)\right)^{T-m}(f(\bm{x}_{m})-f^{*}).$
-Theorem 2 shows that with the adaptive $\delta_{t}$ sequence, the L-RESZO achieves a linear convergence rate that is worse than that of two-point ZO methods for strongly convex functions only by a factor of $C_{d}$.
-We provide strong empirical evidence in Appendix C that $C_{d}$ scales as $\mathcal{O}(\sqrt{d})$ for both ridge and logistic regression problems.
-This implies that, for these problems, the theoretical linear convergence rate of L-RESZO is worse than that of two-point ZO only by a factor of $\sqrt{d}$.
+Suppose that Assumptions 1 and 2 hold, and that $d<m=\mathcal{O}(d)$.
+Suppose that $\tilde{\eta}$ is picked such that Condition 1 is satisfied.
+Then, by choosing $\eta=C\frac{1}{mC_{d}L}$ for some sufficiently small absolute constant $C>0$, for any $T>m$, the L-ReSZO algorithm achieves
+$f(\bm{x}_{T})-f^{*}\leq\left(1-\mathcal{O}\big(\frac{\mu}{C_{d}Ld}\big)\right)^{T-m}(f(\bm{x}_{m})-f^{*}).$ (26)
+Theorem 2 shows that the L-ReSZO algorithm achieves a linear convergence rate for strongly convex functions, which is slower than that of two-point ZO methods only by a factor of $C_{d}$.
 5 Numerical Experiments
-This section demonstrates the performance of the proposed RESZO methods in comparison with classic ZO methods through numerical experiments [^2] All numerical experiments are coded in Matlab and implemented in a computing environment with Intel(R) Core(TM) i7-1185G7 CPUs running at 3.00 GHz and with 16 GB RAM., and studies the impact of adding perturbations.
+This section demonstrates the performance of the proposed ReSZO methods in comparison with classic ZO methods through numerical experiments across various optimization problems, and studies the impact of adding perturbations.
 5.1 Comparison with Classic ZO Methods
@@
 Case (a) Ridge Regression. Consider solving the ridge regression problem (27) [24]:
-$\min_{\bm{x}\in\mathbb{R}^{d}}\,f(\bm{x})=\frac{1}{2}||\bm{y}-H\bm{x}||_{2}^{2}+\frac{\lambda}{2}||\bm{x}||_{2}^{2}.$ (26)
+[…]
 In our experiments, each entry of the matrix $H\in\mathbb{R}^{N\times d}$ is independently drawn from the standard Gaussian distribution $\mathcal{N}(0,1)$.
@@
 The regularization parameter is fixed as $\lambda=0.1$, and the initial point is set to $\bm{x}_{0}=\bm{0}_{d}$.
-We set $N=1000$ and the problem dimension $d=100$.
+We set $N=1500$ and the problem dimension $d=500$.
 Case (b) Logistic Regression. Consider solving the logistic regression problem (28) [24]:
-$\min_{\bm{x}\in\mathbb{R}^{d}}f(\bm{x})=\frac{1}{2}\sum_{i=1}^{N}\log\big{(}1+\exp(-y_{i}\cdot\bm{s}_{i}^{\top}\bm{x})\big{)}+\frac{\lambda}{2}||\bm{x}||_{2}^{2},$ (27)
-where $\bm{s}_{i}\in\mathbb{R}^{d}$ is one of the data samples, $y_{i}\in\{-1,1\}$ is the corresponding class label, and $N=1000$ is the sample size.
-In our experiments, each element of a data sample $\bm{s}_{i}$ is independently drawn from the uniform distribution $\mathrm{Unif}([-1,1])$, and the corresponding label is computed as $y_{i}=\mathrm{sign}(\frac{1}{2}\bm{s}_{i}^{\top}\mathbf{1}_{d})$.
-The regularization parameter is fixed as $\lambda=0.1$, and we set $d=100$.
+[…]
+where $\bm{s}_{i}\in\mathbb{R}^{d}$ is one of the data samples, and $y_{i}\in\{-1,1\}$ is the corresponding class label.
+Each element of a data sample $\bm{s}_{i}$ is independently drawn from the uniform distribution $\mathrm{Unif}([-1,1])$, and the corresponding label is computed as $y_{i}=\mathrm{sign}(\frac{1}{2}\bm{s}_{i}^{\top}\mathbf{1}_{d})$.
+The regularization parameter is fixed as $\lambda=0.1$.
+We set $N=1000$ and $d=100$.
 Case (c) Rosenbrock Function. Consider minimizing the nonconvex Rosenbrock function (29):
@@
 The global optimal solution and objective value are $\bm{x}^{*}\!=\!\bm{0}$ and $f^{*}\!=\!0$.
-We set $d\!=\!200$ and the initial point $\bm{x}_{0}\!=\!0.5\bm{1}_{d}$.
+We set d=200d\!=\!200 and the initial point $\bm{x}_{0}\!=\!0.5\bm{1}_{d}$.
 Case (d) Neural Network Training. Consider training a three-layer fully connected neural network by minimizing the loss function (30):
 $\min_{\bm{x}\in\mathbb{R}^{d}}f(\bm{x})=\sum_{i=1}^{N}\Big[\bm{w}_{o}^{\top}\sigma(W_{3}\sigma(W_{2}\sigma(W_{1}\bm{s}_{i}+\bm{b}_{1})+\bm{b}_{2})+\bm{b}_{3})-y_{i}\Big]^{2},$ (30)
-where the decision variable is $\bm{x}\coloneq(W_{1},W_{2},W_{3},\bm{b}_{1},\bm{b}_{2},\bm{b}_{3},\bm{w}_{o})$, and $\sigma(\cdot)$ denotes the Sigmoid activation function.
+where $\bm{x}\coloneq(W_{1},W_{2},W_{3},\bm{b}_{1},\bm{b}_{2},\bm{b}_{3},\bm{w}_{o})$, and $\sigma(\cdot)$ denotes the Sigmoid activation function.
 Let $\bm{s}_{i}\in\mathbb{R}^{n}$ and $y_{i}\in\mathbb{R}$ be the input and output of each sample $i$ with $n=6$.
@@
 Let the initial point be $\bm{x}_{0}\!=\!\bm{x}^{*}+\bm{\varepsilon}$, and each entry of $\bm{\varepsilon}$ is randomly drawn from the uniform distribution $\mathrm{Unif}([-1,1])$.
-For each ZO method, we optimize the stepsize $\eta$ and the smoothing radius $\delta$ via grid search to achieve the fastest convergence, and the selected parameters are summarized in Table 2.
-The same sample size $m$ is used for L-RESZO and Q-RESZO, which is set to 110, 110, 210, and 6 in Cases (a)-(d) [^3] In Cases (a)-(c), the sample size is chosen as $m=d+10$ to be slightly larger than the problem dimension $d$.
-However, in Case (d), a smaller sample size $m$ yields better convergence, and thus a very small sample size $m=6$ is used.
-It implies that, for certain problems, a small sample size might be preferable and only few recent function evaluations are needed to construct an effective gradient estimate with the proposed RESZO method., respectively.
-For each of the four optimization tasks described above, we run each ZO method 100 times and calculate the mean and $80\%$-confidence interval (CI) of the optimality gap $f(\bm{x})\!-\!f^{*}$.
+We optimize the stepsize $\eta$ for each ZO method and the smoothing radius $\delta$ for RSZO and TZO via grid search to achieve the fastest convergence.
+The adaptive smoothing radius scheme (10) is employed in L-ReSZO and Q-ReSZO.
+The selected parameters of $\eta$ and $\delta$, as well as the sample size $m$ for regression, are presented in Table 2 in Appendix A.
+For each of the four optimization tasks described above, we run each ZO method 50 times and calculate the mean and $80\%$-confidence interval (CI) of the optimality gap $f(\bm{x})\!-\!f^{*}$.
 Figure 1 illustrates the convergence results using the TZO, RSZO, L-ReSZO, and Q-ReSZO methods, evaluated based on the number of function queries.
-Table 2: Selected parameters of smoothing radius $\delta$ and stepsize $\eta$ for the ZO methods.
+Figure 1: The convergence and 80% confidence intervals (CI) of TZO (4), RSZO (2), L-ReSZO (Algorithm 1) and Q-ReSZO (19) for solving ridge regression (27), logistic regression (28), Rosenbrock function minimization (29), and neural network training (30).
+As shown in Figure 1, the proposed L-ReSZO and Q-ReSZO methods converge two to three times faster than TZO in terms of function query complexity and significantly outperform RSZO across all test cases.
+With respect to iteration complexity, L-ReSZO and Q-ReSZO achieve convergence rates comparable to, or slightly faster than, TZO.
+Since L-ReSZO and Q-ReSZO require only one function query per iteration, they are approximately two to three times more efficient than TZO in terms of function query complexity.
+In addition, Q-ReSZO exhibits slightly faster convergence than L-ReSZO, which can be attributed to its more accurate gradient estimation enabled by quadratic surrogate functions.
+5.2 Effect of Perturbations on ReSZO
+To analyze the effect of perturbations used in function evaluations, we evaluate the proposed L-ReSZO and Q-ReSZO methods under varying levels of the smoothing radius $\delta$ when solving the ridge regression problem (27).
+The experimental setup remains the same as that described in the previous subsection.
+Specifically, $\delta$ is progressively reduced from 1, 0.1, 0.01, 0.001, to zero (no perturbation), and we also perform the adaptive smoothing radius scheme (10).
+The test results are illustrated in Figure 2.
+It is observed that a smaller smoothing radius $\delta$ generally leads to improved convergence precision.
+This is because ZO methods estimate the gradient at a perturbed point, i.e., $\nabla f(\hat{\bm{x}})=\nabla f(\bm{x}+\delta\bm{u}_{t})$, and thus a smaller $\delta$ results in lower gradient estimation errors.
+However, the introduction of random perturbations is necessary for the stability of ReSZO.
+As shown in Figure 2, when no perturbation is applied ($\delta=0$) or when the perturbation is excessively small ($\delta=0.001$), the L-ReSZO method diverges, and the Q-ReSZO method exhibits large oscillations in the former case.
+The proposed adaptive smoothing radius scheme effectively addresses these issues by automatically adjusting $\delta$, achieving both algorithmic stability and high convergence precision.
+Figure 2: Convergence of L-ReSZO and Q-ReSZO for solving the ridge regression problem (27) under different levels of perturbations and using the adaptive smoothing radius scheme (10).
+6 Conclusion
+This paper proposes a novel yet simple single-point zeroth-order optimization (SZO) framework termed Regression-based SZO (ReSZO), which significantly improves gradient estimation by leveraging historical function evaluations from previous iterations.
+We theoretically establish that the ReSZO method achieves convergence rates comparable to those of the two-point ZO method for both smooth nonconvex and strongly convex functions.
+Furthermore, extensive numerical experiments demonstrate that ReSZO not only matches but empirically converges approximately two to three times faster than two-point ZO methods in terms of function query complexity across all test cases.
+To our knowledge, this is the first single-point zeroth-order method that attains convergence rates comparable to two-point methods while achieving superior empirical performance.
+Future work will explore strategies to further enhance the ReSZO framework, including more expressive surrogate models, weighted regression schemes, and extensions to noisy function evaluations that better reflect practical settings.
+Appendix A Numerical Experiment Settings
+For the numerical experiments in Section 5.1, we optimize the stepsize $\eta$ for each ZO method and the smoothing radius $\delta$ for RSZO and TZO via grid search to achieve the fastest convergence.
+The adaptive smoothing radius scheme (10) is employed in L-ReSZO and Q-ReSZO.
+The selected parameters of $\eta$ and $\delta$, as well as the sample size $m$ for regression, are presented in Table 2.
+The same sample size $m$ is used for both L-ReSZO and Q-ReSZO, with $m$ set to 510, 110, 210, and 6 for Cases (a)–(d), respectively.
+In Cases (a)–(c), the sample size is chosen as $m=d+10$, slightly exceeding the problem dimension $d$.
+In Case (d), a smaller sample size yields better convergence, and thus $m=6$ is adopted.
+This observation suggests that, for certain problems, only a small number of recent function evaluations is sufficient to construct an accurate gradient estimate using the proposed ReSZO method.
+Table 2: The selected smoothing radius $\delta$, stepsize $\eta$, and sample size $m$ for the ZO methods in Cases (a)-(d) in Section 5.1.
 Methods Ridge Reg. Logistic Reg. Rosenbrock Func. Neural Network
-$\boldsymbol{\delta}$ $\boldsymbol{\eta}$ $\boldsymbol{\delta}$ $\boldsymbol{\eta}$ $\boldsymbol{\delta}$ $\boldsymbol{\eta}$ $\boldsymbol{\delta}$ $\boldsymbol{\eta}$
-TZO 0.002 $1.1\times 10^{-5}$ 0.01 $1.6\times 10^{-3}$ 0.01 $4.5\times 10^{-6}$ 0.01 $3.8\times 10^{-4}$
-RSZO 0.2 $2.5\times 10^{-6}$ 2 $5\times 10^{-4}$ 0.5 $2\times 10^{-6}$ 0.05 $1.1\times 10^{-4}$
-L-RESZO 0.002 $8\times 10^{-6}$ 0.1 $2\times 10^{-3}$ 0.02 $4.2\times 10^{-6}$ 0.001 $1.7\times 10^{-3}$
-Q-RESZO 0.002 $1.6\times 10^{-5}$ 0.01 $5\times 10^{-3}$ 0.02 $1\times 10^{-5}$ 0.001 $1.7\times 10^{-3}$
-As shown in Figure 1, the proposed L-RESZO and Q-RESZO methods outperform the TZO method in terms of function query complexity across all test cases and converge significantly faster than the RSZO method.
-Regarding iteration complexity, L-RESZO and Q-RESZO achieve a convergence rate comparable to that of TZO, which is consistent with our theoretical analysis.
-However, since L-RESZO and Q-RESZO require only one function query per iteration, they are approximately twice as fast as TZO with respect to function query complexity.
-Q-RESZO exhibits slightly faster convergence than L-RESZO, attributable to more accurate gradient estimation enabled by quadratic surrogate functions.
-Nevertheless, TZO ultimately attains higher convergence precision than L-RESZO and Q-RESZO.
-As illustrated in the next subsection, the convergence precision of L-RESZO and Q-RESZO is controlled by the smoothing radius $\delta$.
-Figure 1: The convergence results of applying the TZO (4), RSZO (2), L-RESZO (Algorithm 1) and Q-RESZO (18) methods to solve ridge regression (26), logistic regression (27), Rosenbrock function minimization (28), and neural network training (29).
-5.2 Impact of Random Perturbations on RESZO
-Figure 2: Convergence of L-RESZO and Q-RESZO for solving the ridge regression problem (26) under different levels of perturbations and using the adaptive $\delta$ scheme (20).
-To analyze the impact of perturbations added for function evaluations, we test the proposed L-RESZO and Q-RESZO methods under varying levels of the smoothing radius $\delta$ when solving the ridge regression problem (26).
-The experimental settings remain the same as those described in the previous subsection.
-Specifically, the smoothing radius $\delta$ is progressively reduced from 0.01, 0.005, 0.002, 0.001, to zero (no perturbation), and we also perform the adaptive smoothing radius scheme given by (20).
-The test results are illustrated in Figure 2.
-It is observed that a smaller smoothing radius $\delta$ leads to improved convergence precision.
-That is because the RESZO method essentially estimates the gradient at the perturbed point, i.e., $\nabla f(\hat{\bm{x}})=\nabla f(\bm{x}+\delta\bm{u}_{t})$, and thus a smaller $\delta$ results in lower estimation errors.
-However, the introduction of random perturbations is necessary for the stability of RESZO.
-Figure 2 shows that, when no perturbation is added ($\delta=0$), the Q-RESZO method exhibits large oscillations and poor convergence precision, while the L-RESZO method even diverges.
-Moreover, corroborating our theoretical results, our adaptive $\delta$ scheme given by (20) can effectively enhance convergence precision by adjusting $\delta$ based on the updated gradient estimate.
-6 Conclusion
-In this paper, we propose a novel yet simple Single-point Zeroth-order Optimization (SZO) framework, termed Regression-Based SZO (RESZO), which significantly improves gradient estimation by leveraging historical function evaluations from previous iterations.
-We theoretically establish that the RESZO method achieves convergence rates comparable to those of the two-point ZO method for both smooth nonconvex and strongly convex functions.
-Moreover, extensive numerical experiments demonstrate that RESZO not only matches but achieves approximately twice as fast convergence as the two-point ZO method in terms of function query complexity in all test cases.
-To the best of our knowledge, this is the first work to present single-point ZO that achieves a comparable convergence rate to two-point ZO and empirically outperforms it.
-Limitations.
-Despite showing comparable convergence rates to two-point ZO, the theoretical analysis in this paper can be further enhanced by (i) rigorously bounding the term $C_{d}$ in Assumption 2 and (ii) improving the convergence rate bounds to exactly match those of two-point ZO, as observed in practice.
-Additionally, this work focuses exclusively on the noiseless setting; extending the analysis and algorithms to handle noisy function evaluations would better reflect practical scenarios.
-Future studies can be conducted to explore various strategies to further improve the RESZO framework, such as fitting more advanced surrogate functions, incorporating weighting schemes for historical function evaluations, adopting adaptive stepsize rules, and integrating trust region methods, etc.
-Appendix A Proofs for Theoretical Convergence Analysis
-A.1 More on Assumption 2
+$\boldsymbol{\delta}$ $\boldsymbol{\eta}$ $\boldsymbol{m}$ $\boldsymbol{\delta}$ $\boldsymbol{\eta}$ $\boldsymbol{m}$ $\boldsymbol{\delta}$ $\boldsymbol{\eta}$ $\boldsymbol{m}$ $\boldsymbol{\delta}$ $\boldsymbol{\eta}$ $\boldsymbol{m}$
+TZO 0.01 $1.2\times 10^{-6}$ 0.01 $1.6\times 10^{-3}$ 0.01 $4.5\times 10^{-6}$ 0.01 $3.8\times 10^{-4}$
+RSZO 0.2 $3\times 10^{-7}$ 2 $5\times 10^{-4}$ 0.5 $2\times 10^{-6}$ 0.05 $1.1\times 10^{-4}$
+L-ReSZO adaptive $1.5\times 10^{-6}$ 510 adaptive $8\times 10^{-3} 11$0 adaptive $4.2\times 10^{-6}$ 210 adaptive $1.7\times 10^{-3}$ 6
+Q-ReSZO adaptive $2\times 10^{-6}$ 510 adaptive $5\times 10^{-3}$ 110 adaptive $1\times 10^{-5}$ 210 adaptive $1.8\times 10^{-3}$ 6
+Appendix B Proofs for Theoretical Convergence Analysis
+B.1 Theoretical justification for Assumption 2
 To facilitate our analysis, we first show the following result which provides an alternative characterization of the approximation error term $\|\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})\|$ that appears in Assumption 2.
@@
 Define the errors of the first-order Taylor expansion as:
-$\epsilon_{t}:=\begin{bmatrix}f(\hat{\bm{x}}_{t-1})-(f(\hat{\bm{x}}_{t})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t}\rangle)\\ \vdots\\ f(\hat{\bm{x}}_{t-m+1})-(f(\hat{\bm{x}}_{t})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t}\rangle)\end{bmatrix}.$ (30)
+$\epsilon_{t}:=\begin{bmatrix}f(\hat{\bm{x}}_{t-1})-(f(\hat{\bm{x}}_{t})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t}\rangle)\\ \vdots\\ f(\hat{\bm{x}}_{t-m+1})-(f(\hat{\bm{x}}_{t})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t}\rangle)\end{bmatrix}$.
+(31)
 Suppose $S_{t}S_{t}^{\top}\in\mathbb{R}^{d\times d}$ is invertible. Then,
@@
 as desired. ∎
-We first prove the following result (under a very mild condition stated in (31)) showing that Assumption 2 holds with $C_{d}=1$ in the case $d=1$. Define $[N]\coloneq\{1,2,\cdots,N\}$.
+We can further simplify the above result to an even simpler characterization.
 Lemma 4.
-Suppose Assumption 1 holds. Suppose $d=1$. Suppose also that there exists some $i\in[m-1]$ such that
-$|\hat{\bm{x}}_{t,i}-\hat{\bm{x}}|>0.$ (31)
-Then, with $\bm{g}_{t}$ being defined as in (19), for any $t\geq m$, Assumption 2 holds with $C_{d}=1$, i.e.
-$|\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})|\leq\frac{L}{2}\max_{i=1:m-1}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|.$
+Consider the cumulative sum matrix $C\in\mathbb{R}^{m\times m}$, defined as
+$C\;=\;\begin{pmatrix}1&0&0&\cdots&0\\ 1&1&0&\cdots&0\\ 1&1&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 1&1&1&\cdots&1\end{pmatrix}$.
+Define also the matrix
+$D_{t}^{\top}:=\begin{bmatrix}(\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t})^{\top}\\ (\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1})^{\top}\\ \vdots\\ (\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2})^{\top}\\ \end{bmatrix}$,
+as well as the following error vector
+$\hat{\epsilon}_{t}:=\begin{bmatrix}f(\hat{\bm{x}}_{t-1})-(f(\hat{\bm{x}}_{t})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t}\rangle)\\ f(\hat{\bm{x}}_{t-2})-\left(f(\hat{\bm{x}}_{t-1})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1}\rangle\right)\\ \vdots\\ f(\hat{\bm{x}}_{t-m+1})-\left(f(\hat{\bm{x}}_{t-m+2})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2}\rangle\right)\end{bmatrix}$.
+(32)
+Suppose $D$ has full rank. Then,
+$\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})=(D_{t}D_{t}^{\top})^{-1}D_{t}\hat{\epsilon}_{t}.$
 Proof.
-Note that $d=1$.
-Since we assume there exists some $i\in[m-1]$ such that $|\hat{\bm{x}}_{t,i}-\hat{\bm{x}}|>0$, we know that $S_{t}S_{t}^{\top}\in\mathbb{R}$ is invertible, satisfying the condition in Lemma 3.
-Thus, using Lemma 3, we note that the result is equivalent to showing that
-$|(S_{t}S_{t}^{\top})^{-1}S_{t}\epsilon_{t}|\leq\frac{L}{2}\max_{i=1:m-1}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|.$
-Let $\epsilon_{t,i}$ denote the $i$-th entry of $\epsilon_{t}\in\mathbb{R}^{m-1}$. By $L$-smoothness of $f$, we observe that for any $i\in[m-1]$, we have
-$|\epsilon_{t,i}|=\ |f(\hat{\bm{x}}_{t-i})-(f(\hat{\bm{x}}_{t})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\rangle)|\leq\ \frac{L}{2}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|^{2}.$
-Using the above result, in the case $d=1$, we observe that
-$|(S_{t}S_{t}^{\top})^{-1}S_{t}\epsilon_{t}|=$ $\ \Big{|}\frac{1}{\sum_{i=1}^{m-1}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|^{2}}\sum_{i=1}^{m-1}(\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t})\epsilon_{t,i}\Big{|}$
-$\leq$ $\ \frac{1}{\sum_{i=1}^{m-1}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|^{2}}\max_{i=1:m-1}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|\sum_{i=1}^{m-1}|\epsilon_{t,i}|$
-$\leq$ $\ \frac{1}{\sum_{i=1}^{m-1}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|^{2}}\max_{i=1:m-1}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|\sum_{i=1}^{m-1}\frac{L}{2}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|^{2}$
-$\leq$ $\ \frac{L}{2}\max_{i=1:m-1}|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}|,$
-as desired. ∎
-A.1.1 Technical Challenges of Proving Assumption 2 for General Cases with $d>1$
-From Lemma 3, suppose that $S_{t}S_{t}^{\top}$ is invertible, proving Assumption 2 is equivalent to showing that
-$\|(S_{t}S_{t}^{\top})^{-1}S_{t}\epsilon_{t}\|\leq\frac{C_{d}L}{2}\max_{i=1:m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|,$
-for some $C_{d}>0$. A natural approach to proving the above inequality is to use the standard bound
-$\|(S_{t}S_{t}^{\top})^{-1}S_{t}\epsilon_{t}\|\leq\|(S_{t}S_{t}^{\top})^{-1}\|_{2}\|S_{t}\epsilon_{t}\|.$
-However, it is (i) unclear how we might theoretically bound $\|(S_{t}S_{t}^{\top})^{-1}\|_{2}$, and (ii), empirically, when $d<m=\mathcal{O}(d)$ regression samples are used, on the ridge regression, we observe that $\|(S_{t}S_{t}^{\top})^{-1}\|_{2}$ scales as [^4] We recall that the notation $\Omega(\cdot)$ is defined as follows: for any non-negative variable $n$, saying that a positive function $f(n)$ is $\Omega(g(n))$, where $g(n)$ is some positive funcion of $n$, means that there exists some absolute constant $C_{1}>0$ and a number $n_{0}$ such that for all $n\geq n_{0}$, we have $f(n)\geq C_{1}g(n)>0$.
-$\|(S_{t}S_{t}^{\top})^{-1}\|_{2}=\Omega\left(\frac{d^{3}}{\sum_{i=1}^{m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|^{2}}\right).$
-Combining this with the argument in Lemma 4 bounding $\|\epsilon_{t,i}\|$ as
-$|\epsilon_{t,i}|=\ |f(\hat{\bm{x}}_{t-i})-(f(\hat{\bm{x}}_{t})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\rangle)|\leq\ \frac{L}{2}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|^{2},$
-it yields
-$\|(S_{t}S_{t}^{\top})^{-1}S_{t}\epsilon_{t}\|\leq$ $\ \|(S_{t}S_{t}^{\top})^{-1}\|_{2}\|S_{t}\epsilon_{t}\|$
-$\leq$ $\ \Omega\left(\frac{d^{3}}{\sum_{i=1}^{m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|^{2}}\right)\left\|\sum_{i=1}^{m-1}(\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t})\epsilon_{t,i}\right\|$
-$\leq$ $\ \Omega\left(\frac{d^{3}}{\sum_{i=1}^{m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|^{2}}\right)\max_{i=1:m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|\sum_{i=1}^{m-1}\frac{L}{2}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|^{2}$
-$\leq$ $\ \Omega\left(d^{3}\right)\frac{L}{2}\max_{i=1:m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|,$
-yielding $C_{d}=\Omega(d^{3})$, which as we later demonstrate empirically in Appendix C, vastly overestimates the true size of $C_{d}$ that tends to scale as $\mathcal{O}(\sqrt{d})$.
-This suggests that the natural approach of estimating $C_{d}$ by first bounding $\|(S_{t}S_{t}^{\top})^{-1}\|_{2}$ does not yield tight bounds; additionally, it is in the first place also non-trivial to theoretically prove an upper bound for $\|(S_{t}S_{t}^{\top})^{-1}\|_{2}$.
-Thus, we regard the task of accurately bounding $\|(S_{t}S_{t}^{\top})^{-1}S_{t}\epsilon_{t}\|$ as a very challenging theoretical problem which we leave to future work.
-A.2 Proofs for Lemma 1, Theorem 1, and Theorem 2
-Throughout the analysis, without loss of generality, we assume that $C_{d}\geq 1$.
-We also recall what we stated earlier, which is that we will study a variant of the L-RESZO algorithm with an adaptive choice of smoothing radius $\delta_{t}$ as (20),
-$\delta_{t}:=\eta\|\bm{g}_{t-1}\|.$
+From the earlier result, we saw that when $S$ is full rank (equivalent to $D$ being full rank), we have
+$\bm{g}_{t}=$ $\ (S_{t}S_{t}^{\top})^{-1}S_{t}\Delta_{t}^{f}=\ (S_{t}S_{t}^{\top})^{-1}S_{t}\left(S_{t}^{\top}\nabla f(\hat{\bm{x}}_{t})+\epsilon_{t}\right)$
+$=$ $\ \nabla f(\hat{\bm{x}}_{t})+(S_{t}S_{t}^{\top})^{-1}S_{t}\epsilon_{t}.$
+It can be easily verified that $S_{t}^{\top}=CD_{t}^{\top}$. Since $(S_{t}^{\top})^{\dagger}=(CD_{t}^{\top})^{\dagger}=(D_{t}^{\top})^{\dagger}C^{-1}$, we have that
+$\bm{g}_{t}=$ $\ \nabla f(\hat{\bm{x}}_{t})+(S_{t}S_{t}^{\top})^{-1}S_{t}\epsilon_{t}$
+$=$ $\ \nabla f(\hat{\bm{x}}_{t})+(D_{t}D_{t}^{\top})^{-1}D_{t}C^{-1}\epsilon_{t}.$
+Since $C^{-1}$ is the following matrix
+$C^{-1}\;=\;\begin{pmatrix}1&0&0&\cdots&0\\ -1&1&0&\cdots&0\\ 0&-1&1&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&0\\ 0&\cdots&0&-1&1\end{pmatrix}$,
+the desired result follows since $C^{-1}\epsilon_{t}=\hat{\epsilon}_{t}.$
+In order for Assumption 2 to hold, we need to rely on the following assumption on the smallest singular value of $D_{t}$.
+While difficult to prove since it involves computing the smallest singular value of random matrices with mild correlation across rows, we provide heuristic support for its validity.
+Assumption 3.
+Let $D_{t}^{\top}\in\mathbb{R}^{(m-1)\times d}$ be the matrix
+$D_{t}^{\top}:=\begin{bmatrix}(\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t})^{\top}\\ (\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1})^{\top}\\ \vdots\\ (\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2})^{\top}\\ \end{bmatrix}$.
+Then, the smallest singular value of $D_{t}^{\top}$ satisfies the following bound
+$\sigma_{\min}(D_{t}^{\top})\geq\Omega\left(\frac{1}{\Gamma_{d,m}}\sqrt{\sum_{i=1}^{m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t-i+1}\|^{2}}\right).$
+Here $\Gamma_{d,m}$ is a dimension-dependent constant satisfying
+$\Gamma_{d,m}=\Theta\!\left(\sqrt{d}\sqrt{m}\right).$
+Theoretical justification for Assumption 3.
+We note that the matrix $D_{t}^{\top}$ can also be written as
+$D_{t}^{\top}=$ $\ \begin{bmatrix}(\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t})^{\top}\\ (\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1})^{\top}\\ \vdots\\ (\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2})^{\top}\\ \end{bmatrix}$
+$=$ $\ \begin{bmatrix}(\hat{\bm{x}}_{t-1}-{\bm{x}}_{t-1})^{\top}+({\bm{x}}_{t-1}-{\bm{x}}_{t})^{\top}+({\bm{x}}_{t}-\hat{\bm{x}}_{t})^{\top}\\ (\hat{\bm{x}}_{t-2}-{\bm{x}}_{t-2})^{\top}+({\bm{x}}_{t-2}-{\bm{x}}_{t-1})^{\top}+({\bm{x}}_{t-1}-\hat{\bm{x}}_{t-1})^{\top}\\ \vdots\\ (\hat{\bm{x}}_{t-m+1}-{\bm{x}}_{t-m+1})^{\top}+({\bm{x}}_{t-m+1}-{\bm{x}}_{t-m+2})^{\top}+({\bm{x}}_{t-m+2}-\hat{\bm{x}}_{t-m+2})^{\top}\\ \end{bmatrix}$
+$=$ $\ \begin{bmatrix}(\bm{u}_{t-1})^{\top}+({\bm{x}}_{t-1}-{\bm{x}}_{t})^{\top}+(\bm{u}_{t})^{\top}\\ (\bm{u}_{t-2})^{\top}+({\bm{x}}_{t-2}-{\bm{x}}_{t-1})^{\top}+(\bm{u}_{t-1})^{\top}\\ \vdots\\ (\bm{u}_{t-m+1})^{\top}+({\bm{x}}_{t-m+1}-{\bm{x}}_{t-m+2})^{\top}+(\bm{u}_{t-m+2})^{\top}\\ \end{bmatrix}$
+$=$ $\ G_{t}^{\top}+H_{t}^{\top},$
+where
+$G_{t}^{\top}=$ $\ \begin{bmatrix}({\bm{x}}_{t-1}-{\bm{x}}_{t})^{\top}\\ ({\bm{x}}_{t-2}-{\bm{x}}_{t-1})^{\top}\\ \vdots\\ ({\bm{x}}_{t-m+1}-{\bm{x}}_{t-m+2})^{\top}\end{bmatrix},\quad H_{t}^{\top}=\begin{bmatrix}(\bm{u}_{t-1})^{\top}-(\bm{u}_{t})^{\top}\\ (\bm{u}_{t-2})^{\top}-(\bm{u}_{t-1})^{\top}\\ \vdots\\ (\bm{u}_{t-m+1})^{\top}-(\bm{u}_{t-m+2})^{\top}\\ \end{bmatrix}$,
+with $\bm{u_{j}}^{\top}$ is sampled independently and uniformly from the sphere with norm $\|\bm{x}_{j}-\bm{x}_{j-1}\|$.
+To find the smallest singular value of $D_{t}^{\top}$, we thus have to find the smallest eigenvalue of the matrix $(D_{t}D_{t}^{\top})=(G_{t}G_{t}^{\top}+G_{t}H_{t}^{\top}+H_{t}G_{t}^{\top}+H_{t}H_{t}^{\top})$.
+While $G_{t}$ and $H_{t}$ are slightly correlated, the cross-terms should display concentration effects and thus dominated by the effects of the terms $G_{t}G_{t}^{\top}+H_{t}H_{t}^{\top}$.
+To see this, taking $G_{t}H_{t}^{\top}$ for example, we observe that
+$G_{t}H_{t}^{\top}=\sum_{j=1}^{m-1}(\bm{x}_{t-j}-\bm{x}_{t-j+1})(\bm{u}_{t-j}-\bm{u}_{t-j+1})^{\top}.$ (33)
+Since $\bm{u}_{t-j+1}$ is independent of $\bm{x}_{t-j}-\bm{x}_{t-j+1}$, the term $\sum_{j=1}^{m-1}(\bm{x}_{t-j}-\bm{x}_{t-j+1})(\bm{u}_{t-j+1})^{\top}$ is mean zero and exhibits concentration of measure effects.
+Similarly, any correlation between $\bm{x}_{t-j}-\bm{x}_{t-j+1}$ and $\bm{u}_{t-j}$ is likely mild; when the iterates are generated before the regression starts, e.g., via one-point method, $\bm{x}_{t-j}-\bm{x}_{t-j+1}$ will have no relation to $\bm{u}_{t-j}$, since the random direction at iterate tjt-j is independent of the perturbation $\bm{u}_{t-j}$ which is only to obtain the function value evaluation.
+When $\bm{x}_{t-j}-\bm{x}_{t-j+1}=\eta\bm{g}_{t-j}$ is the result of a regression-based gradient estimation update, $\bm{g}_{t-j}$ is affected by $\bm{u}_{t-j}$, but only in one row, and thus unlikely to have significant correlation.
+We may thus assume that
+$\sigma_{\min}(D_{t}D_{t}^{\top})=\Theta(\sigma_{\min}(G_{t}G_{t}^{\top}+H_{t}H_{t}^{\top}))\geq\Theta(\sigma_{\min}(H_{t}H_{t}^{\top})).$
+Fully characterizing the distribution of $\sigma_{\min}(H_{t}H_{t}^{\top})$ is highly non-trivial since the norm of each row of $H_{t}^{\top}$ has a different variance, namely that the $j$-th row of $H_{t}^{\top}$, which is equal to $(\bm{u}_{t-j}-\bm{u}_{t-j+1})^{\top}$, has variance $\|\bm{u}_{t-j}\|^{2}+\|\bm{u}_{t-j+1}\|^{2}=\|\bm{x}_{t-j}-\bm{x}_{t-j-1}\|^{2}+\|\bm{x}_{t-j+1}-\bm{x}_{t-j}\|^{2}$.
+To proceed, note that $H_{t}^{\top}\in\mathbb{R}^{(m-1)\times d}$ has rows $h_{j}^{\top}:=(\bm{u}_{t-j}-\bm{u}_{t-j+1})^{\top}$ for $j=1,\dots,m-1$.
+To build intuition for the scale of $\sigma_{\min}(H_{t}^{\top})$, consider the Gram matrix
+$H_{t}H_{t}^{\top}\;=\;\sum_{j=1}^{m-1}h_{j}h_{j}^{\top}\;=\;\sum_{j=1}^{m-1}(\bm{u}_{t-j}-\bm{u}_{t-j+1})(\bm{u}_{t-j}-\bm{u}_{t-j+1})^{\top}.$
+Expanding the summand gives
+$(\bm{u}_{t-j}-\bm{u}_{t-j+1})(\bm{u}_{t-j}-\bm{u}_{t-j+1})^{\top}=\bm{u}_{t-j}\bm{u}_{t-j}^{\top}+\bm{u}_{t-j+1}\bm{u}_{t-j+1}^{\top}-\bm{u}_{t-j}\bm{u}_{t-j+1}^{\top}-\bm{u}_{t-j+1}\bm{u}_{t-j}^{\top}.$
+Summing over $j$ yields
+$H_{t}H_{t}^{\top}=\underbrace{\sum_{j=1}^{m-1}\bm{u}_{t-j}\bm{u}_{t-j}^{\top}+\sum_{j=1}^{m-1}\bm{u}_{t-j+1}\bm{u}_{t-j+1}^{\top}}_{\text{``diagonal'' terms}}\;-\;\underbrace{\sum_{j=1}^{m-1}\big(\bm{u}_{t-j}\bm{u}_{t-j+1}^{\top}+\bm{u}_{t-j+1}\bm{u}_{t-j}^{\top}\big)}_{\text{cross terms}}.$
+The diagonal part is essentially a weighted sum of rank-one matrices of the form $\bm{u}_{k}\bm{u}_{k}^{\top}$.
+The cross terms are mean-zero (conditionally on $\{\bm{x}_{k}\}$) because the directions $\bm{u}_{k}$ are independent across $k$ and isotropic, so $\mathbb{E}[\bm{u}_{t-j}\bm{u}_{t-j+1}^{\top}]=0$.
+Consequently, by concentration of measure, the cross terms exhibit cancellations and are typically of smaller spectral norm compared to the diagonal terms.
+This motivates the approximation
+$\sigma_{\min}(H_{t}H_{t}^{\top})\;\approx\;\sigma_{\min}\!\left(\sum_{k=t-m+2}^{t-1}\bm{u}_{k}\bm{u}_{k}^{\top}\right).$
+Next, assume that $\|\bm{u}_{k}\|^{2}$ are comparable over the window, i.e., $\|\bm{u}_{k}\|^{2}\asymp r^{2}$ for $k=t-m+2,\dots,t-1$.
+Under either the Gaussian model $\bm{u}_{k}\sim\mathcal{N}(0,\tfrac{r^{2}}{d}I_{d})$ or the uniform-sphere model (with comparable sub-Gaussian tails), the matrix with rows proportional to $\bm{u}_{k}^{\top}$ is a (sub-)Gaussian design with unit-variance entries up to scaling.
+Therefore, by classic results on the smallest singular value of sub-Gaussian random matrices (e.g., Theorem 1.1 in [21]), we have heuristically that
+$\sigma_{\min}\!\left(\sum_{k=t-m+2}^{t-1}\bm{u}_{k}\bm{u}_{k}^{\top}\right)\;=\;\Theta\!\left(\frac{r^{2}}{d}\right),$
+which implies
+$\sigma_{\min}(H_{t}^{\top})\;=\;\sqrt{\sigma_{\min}(H_{t}H_{t}^{\top})}\;\approx\;\Theta\!\left(\frac{r}{\sqrt{d}}\right).$
+Finally, using $\|\bm{u}_{k}\|^{2}\asymp r^{2}$ and m1m-1 terms in the window,
+$r\;\asymp\;\sqrt{\frac{1}{m-1}\sum_{k=t-m+2}^{t-1}\|\bm{u}_{k}\|^{2}},$
+so the above scaling can be rewritten as
+$\sigma_{\min}(H_{t}^{\top})\;\approx\;$ $\ \Theta\!\left(\frac{1}{\sqrt{d}}\sqrt{\frac{1}{m-1}\sum_{k=t-m+2}^{t-1}\|\bm{u}_{k}\|^{2}}\right)\;=\;\Theta\!\left(\frac{1}{\sqrt{d(m-1)}}\sqrt{\sum_{k=t-m+2}^{t-1}\|\bm{u}_{k}\|^{2}}\right)$
+$=$ $\ \Omega\left(\frac{1}{\sqrt{d}\sqrt{m}}\sqrt{\sum_{i=1}^{m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t-i+1}\|^{2}}\right).$
+We are now ready to prove Assumption 2, assuming that Assumption 3 holds.
+Proof of Assumption 2.
+Suppose Assumption 3 holds. Recall from Lemma 4 that
+$\bm{g}_{t}-\nabla f(\hat{\bm{x}}_{t})=(D_{t}D_{t}^{\top})^{-1}D_{t}\hat{\epsilon}_{t},$
+where
+$D_{t}^{\top}:=\begin{bmatrix}(\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t})^{\top}\\ (\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1})^{\top}\\ \vdots\\ (\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2})^{\top}\\ \end{bmatrix}$,
+and
+$\hat{\epsilon}_{t}:=\begin{bmatrix}f(\hat{\bm{x}}_{t-1})-(f(\hat{\bm{x}}_{t})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t}\rangle)\\ f(\hat{\bm{x}}_{t-2})-\left(f(\hat{\bm{x}}_{t-1})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1}\rangle\right)\\ \vdots\\ f(\hat{\bm{x}}_{t-m+1})-\left(f(\hat{\bm{x}}_{t-m+2})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2}\rangle\right)\end{bmatrix}$.
+From Assumption 3, we have that
+$\sigma_{\min}(D_{t}^{\top})\geq\Omega\left(\frac{1}{\Gamma_{d,m}}\sqrt{\sum_{i=1}^{m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t-i+1}\|^{2}}\right).$
+It follows then that
+$\|\bm{g}_{t}-\nabla f(\hat{x}_{t})\|\leq\frac{1}{\sigma_{\min}(D_{t})}\|\hat{\epsilon}_{t}\|\leq\frac{\mathcal{O}(\sqrt{d}\sqrt{m})}{\sqrt{\sum_{i=1}^{m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t-i+1}\|^{2}}}\|\hat{\epsilon}_{t}\|.$
+Using Taylor expansion, note we can also write $\hat{\epsilon}_{t}$ as
+$\hat{\epsilon}_{t}:=$ $\ \begin{bmatrix}f(\hat{\bm{x}}_{t-1})-(f(\hat{\bm{x}}_{t})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t}\rangle)\\ f(\hat{\bm{x}}_{t-2})-\left(f(\hat{\bm{x}}_{t-1})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1}\rangle\right)\\ \vdots\\ f(\hat{\bm{x}}_{t-m+1})-\left(f(\hat{\bm{x}}_{t-m+2})+\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2}\rangle\right)\\ \end{bmatrix}$
+$=$ $\ \begin{bmatrix}\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t}\rangle+\mathcal{O}(\frac{L}{2})\|\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t}\|^{2}-(\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t}\rangle)\\ \langle\nabla f(\hat{\bm{x}}_{t-1}),\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1}\rangle+\mathcal{O}(\frac{L}{2})\|\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1}\|^{2}-(\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1}\rangle)\\ \vdots\\ \langle\nabla f(\hat{\bm{x}}_{t-m+2}),\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2}\rangle+\mathcal{O}(\frac{L}{2})\|\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2}\|^{2}-(\langle\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2}\rangle)\\ \end{bmatrix}$
+$=$ $\ \begin{bmatrix}\langle\nabla f(\hat{\bm{x}}_{t})-\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t}\rangle+\mathcal{O}(\frac{L}{2})\|\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t}\|^{2}\\ \langle\nabla f(\hat{\bm{x}}_{t-1})-\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1}\rangle+\mathcal{O}(\frac{L}{2})\|\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1}\|^{2}\\ \vdots\\ \langle\nabla f(\hat{\bm{x}}_{t-m+2})-\nabla f(\hat{\bm{x}}_{t}),\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2}\rangle+\mathcal{O}(\frac{L}{2})\|\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2}\|^{2}\\ \end{bmatrix}$
+$=$ $\ \begin{bmatrix}\mathcal{O}(\frac{L}{2})\|\hat{\bm{x}}_{t-1}-\hat{\bm{x}}_{t}\|^{2}\\ \mathcal{O}(L\|\hat{\bm{x}}_{t-1})-\nabla f(\hat{\bm{x}}_{t})\|\|\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1}\|)+\mathcal{O}(\frac{L}{2})\|\hat{\bm{x}}_{t-2}-\hat{\bm{x}}_{t-1}\|^{2}\\ \vdots\\ \mathcal{O}(L\|\hat{\bm{x}}_{t-m+1})-\nabla f(\hat{\bm{x}}_{t})\|\|\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2}\|)+\mathcal{O}(\frac{L}{2})\|\hat{\bm{x}}_{t-m+1}-\hat{\bm{x}}_{t-m+2}\|^{2}\\ \end{bmatrix}$.
+Thus it follows that
+$\|\hat{\epsilon}_{t}\|\leq\max\{\max_{i=1:m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|,\max_{i=1:m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t-i+1}\|\}\sqrt{\sum_{i=1}^{m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t-i+1}\|^{2}}$
+It follows then that
+$\|\bm{g}_{t}-\nabla f(\hat{x}_{t})\|\leq$ $\ \frac{1}{\sigma_{\min}(D_{t})}\|\hat{\epsilon}_{t}\|\leq\frac{\mathcal{O}(\sqrt{d}\sqrt{m})}{\sqrt{\sum_{i=1}^{m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t-i+1}\|^{2}}}\|\hat{\epsilon}_{t}\|$
+$\leq$ $\ \mathcal{O}(\sqrt{d}\sqrt{m})\max\{\max_{i=1:m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t}\|,\max_{i=1:m-1}\|\hat{\bm{x}}_{t-i}-\hat{\bm{x}}_{t-i+1}\|\},$
+i.e.
+the relation $C_{d,m}=\mathcal{O}(\sqrt{d}\sqrt{m})$ holds.
+[^2] We note that this is slightly different from what we stated in Assumption 2, but this result is identical to that in Assumption 2 for the purposes of our proofs.
+B.2 Proofs for Lemma 1, Theorem 1, and Theorem 2
+Throughout the analysis, without loss of generality, we denote $C_{d}:=C_{d,m}$ and assume that $C_{d}\geq 1$.
 We first prove the following useful helper result, which bounds the change in gradient norm across $m$ iterations whenever the gradient estimation bias can be bounded in terms of the true gradient norm.
 Lemma 5.
-Suppose Assumption 1 holds. Suppose $2\leq m=\mathcal{O}(d)$. Consider any $t\geq m$. Suppose that
+Suppose Assumption 1 holds. Suppose $m>d$. Consider any $t\geq m$. Suppose that
 $\|\bm{x}_{j+1}-\bm{x}_{j}\|\leq 2\eta\|\nabla f(\bm{x}_{j})\|$ (34)
 for all $j<t$.
-Then, by picking $\eta=\mathcal{O}\left(\frac{1}{dL}\right)$, we can ensure that
+Then, by picking $\eta=\mathcal{O}\left(\frac{1}{mL}\right)$, we can ensure that
 $\|\nabla f(\bm{x}_{t-j})\|\leq 4\|\nabla f(\bm{x}_{t})\|$
@@
 we have that
-[…]
+$\|\nabla f(\bm{x}_{j+1})-\nabla f(\bm{x}_{j})\|\leq\ L\|\bm{x}_{j+1}-\bm{x}_{j}\|\leq\ 2\eta L\|\nabla f(\bm{x}_{j})\|,$ (35)
 where the first line utilizes $L$-smoothness and the second line utilizes our assumption in (34). Continuing, we observe that
@@
 where the final inequality follows from the fact that the sequence $(\frac{n}{n-1})^{n}\leq 4$ for all $n\geq 2$. ∎
-Next, we restate and prove Lemma 1, which shows that for any $10<c<$ 1, we can inductively guarantee that the gradient estimation bias $\|\xi_{t}\|$ always satisfies $\|\xi_{t}\|\leq c\|\nabla f(\bm{x}_{t})\|$ for all $t\geq m$.
+Next, we restate and prove Lemma 1, which shows that for any $0<c<1$, we can inductively guarantee that the gradient estimation bias $\|\xi_{t}\|$ always satisfies $\|\xi_{t}\|\leq c\|\nabla f(\bm{x}_{t})\|$ for all $t\geq m$.
 See 1
@@
 $\|\bm{x}_{j+1}-\bm{x}_{j}\|\leq 2\eta\|\nabla f(\bm{x}_{j})\|$
-holds for all $j<t$. Then, for any $10<c<$ 1, by picking $\eta=C\frac{1}{dC_{d}L}$ for some sufficiently small absolute constant $C>0$, we will show that the inductive claim implies that
+holds for all $j<t$. Then, for any $0<c<1$, by picking $\eta=C\frac{1}{mC_{d}L}$ for some sufficiently small absolute constant $C>0$, we will show that the inductive claim implies that
 $\ \|\xi_{t}\|\leq c\|\nabla f(\bm{x}_{t})\|$
@@
 $=$ $\ \frac{C_{d}L}{2}\max_{i=1:m-1}\left\|\sum_{j=1}^{i}(\bm{x}_{t-j}-\bm{x}_{t-j+1}+\eta_{t-j-1}\|\bm{g}_{t-j-1}\|\bm{u}_{t-j}-\eta_{t-j}\|\bm{g}_{t-j}\|\bm{u}_{t-j+1})\right\|$
-$\ \quad\quad+\eta_{t-1}L\|\bm{g}_{t-1}\|$
+$\ \qquad\qquad+\eta_{t-1}L\|\bm{g}_{t-1}\|$
 $\stackrel{{\scriptstyle\textnormal{(iv)}}}{{\mathstrut{\leq}}}$ $\ \frac{C_{d}L}{2}\max_{i=1:m-1}\left(\sum_{j=1}^{i}\left(\eta_{t-j}\left\|\bm{g}_{t-j}\right\|+\eta_{t-j-1}\|\bm{g}_{t-j-1}\|+\eta_{t-j}\|\bm{g}_{t-j}\|\right)\right)+\eta_{t-1}L\|\bm{g}_{t-1}\|$
@@
 $\leq$ $\ \frac{C_{d}L}{2}\sum_{j=1}^{m}4\eta_{t-j}\|\bm{g}_{t-j}\|$
-=$\stackrel{{\scriptstyle\textnormal{(vi)}}}{{\mathstrut{=}}}$ $\ \frac{C_{d}L}{2}\sum_{j=1}^{m}4\|\bm{x}_{t-j+1}-\bm{x}_{t-j}\|$
+$\stackrel{{\scriptstyle\textnormal{(vi)}}}{{\mathstrut{=}}}$ $\ \frac{C_{d}L}{2}\sum_{j=1}^{m}4\|\bm{x}_{t-j+1}-\bm{x}_{t-j}\|$
 $\stackrel{{\scriptstyle\textnormal{(vii)}}}{{\mathstrut{\leq}}}$ $\ 2C_{d}L\sum_{j=1}^{m}2\eta\|\nabla f(\bm{x}_{t-j})\|,$
-where (A.2) holds by $L$-smoothness of $f$, (A.2) holds by definition of $\hat{\bm{x}}_{t}:=\hat{\bm{x}}_{t}+\eta_{t-1}\|\bm{g}_{t-1}\|\bm{u}_{t}$ due to the adaptive selection of the smoothing radius $\delta_{t}:=\eta\|\bm{g}_{t-1}\|$ and the fact that each $\bm{u}_{t}$ has unit norm, (A.2) utilizes Assumption 2, (A.2) again uses the fact that each $\bm{u}_{t}$ has a unit norm, (A.2) holds by our assumption at the start of this subsection that $C_{d}\geq 1$, (A.2) holds since $\bm{x}_{t-j+1}=\bm{x}_{t-j}+\eta_{t-j}\bm{g}_{t-j}$, and (A.2) follows by our inductive assumption.
+where (B.2) holds by $L$-smoothness of $f$, (B.2) holds by definition of $\hat{\bm{x}}_{t}:=\hat{\bm{x}}_{t}+\eta_{t-1}\|\bm{g}_{t-1}\|\bm{u}_{t}$ due to the adaptive selection of the smoothing radius $\delta_{t}:=\eta\|\bm{g}_{t-1}\|$ and the fact that each $\bm{u}_{t}$ has unit norm, (B.2) utilizes Assumption 2, (B.2) again uses the fact that each $\bm{u}_{t}$ has a unit norm, (B.2) holds by our assumption at the start of this subsection that $C_{d}\geq 1$, (B.2) holds since $\bm{x}_{t-j+1}=\bm{x}_{t-j}+\eta_{t-j}\bm{g}_{t-j}$, and (B.2) follows by our inductive assumption.
 Now, utilizing our inductive assumption, we may also use Lemma 5 to find that for any $j\in[m]$, $\|\nabla f(\bm{x}_{t-j})\|\leq 4\|\nabla f(\bm{x}_{t})\|$. Thus, continuing from above, we find that
@@
 $\leq$ $\ 16C_{d}L\eta\sum_{j=1}^{m}\|\nabla f(\bm{x}_{t})\|.$
-Now, (recalling that $m=\mathcal{O}(d)$), by picking $\eta=\frac{C}{dC_{d}L}$ for a sufficiently small $C>0$, we may conclude that
+Now, by picking $\eta=\frac{C}{mC_{d}L}$ for a sufficiently small $C>0$, we may conclude that
 $\ \|\xi_{t}\|\leq c\|\nabla f(\bm{x}_{t})\|,$
@@
 as desired. ∎
+We next restate and prove Lemma 2.
+See 2
+Proof.
+Consider any positive integer $t\geq m$.
+Fix any $0<c<1$.
+By our assumptions, applying Lemma 1, we find that there exists a sufficiently small absolute constant $C>0$ such that when $\eta=\frac{C}{mC_{d}L}$, we have
+$\|\xi_{t}\|\leq c\|\nabla f(\bm{x}_{t})\|$
+for any $t\geq m$.
+Next, observe that
+$\ f(\bm{x}_{t+1})-f(\bm{x}_{t})$
+$\stackrel{{\scriptstyle\textnormal{(i)}}}{{\mathstrut{\leq}}}$ $\ -\eta\langle\nabla f(\bm{x}_{t}),\bm{g}_{t}\rangle+\frac{L}{2}\|\eta\bm{g}_{t}\|^{2}$
+$=$ $\ -\eta\langle\nabla f(\bm{x}_{t}),\nabla f(\bm{x}_{t})\rangle-\eta\langle\nabla f(\bm{x}_{t}),\bm{g}_{t}-\nabla f(\bm{x}_{t})\rangle$
+$\quad\quad+\frac{L}{2}\|\eta(\bm{g}_{t}-\nabla f(\bm{x}_{t})+\nabla f(\bm{x}_{t}))\|^{2}$
+$\stackrel{{\scriptstyle\textnormal{(ii)}}}{{\mathstrut{\leq}}}$ $\ -\eta\|\nabla f(\bm{x}_{t})\|^{2}+\frac{\eta}{2}\left(\|\nabla f(\bm{x}_{t})\|^{2}+\|\xi_{t}\|^{2}\right)$
+$\quad\quad+\eta^{2}L\left(\|\nabla f(\bm{x}_{t})\|^{2}+\|\xi_{t}\|^{2}\right)$
+$\leq$ $\ -\frac{\eta}{4}\|\nabla f(\bm{x}_{t})\|^{2}+\eta\|\xi_{t}\|^{2},$ (36)
+where (B.2) follows by $L$-smoothness of $f$, (B.2) follows by application of Young’s equality as well as the notation $\xi_{t}:=\bm{g}_{t}-\nabla f(\bm{x}_{t})$, and the final inequality follows by choosing $\eta\leq\frac{1}{4L}$.
+Thus, continuing from (36), by picking $c$ to be smaller than $\frac{1}{\sqrt{8}}$, such that
+$\|\xi_{t}\|\leq\frac{1}{\sqrt{8}}\|\nabla f(\bm{x}_{t})\|,$
+we find that
+$f(\bm{x}_{t+1})-f(\bm{x}_{t})\leq$ $\ -\frac{\eta}{4}\|\nabla f(\bm{x}_{t})\|^{2}+\eta\|\xi_{t}\|^{2}.$
+$\leq$ $\ -\frac{\eta}{4}\|\nabla f(\bm{x}_{t})\|^{2}+\frac{\eta}{8}\|\nabla f(\bm{x}_{t})\|^{2}$
+$\leq$ $\ -\frac{\eta}{8}\|\nabla f(\bm{x}_{t})\|^{2},$
+as desired.
 We next restate and prove Theorem 1, which provides a convergence result for smooth nonconvex functions.
@@
 $f(\bm{x}_{t+1})-f(\bm{x}_{t})\leq-\frac{\eta}{8}\|\nabla f(\bm{x}_{t})\|^{2}.$
-Summing this from $t=m$ to $t=T$ and plugging in our choice of $\eta=\mathcal{O}\left(\frac{1}{dC_{d}L}\right)$, we observe that
-$\ f(\bm{x}_{T})-f(\bm{x}_{m})\leq-\mathcal{O}\left(\frac{1}{dC_{d}L}\sum_{t=m}^{T-1}\|\nabla f(\bm{x}_{t})\|^{2}\right)$
-$\implies$ $\ \mathcal{O}\left(\frac{1}{dC_{d}L}\sum_{t=m}^{T-1}\|\nabla f(\bm{x}_{t})\|^{2}\right)\leq f(\bm{x}_{m})-f(\bm{x}_{T})$
-$\implies$ $\ \frac{1}{T}\sum_{t=m}^{T-1}\|\nabla f(\bm{x}_{t})\|^{2}\leq\mathcal{O}\left(\frac{dC_{d}L\left(f(\bm{x}_{m})-f^{*}\right)}{T}\right),$
+Summing this from $t=m$ to $t=T$ and plugging in our choice of $\eta=\mathcal{O}\left(\frac{1}{mC_{d}L}\right)$, we observe that
+$\ f(\bm{x}_{T})-f(\bm{x}_{m})\leq-\mathcal{O}\left(\frac{1}{mC_{d}L}\sum_{t=m}^{T-1}\|\nabla f(\bm{x}_{t})\|^{2}\right)$
+$\implies$ $\ \mathcal{O}\left(\frac{1}{mC_{d}L}\sum_{t=m}^{T-1}\|\nabla f(\bm{x}_{t})\|^{2}\right)\leq f(\bm{x}_{m})-f(\bm{x}_{T})$
+$\implies$ $\ \frac{1}{T}\sum_{t=m}^{T-1}\|\nabla f(\bm{x}_{t})\|^{2}\leq\mathcal{O}\left(\frac{mC_{d}L\left(f(\bm{x}_{m})-f^{*}\right)}{T}\right),$
 where the final inequality uses the simple fact that $f(\bm{x}_{t})\geq f^{*}$. ∎
@@
 where the fourth inequality follows by $\mu$-strong convexity which ensures that $\|\nabla f(\bm{x}_{t})\|^{2}\geq 2\mu(f(\bm{x}_{t})-f^{*})$. Our desired result then follows. ∎
-Appendix B Empirical Study of $C_{d}$ in Assumption 2
+Appendix C Empirical Study of $C_{d}$ in Assumption 2
+We assume throughout this section that $m=\mathcal{O}(d)$, so we may denote $C_{d}:=C_{d,m}$.
 We recall from Assumption 2 that the $C_{d}$ term is defined as a positive ratio term such that
@@
 and use our analytical knowledge of the ridge and logistic regression problems in (27) and (28) to compute the smoothness constant $L$.
-B.1 Ridge Regression
+C.1 Ridge Regression
 We experimented on the ridge regression problem in (27) for three values for the dimension $d$: $d=100,400,900$.
@@
 (c) $d=900$
-Figure 4: Convergence of L-RESZO on the ridge-regression problem in (26) for dimensions $d=100,400,900$. The $C_{d}$ ratios shown in Figure 3 are based on the algorithmic trajectories here.
-B.2 Logistic Regression
+Figure 4: Convergence of L-ReSZO on the ridge‐regression problem in (27) for dimensions $d=100,400,900$. The $C_{d}$ ratios shown in Figure 3 are based on the algorithmic trajectories here.
+C.2 Logistic Regression
 We experimented on the logistic regression problem (28) for three values for the dimension $d$: $d=100,400,900.$ We computed the $C_{d}$ ratio statistics for a trial of the algorithm for these three dimensions, and display this information in Table 4, which is based on the data from Figure 5; the corresponding convergence behavior is shown in Figure 6.